### v1

# Root Number Equidistribution for Self-Dual Automorphic Representations on $\mathrm{GL}_{N}$

Abstract.
Let $F$ be a totally real field. We study the root numbers $\epsilon(1/2,\pi)$ of self-dual cuspidal automorphic representations $\pi$ of $\mathrm{GL}_{2N}/F$ with conductor $\mathfrak{n}$ and regular integral infinitesimal character $\lambda$. If $\pi$ is orthogonal, then $\epsilon(1/2,\pi)$ is known to be identically one. We show that for symplectic representations, the root numbers $\epsilon(1/2,\pi)$ equidistribute between ±$\pm 1$ as $\lambda\to\infty$, provided that there exists a prime dividing $\mathfrak{n}$ with power $>N$.We also study conjugate self-dual representations with respect to a CM extension $E/F$, where we obtain a similar result under the assumption that $\mathfrak{n}$ is divisible by a large enough power of a ramified prime and provide evidence that equidistribution does not hold otherwise. In cases where there are known to be associated Galois representations (e.g by [Shi11]), we deduce root number equidistribution results for the corresponding families of $N$-dimensional Galois representations.
The proof generalizes a classical argument for the the case of $\mathrm{GL}_{2}/\mathbb{Q}$ by using Arthur’s trace formula and the endoscopic classification for quasisiplit classical groups similarly to a previous work ([DGG22]). The main new technical difficulty evaluating endoscopic transfers of the required test functions at central elements.
1. Introduction
1.1. Context
Recall a classical open question: consider the set of elliptic curves $E/\mathbb{Q}$ with conductor $\leq n$. The BSD conjecture leads us to expect that the parity of the rank of the group of $\mathbb{Q}$-points of $E$ depends on the root number $\epsilon(1/2,E)=\pm 1$ of the corresponding $L$-function. We can therefore ask whether this root number is equidistributed between ±$\pm 1$ as $n\to\infty$. This paper studies an automorphic , weight-aspect variant of this question for higher-rank groups. Under very mild technical assumptions, we prove that equidistribution holds in all cases where it should be expected, see Theorem 1.2.1.
1.1.1. Automorphic Translation
By the modularity theorem, every elliptic curve $E/\mathbb{Q}$ of conductor $n$ corresponds to a cuspidal, weight-$2$ modular (eigen)newform $f$ of level $n$ with matching $L$-function/root number: $\epsilon(1/2,E)=\epsilon(1/2,f)$. We can therefore ask instead if the root numbers of cuspidal weight-$2$ modular newforms of level $n$ equidistribute as $n\to\infty$. This does not amount to computing the same statistic for rational elliptic curves since there are many extra modular forms. However, it is more natural from some perspectives : by the modularity theorem, cuspidal weight-$2$ newforms are in bijection with irreducible $2$-dimensional representations of $\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ with Hodge-Tate weights $\{0,1\}$. For some applications, counting these Galois representations may be more natural than counting elliptic curves.e
Most importantly for our purposes, counting modular forms opens the way for the application of trace formula techniques. These have been used to prove equidistribution in [SZ88, Mar18, Mar23] and are effective enough to compute finer statistics such as lower-order biases as in the results of Martin. Trace formulas even allow for computing subtle correlations between root numbers and Hecke eigenvalues depending on second-order estimates—for example, the “murmurations” phenomenon proven in [Zub23].
1.1.2. Generalization to Higher Dimensions
Recall that cuspidal modular newforms correspond to certain automorphic representations $\pi$ on $\mathrm{GL}_{2}/\mathbb{Q}$ with regular, integral infinitesimal character at infinity. Any automorphic representation factors as a restricted direct product over places $v$ of $\mathbb{Q}$:
$\pi={\bigotimes_{v}}^{\prime}\pi_{v},$
with each $\pi_{v}$ an irrep of $\mathrm{GL}_{2}(\mathbb{Q}_{v})$. In this context, the generalization of weight of a modular form is the so-called infinitesimal character of $\pi_{\infty}$. The automorphic representations corresponding to holomorphic modular forms are those where $\pi_{\infty}$ has regular, integral infinitesimal character—i.e. the same as a finite-dimensional irrep and therefore as “simple” as possible. Furthermore, as pointed out in [BG14], having integral infinitesimal character is exactly the expected condition needed to have an associated Galois representation. Thus automorphic representations of $GL_{N}/\mathbb{Q}$ with regular, integral infinitesimal character appear to be a natural generalization of holomorphic modular forms.
Generalizations of modular forms actually satisfy one more condition. For automorphic representations $\pi$ on $\mathrm{GL}_{2}$, having regular integral infinitesimal character at infinity is equivalent to $\pi_{\infty}$ being either discrete series or trivial. This discrete-at-infinity requirement can be thought of as a “niceness” condition; that these automorphic representations are the simplest possible in the same way that modular forms of weight $\geq 2$ are simpler than Maass forms or weight-$1$ forms. Unfortunately, for $N>2$$\mathrm{GL}_{N}(\mathbb{R})$ has no discrete series, so it at first appears that modular forms do not actually generalize to higher dimension.
A good resolution is to consider automorphic representations $\pi$ that are either:
  • self-dual : $\pi$ is an automorphic representation for $\mathrm{GL}_{N}/F$ for $F$ a totally real number field and $\pi^{\vee}\cong\pi$.
  • conjugate self-dual : $\pi$ is an automorphic representation for $\mathrm{GL}_{N}/E$ for $E/F$ a CM quadratic extension of number fields and $\pi^{\vee}\cong\bar{\pi}$ (i.e. the $E/F$-conjugate).
Any (conjugate) self-dual $\pi$ has component $\pi_{\infty}$ that is also (conjugate) self-dual. Moreover, such $\pi_{\infty}$ with regular enough and integral infinitesimal character are discrete within the (conjugate) self-dual spectrum. Lastly, these are also the automorphic representations to which, even in higher dimension, there are expected to be attached Galois representations—e.g. by work of Shin and collaborators [Shi11], [KS20], [KS23] in some cases and [Shi24] for a “weak transfer” result in all cases.
Self-dual automorphic representations can be classified into two types: orthogonal- or symplectic-type. This should be thought of as encoding whether the conjectural associated Galois representations preserve a symmetric or skew-symmetric bilinear form (the precise definition will be recalled in §2.3 and depends on [Art13]). One of the outputs of the endoscopic classification [Art13] is that orthogonal-type automorphic representations have root number determined by their central character, and as such, easily understood. In the self-dual case, the most interesting statistics are those of the the symplectic-type automorphic representations. We similarly restrict our attention to the conjugate-symplectic case for conjugate self-dual representations, this time justified by results in the unitary endoscopic classification [Mok15].
Note that in the self-dual case, restriction to symplectic-type is also natural from the Galois side. The natural generalization of elliptic curves is higher-dimensional abelian varieties. Just like elliptic curves, these are conjecturally associated to automorphic representations through their Galois representations. However, Galois representations of abelian varieties always preserve the symplectic Weil pairing, so their associated automorphic representations should be symplectic-type.
In summary, it is most appropriate to consider the following generalization of modular forms: (conjugate) symplectic-type automorphic representations of $\mathrm{GL}_{N}$ with regular, integral infinitesimal character at infinity.
1.2. Results
1.2.1. Setup
To make the results precise, fix a totally real number field $F$. We consider two cases:
  • self-dual: define $G_{N}=\mathrm{GL}_{N}/F$,
  • conjugate self-dual: define $G_{N}=\operatorname{Res}^{E}_{F}\mathrm{GL}_{N}$.
Let $\mathcal{AR}_{\mathrm{cusp}}(G_{N})$ be the set of cuspidal automorphic representations for $G_{N}$—i.e those on $\mathrm{GL}_{N}(F)\backslash\mathrm{GL}_{N}(\mathbb{A}_{F})$ in the self-dual case and those on $\mathrm{GL}_{N}(E)\backslash\mathrm{GL}_{N}(\mathbb{A}_{E})$ in the conjugate self-dual case.
As discussed in §1.1.2 and made precise in §2.3, there is a subset of $\mathcal{AR}_{\mathrm{cusp}}(G_{N})$ of (conjugate) self-dual representations which can be partitioned in to (conjugate) symplectic and (conjugate) orthogonal pieces. Call these $\mathcal{AR}^{\mathrm{symp}}_{\mathrm{cusp}}(G_{N})$ and $\mathcal{AR}^{\mathrm{orth}}_{\mathrm{cusp}}(G_{N})$ respectively.
[…]
Finally, every (conjugate) self-dual representation $\pi$ has a root number invariant $\epsilon(1/2,\pi)=\pm 1$ (see §3.3.2 for normalization details). We are interested in the average value
$\lim_{\begin{subarray}{c}\lambda\to\infty\\ \lambda\text{ integral}\end{subarray}}\frac{1}{|\mathcal{AR}^{\star}_{N}(\lambda,\mathfrak{n})|}\sum_{\pi\in\mathcal{AR}^{\star}_{N}(\lambda,\mathfrak{n})}\epsilon(1/2,\pi).$
for $\star=\mathrm{orth},\mathrm{symp}$ and where $\lambda\to\infty$ is as a dominant weight of $G_{\infty}$, i.e. so that the minimum difference between elements in each tuple $\lambda_{v}$ goes to $\infty$; see §5.6.
1.2.2. Main Theorem
First, by the endoscopic classifications [Art13] and [Mok15], the root number $\epsilon(1/2,\pi)$ of $\pi\in\mathcal{AR}^{\mathrm{orth}}_{\mathrm{cusp}}(G_{N})$ depends only on the central character of $\pi$. Therefore we focus on the symplectic-type representations. We also assume that $N$ is even: this is only a restriction in the conjugate self-dual case since there are no self-dual, symplectic-type representations with $N$ odd.
Our main result is that $\epsilon(1/2,\pi)$ equidistributes under these conditions:
Theorem 1.2.1 (10.2.1, 10.3.2).
Let $N$ be even and let $\mathfrak{n}=\prod_{v}\mathfrak{p}_{v}^{e_{v}}$ be such that:
  • One of the $e_{v}$ is either odd or greater than $N$ (self-dual case),
  • One of the $e_{v}$ is large enough for one of the ramified $\mathfrak{p}_{v}$ and $\mathfrak{n}$ is coprime to $2$ (conjugate self-dual case).
Then
$\lim_{\begin{subarray}{c}\lambda\to\infty\\ \lambda\text{ integral}\end{subarray}}\frac{1}{|\mathcal{AR}^{\mathrm{symp}}_{N}(\lambda,\mathfrak{n})|}\sum_{\pi\in\mathcal{AR}^{\mathrm{symp}}_{N}(\lambda,\mathfrak{n})}\epsilon(1/2,\pi)\to 0.$
Furthermore, in the self-dual case, the conditions on $\mathfrak{n}$ are also necessary.
Theorems 10.2.1 and 10.3.2 are in fact stronger: they show that averages of $\epsilon(1/2,\pi)$ are 00 even when weighted by $\operatorname{tr}_{\pi_{S}}(f_{S})$ for any “positive” unramified test function $f_{S}$ at a finite set of places $S$ coprime to $\mathfrak{n}$, see Definition 10.1.2. This is analogous to weighting the average on $\mathrm{GL}_{2}$ by a classical Hecke eigenvalue. The main theorems also provide a bound on the rate of convergence in terms of this weighting.
Finally, by Conjecture 8.4.3, we expect the precise conditions on $\mathfrak{n}$ in Theorem 10.3.2 to also be necessary conditions as long as there is no obstruction to the existence of $\pi\in\mathcal{AR}^{\mathrm{symp}}_{N}(\lambda,\mathfrak{n})$ coming from the central character of $\pi^{\infty}$.
Note.
It is somewhat surprising that equidistribution in this conjugate self-dual setting seems to only holds at ramified level. In fact, checking examples on [LMF] seems to suggest that the root number of a local conjugate self-dual representation over an unramified quadratic extension is determined by its conductor.
1.2.3. Side Effects
Our techniques have two interesting side effects:
First, one of the key technical tools is a “weight-aspect” version of the bounds in [DGG22] that also applies to the self-dual setting (§5, 6). This is of independent interest—for example, Theorem 6.5.1 is a key input into the proving that the gate sets for quantum circuits constructed in [DEP] achieve the bounds needed to be so-called “golden gates”.
Second, we prove a statement (Theorem 9.1.2) roughly stating that the Plancherel measure on a $p$-adic quasisplit classical group $G_{v}$ is “constant” on $L$-packets. In particular, this shows that all elements of a given supercuspidal $L$-packet for $G_{v}$ have the same formal degree (Corollary 9.1.3).
1.2.4. Methods and Potential Extensions
We rely on generalizations of the the classical trace formulas that were so effective in studying root numbers of modular forms [Mar23, Zub23]. Specifically, through the stabilization of the twisted trace formula as used in the endoscopic classifications of [Art13] and [Mok15], we produce an expression for our desired statistic made up of terms as in [Art89], which have further been studied and bounded in [ST16]. This analysis builds on the authors’ previous works [Dal22] and [DGG22]. It should be compared to an alternate, non-endoscopic strategy for analyzing the twisted trace formula in [TW24].
These more intricate methods currently come with serious costs (beyond just the length of the argument). First, the very precise bounds needed for the finer estimates of [Mar23] and especially [Zub23] depend on interpretations of terms on the geometric side of the trace formula in terms of $L$-values/class numbers. These are currently only available in the case of untwisted $\mathrm{GL}_{N}$ from [Yun13].
Second, it would be desirable to have asymptotics in the level aspect $\mathfrak{n}\to\infty$ in addition to $\lambda\to\infty$. The complexity of the test functions $\widetilde{C}^{\infty}_{\mathfrak{n}}$ and $\widetilde{E}^{\infty}_{\mathfrak{n}}$ constructed in §4 make this hard: we would require explicit asymptotics in $\mathfrak{n}$ of their orbital integrals instead of just the very coarse information computed in §§7-9. We are still hopeful that the $\mathfrak{n}\to\infty$ limit is tractable with more work and leave it for a future paper, possibly using techniques from [KY12] and [TW24].
Finally, our techniques fall short in the conjugate self-dual case for $N$ odd. Conceptually, we are really studying $(N-1)$st powers of root numbers by taking advantage of the particular simplicity of the $N\times(N-1)$ Rankin-Selberg integral representation: that the involution on functions that proves this functional equation is “geometric”—i.e. comes from an involution of the group instead of something like a Fourier transform in the case of Godement-Jacquet. Our parity restriction is also related to the miraculous vanishing of signs in twisted endoscopic transfers for odd $N$ that was key to the main result of [AOY22]. In our context, these signs exactly carry the information of root numbers we want to probe, so their vanishing is a tragedy instead of a miracle.
1.3. Application to Galois Representations
In the conjugate self-dual case when $N$ is even and $\lambda$ is integral, [Shi11] attaches to each $\pi\in\mathcal{AR}^{\mathrm{symp}}_{N}(\mathfrak{n},\lambda)$ a Galois representation of conductor $\mathfrak{n}$, Hodge-Tate weights matching $\lambda$, and the same root number. Let $\overline{\Phi}^{\mathrm{symp}}_{N}(E/F,\mathfrak{n},\lambda)$ be the resulting set of conjugate symplectic-type Galois representations $\Gamma_{F}\to\mathrm{GL}_{N}$.
We deduce from our main theorems an equidistribution result for the family $\overline{\Phi}^{\mathrm{symp}}_{N}(E/F,\mathfrak{n},\lambda)$:
Corollary 1.3.1.
Let $N$ be even and assume that $\mathfrak{n}$ is divisible by a high enough power of a ramified prime. Then
$\lim_{\begin{subarray}{c}\lambda\to\infty\\ \lambda\text{ integral}\end{subarray}}\frac{1}{|\overline{\Phi}^{\mathrm{symp}}_{N}(E/F,\mathfrak{n},\lambda)|}\sum_{\varphi\in\overline{\Phi}^{\mathrm{symp}}_{N}(E/F,\mathfrak{n},\lambda)}\epsilon(1/2,\varphi)\to 0.$
Assuming the Langlands correspondence, we expect $\overline{\Phi}^{\mathrm{symp}}_{N}(E/F,\mathfrak{n},\lambda)$ to be exactly the set of irreducible Galois representations of specific Hodge-Tate weight and conductor. This gives:
Corollary 1.3.2.
Let $N$ be even and assume the Langlands correspondence for conjugate symplectic-type automorphic representations of $\mathrm{GL}_{N}/E$ with regular integral infinitesimal character at infinity.
Then, if $\Phi_{N}^{\mathrm{symp}}(E/F,\mathfrak{n},\lambda)$ is the set of irreducible $N$-dimensional Galois representations that are
  • conjugate symplectic,
  • conductor $\mathfrak{n}$ coprime to $2$ divisible by a high enough power of a ramified prime,
  • Hodge-Tate weights matching $\lambda$,
we have,
$\lim_{\begin{subarray}{c}\lambda\to\infty\\ \lambda\text{ integral}\end{subarray}}\frac{1}{|\Phi^{\mathrm{symp}}_{N}(E/F,\mathfrak{n},\lambda)|}\sum_{\varphi\in\Phi^{\mathrm{symp}}_{N}(E/F,\mathfrak{n},\lambda)}\epsilon(1/2,\varphi)\to 0$
Known analogous self-dual results [KS20], [KS23] do not yet cover the symplectic-type case—in particular, the weak matching of [Shi24] does not yet give matching of root numbers. Therefore we only get a Galois representation corollary conditional on the Langlands correspondence:
Corollary 1.3.3.
Let $N$ be even and assume the Langlands correspondence for symplectic-type automorphic representations of $\mathrm{GL}_{N}/F$ with regular integral infinitesimal character at infinity.
Let $\Phi_{N}^{\mathrm{symp}}(F,\mathfrak{n},\lambda)$ be the set of irreducible symplectic Galois representations $\operatorname{Gal}(\bar{F}/F)\to\mathrm{Sp}_{N}$ that have
  • conductor $\mathfrak{n}$ divisible by a high enough power of some prime,
  • Hodge-Tate weights matching $\lambda$,
Then
$\lim_{\begin{subarray}{c}\lambda\to\infty\\ \lambda\text{ integral}\end{subarray}}\frac{1}{|\Phi^{\mathrm{symp}}_{N}(F,\mathfrak{n},\lambda)|}\sum_{\varphi\in\Phi^{\mathrm{symp}}_{N}(F,\mathfrak{n},\lambda)}\epsilon(1/2,\varphi)\to 0.$
if and only if there is finite place $v$ such that $v(\mathfrak{n})>N$ or $v(\mathfrak{n})$ is odd.
Similar to Theorem 1.2.1, we can weight the averages by functions of Frobenius conjugacy classes at places coprime to $\mathfrak{n}$ and we also have the same control over rates of convergence—see again Theorems 10.2.1 and 10.3.2 for full generality.
1.4. Summary of Argument
Once the outline of the argument in §1.4.2 is understood well, the details of the individual steps are reasonably self-contained and can more-or-less be read in any order.
1.4.1. $\mathrm{GL}_{2}$-model
Our method is very closely modeled on the classical $\mathrm{GL}_{2}/\mathbb{Q}$-argument which we now present in modernized language to emphasize the parallelism.
Let $\mathcal{AR}_{\mathrm{cusp}}(\mathrm{GL}_{2}/\mathbb{Q},\lambda)$ be the set of automorphic representations for $\mathrm{GL}_{2}/\mathbb{Q}$ with regular, integral infinitesimal character $\lambda$ at infinity. We are interested in the asymptotic fractions of $\pi\in\mathcal{AR}_{\mathrm{cusp}}(\mathrm{GL}_{2}/\mathbb{Q},\lambda)$ with conductor $c(\pi)=n$ and $\epsilon(1/2,\pi)=\pm 1$—in other words, in the variation of the ratio
(1.4.1) $\left.\sum_{\begin{subarray}{c}\pi\in\mathcal{AR}_{\mathrm{cusp}}(\mathrm{GL}_{2}/\mathbb{Q},\lambda)\\ c(\pi)=n\end{subarray}}\epsilon(1/2,\pi)\right/\sum_{\begin{subarray}{c}\pi\in\mathcal{AR}_{\mathrm{cusp}}(\mathrm{GL}_{2}/\mathbb{Q},\lambda)\\ c(\pi)=n\end{subarray}}1.$
as $\lambda\to\infty$.
Step 1:
Following the theory of newvectors, there exists a sequence of compact-open subgroups $\Gamma_{0}(n)\subset GL_{2}(\mathbb{A}^{\infty})$ such that
$(\pi^{\infty})^{\Gamma_{0}(n)}=\begin{cases}1&c(\pi)=n\\ 0&c(\pi)>n.\end{cases}$
If $c(\pi)=n$, we call an appropriately normalized element of $(\pi^{\infty})^{\Gamma_{0}(n)}$ the newvector $\varphi$ for $\pi$. The newvector satisfies a key property of being a “test vector” for Rankin-Selberg integral representations.
Step 2:
The conductor $n$ has an associated Atkin-Lehner involution $\iota_{n}$. As a corollary of the proof of the functional equation of the $L$-function $L(s,\pi)$ through its $\mathrm{GL}_{2}\times\mathrm{GL}_{1}$-Rankin-Selberg integral representation (e.g. [DS05, 5.10.2]), we get
$\varphi^{\iota_{n}}=\epsilon^{\infty}(1/2,\pi)\varphi$
(leaving out an archimedean factor). The “test vector” property of the newvector $\varphi$ is a crucial input here.
Step 3:
We can find $g_{n}\in\mathrm{GL}_{2}(\mathbb{A}^{\infty})$ such that $f^{\iota_{n}}=g_{n}f$ for all $f$ in the space of automorphic forms. By taking the trace of automorphic representations $\pi$ against the volume-normalized indicator functions of the corresponding cosets, we get:
$\operatorname{tr}_{\pi^{\infty}}\bar{\mathbf{1}}_{\Gamma_{0}(n)}=\begin{cases}0&c(\pi)>n\\ 1&c(\pi)=n\end{cases},\qquad\operatorname{tr}_{\pi^{\infty}}\bar{\mathbf{1}}_{g_{n}\Gamma_{0}(n)}=\begin{cases}0&c(\pi)>n\\ \epsilon^{\infty}(1/2,\pi)&c(\pi)=n\end{cases},$
where we leave out factors at the archimedean place.
Step 4:
Using an explicit description of the space of oldforms—the $\Gamma_{0}(n^{\prime})$-fixed points in $\pi^{\infty}$ when $c(\pi)=n$ and $n|n^{\prime}$—we compute
$\operatorname{tr}_{\pi^{\infty}}\bar{\mathbf{1}}_{\Gamma_{0}(n^{\prime})}=\dim((\pi^{\infty})^{\Gamma_{0}(n^{\prime})}),\qquad\operatorname{tr}_{\pi^{\infty}}\bar{\mathbf{1}}_{g_{n}\Gamma_{0}(n^{\prime})}.$
The values depend only on $n^{\prime}/n$. This allows us to use linear combination of functions of the type $\bar{\mathbf{1}}_{\Gamma_{0}(n)}$ or $\bar{\mathbf{1}}_{g_{n}\Gamma_{0}(n)}$ to construct functions $E^{\infty}_{n}$ and $C^{\infty}_{n}$ respectively so that
$\operatorname{tr}_{\pi^{\infty}}C^{\infty}_{n}=\begin{cases}1&c(\pi)=n\\ 0&c(\pi)\neq n\\ \end{cases},\qquad\operatorname{tr}_{\pi^{\infty}}E^{\infty}_{n}=\begin{cases}\epsilon^{\infty}(1/2,\pi)&c(\pi)=n\\ 0&c(\pi)\neq n\\ \end{cases}.$
Since $\epsilon_{\infty}(\pi_{\infty},1/2)=:\epsilon_{\infty}(1/2,\lambda)$ is constant for $\pi_{\infty}$ with fixed infinitesimal character, we can write the quantity (1.4.1) which we are trying to estimate as the following ratio of traces:
$\epsilon_{\infty}(1/2,\lambda)\left.\sum_{\pi\in\mathcal{AR}_{\mathrm{cusp}}(\mathrm{GL}_{2}/\mathbb{Q},\lambda)}\operatorname{tr}_{\pi^{\infty}}(E^{\infty}_{n})\right/\sum_{\pi\in\mathcal{AR}_{\mathrm{cusp}}(\mathrm{GL}_{2}/\mathbb{Q},\lambda)}\operatorname{tr}_{\pi^{\infty}}(C^{\infty}_{n}).$
Step 5:
The Eichler-Selberg trace formula gives a tractable expression for
$\sum_{\pi\in\mathcal{AR}_{\mathrm{cusp}}(\mathrm{GL}_{2}/\mathbb{Q},\lambda)}\operatorname{tr}_{\pi^{\infty}}(\psi)$
for any test function $\psi$. Using this expression, we show that if $F_{\lambda}$ is the finite-dimensional representation of $GL_{2}$ with infinitesimal character $\lambda$, then
$(\dim F_{\lambda})^{-1}\sum_{\pi\in\mathcal{AR}_{\mathrm{cusp}}(\mathrm{GL}_{2}/\mathbb{Q},\lambda)}\operatorname{tr}_{\pi^{\infty}}(\psi)\to\psi(1)$
as $\lambda\to\infty$. In particular the expression (1.4.1) has asymptotics:
[…]
Step 6:
We compute that $C_{n}(1)>0$. Furthermore, $E_{n}(1)=0$ if and only if $n$ is not a product of squares of distinct primes. This shows that we have asymptotic equidistribution of $\epsilon(1/2,\pi)$ if and only if $n$ is not product of squares of distinct primes.
1.4.2. Full Argument
We now summarize the general $\mathrm{GL}_{N}$-argument and point out where complications occur. We are interested in the analogous
(1.4.2) $\left.\sum_{\begin{subarray}{c}\pi\in\mathcal{AR}^{\mathrm{symp}}_{\mathrm{cusp}}(\mathrm{GL}_{N}/F,\lambda)\\ c(\pi)=\mathfrak{n}\end{subarray}}\epsilon(1/2,\pi)\right/\sum_{\begin{subarray}{c}\pi\in\mathcal{AR}^{\mathrm{symp}}_{\mathrm{cusp}}(\mathrm{GL}_{N}/F,\lambda)\\ c(\pi)=\mathfrak{n}\end{subarray}}1.$
for totally real $F$ in the self-dual case and
(1.4.3) $\left.\sum_{\begin{subarray}{c}\pi\in\mathcal{AR}^{\mathrm{symp}}_{\mathrm{cusp}}(\operatorname{Res}^{E}_{F}\mathrm{GL}_{N},\lambda)\\ c(\pi)=\mathfrak{n}\end{subarray}}\epsilon(1/2,\pi)\right/\sum_{\begin{subarray}{c}\pi\in\mathcal{AR}^{\mathrm{symp}}_{\mathrm{cusp}}(\operatorname{Res}^{E}_{F}\mathrm{GL}_{N},\lambda)\\ c(\pi)=\mathfrak{n}\end{subarray}}1.$
for $E/F$ CM in the conjugate self-dual case. Note here that over a general number field, the level becomes an ideal $\mathfrak{n}$ of $F$ instead of an integer $n$.
Step 1:
Newvector theory directly generalizes to $\mathrm{GL}_{N}(\mathbb{A}_{F})$ or $\mathrm{GL}_{N}(\mathbb{A}_{E})$ through [JPSS81]. The corresponding level subgroups are the “mirahorics” $K_{1}(\mathfrak{n})$. We summarize the results and the necessary “test vector” properties in §§3.1, 3.2.
Step 2:
Using ideas from [Tom21], we generalize the Atkin-Lehner involution to define involutions  $\iota_{\mathfrak{n}},\bar{\iota}_{\mathfrak{n}}$ in the self-dual or conjugate self-dual cases respectively. To connect this to root numbers, it turns out to be best to look at $\mathrm{GL}_{N}\times\mathrm{GL}_{N-1}$ Rankin-Selberg $L$-functions instead of those associated to $\mathrm{GL}_{N}\times\mathrm{GL}_{1}$. The proof of the functional equation (e.g. [Cog04, §6]) can be algebraically manipulated to involve either $\iota_{\mathfrak{n}}$ or $\bar{\iota}_{\mathfrak{n}}$ instead of the standard involution $g\mapsto g^{-T}$. This roughly shows that for newvectors $\varphi$ in (conjugate) self-dual representations,
(1.4.4) $\varphi^{\iota_{\mathfrak{n}}}\approx\epsilon^{\infty}(1/2,\pi)^{N-1}\varphi.$
After reviewing some background in §3.3, this is work of §§3.4,3.5 culminating in local Theorems 3.4.2 and 3.5.2.
The exponent in (1.4.4) is the reason for our restriction to even $N$ in the conjugate self-dual result. Luckily, there are no odd symplectic-type representations in the self-dual case.
Step 3:
This is where we make our main departure from the classical argument. Since $\iota_{\mathfrak{n}},\bar{\iota}_{\mathfrak{n}}$ are inner to $g\mapsto g^{-T}$ and $g\mapsto\bar{g}^{-T}$ which is outer when $N>2$, the element $g_{\mathfrak{n}}$ such that
$g_{\mathfrak{n}}\varphi=\varphi^{\iota_{\mathfrak{n}}}$
no longer lies $\mathrm{GL}_{N}(\mathbb{A}_{F})$ itself, but instead in a larger twisted group. Consider involutions $\theta,\bar{\theta}$ that are inner to $g\mapsto g^{-T}$, $g\mapsto\bar{g}^{-T}$ respectively. Then $g_{\mathfrak{n}}$ lies in one of the cosets $\widetilde{G}_{N}:=\mathrm{GL}_{N}(\mathbb{A}_{F})\rtimes\theta$ or $\mathrm{GL}_{N}(\mathbb{A}_{E})\rtimes\bar{\theta}$ in a corresponding semidirect product to which the action of $\mathrm{GL}_{N}$ on $\pi$ can be extended. See §4.1 for details.
Therefore, the generalized versions of the building blocks $\bar{\mathbf{1}}_{g_{n}K_{1}(n)}$ for $E_{n}$ lie on the twisted group $\widetilde{G}_{N}$ instead. This is a feature and not a bug since it allows us to use the twisted trace formula and isolate the (conjugate) self-dual spectrum for $\mathrm{GL}_{N}$ as desired.
On the other hand, generalizing the building blocks $\bar{\mathbf{1}}_{K_{1}(n)}$ of $C_{n}$ is much more difficult since we do not have a twisted analogue. We will therefore need to construct a generalized $C_{n}$ in another way.
Step 4:
A result of Reeder [Ree91] gives a description of oldform spaces $(\pi^{\infty})^{K_{1}(\mathfrak{n}^{\prime})}$ for $c(\pi)|\mathfrak{n}^{\prime}$ in terms of Hecke operators for $\mathrm{GL}_{N-1}$. The action of these $\mathrm{GL}_{N-1}$-Hecke operators can be commuted with the $\mathfrak{\iota}_{\mathfrak{n}}$’s to give a formula for $\operatorname{tr}_{\pi^{\infty}}(\mathbf{1}_{g_{\mathfrak{n}^{\prime}}K_{1}(\mathfrak{n}^{\prime})})$ that only depends on $\mathfrak{n}^{\prime}/c(\pi)$ (from local Propositions 4.2.3, 4.2.5).
In particular, we can similarly solve a system of linear equations and find a linear combination $\widetilde{E}^{\infty}_{\mathfrak{n}}$ of the twisted test functions $\bar{\mathbf{1}}_{g_{\mathfrak{n}}K_{1}(\mathfrak{n})}$ on $\widetilde{G}_{N}$ so that roughly:
$\operatorname{tr}_{\pi^{\infty}}(\widetilde{E}^{\infty}_{\mathfrak{n}})\approx\begin{cases}\epsilon^{\infty}(1/2,\pi)&c(\pi)=\mathfrak{n}\\ 0&c(\pi)\neq\mathfrak{n}\end{cases}$
This is the work of §§4.2, 4.3 culminating in Theorems 4.3.5 and 4.3.6.
The twisted generalization of $C^{\infty}_{n}$ requires a completely different method in §4.4. We first show that root numbers are constant on Bernstein components of $\mathrm{GL}_{N}$ intersect the set of (conjugate) self-dual representations of a fixed conductor (Propositions 4.4.2, 4.4.3). Therefore, an abstract Paley-Wiener theorem [Rog81] lets us bootstrap the existence of $\widetilde{E}^{\infty}_{\mathfrak{n}}$ into the existence of a $\widetilde{C}^{\infty}_{\mathfrak{n}}$ such that
$\operatorname{tr}_{\pi^{\infty}}(\widetilde{C}^{\infty}_{\mathfrak{n}})=\begin{cases}1&c(\pi)=\mathfrak{n}\\ 0&c(\pi)\neq\mathfrak{n}\end{cases}$
(Corollary 4.4.4).
Step 5:
The techniques of [DGG22] using the inductive argument of [Taï17] modify in a straightforward way to generate a tractable expression for both
$\sum_{\pi\in\mathcal{AR}_{\mathrm{cusp}}^{\mathrm{symp}}(\mathrm{GL}_{N}/F,\lambda)}\operatorname{tr}_{\pi^{\infty}}(\widetilde{\psi})$
and
$\sum_{\pi\in\mathcal{AR}_{\mathrm{cusp}}^{\mathrm{symp}}(\operatorname{Res}^{E}_{F}\mathrm{GL}_{N},\lambda)}\operatorname{tr}_{\pi^{\infty}}(\widetilde{\psi}),$
for suitable twisted test functions $\widetilde{\psi}$ (Proposition 6.1.1 and §5).
After inputting bounds from [ST16], we obtain that both these expressions have (up to technical issues with centers) have scaling limits that are
$\approx\widetilde{\psi}^{G}(1),$
for $\widetilde{\psi}^{G}$ an endoscopic transfer to the twisted endoscopic group $G=\mathrm{SO}_{N+1}$ of $\widetilde{G}_{N}$ in the self-dual case and $G=U_{N}$ of $\widetilde{G}_{N}$ embedded through base change in the conjugate self-dual case (§6 culminating in Theorem 6.4.1).
In total, we get that, up to center technicalities:
[…]
as $\lambda\to\infty$. Note that [ST16] also gives us a power saving on the error term.
Step 6:
It remains to understand $(\widetilde{E}^{\infty}_{\mathfrak{n}})^{G}(1)$ and $(\widetilde{C}^{\infty}_{\mathfrak{n}})^{G}(1)$. Unlike the $\mathrm{GL}_{2}$ case, this is by far the hardest part of the argument.
Our explicit description for $\widetilde{E}^{\infty}_{\mathfrak{n}}$ allows us to use the orbital integral transfer identities to compute $(\widetilde{E}^{\infty}_{\mathfrak{n}})^{G}(1)$. The identity class of our relevant $G$ is a twisted analogue of the “equisingular” classes studied in [LS07, 2.4.A]. Therefore in §7.1, we recall a Shalika germ argument from Mœglin and Waldspurger’s stabilization of the twisted trace formula [MW16] to get that $(\widetilde{E}^{\infty}_{\mathfrak{n}})^{G}(1)$ equals a particular non-regular twisted $\kappa$-orbital integral on $\widetilde{G}_{N}$ (Corollary 7.1.8). Computing when this integral vanishes is the work of §7.2,8 and is somewhat involved in the conjugate self-dual case.
The function $\widetilde{C}^{\infty}_{\mathfrak{n}}$ is much more daunting since it only exists by an abstract Paley-Wiener theorem. Luckily, we only need to show that $(\widetilde{C}^{\infty}_{\mathfrak{n}})^{G}(1)>0$ (up to details with the center). By the endoscopic character identities and the spectral definition of $\widetilde{C}^{\infty}_{\mathfrak{n}}$, we can show that a formally defined “stable Fourier transform” of $(\widetilde{C}^{\infty}_{\mathfrak{n}})^{G}$ is always positive and somewhere non-vanishing. This allows us to apply a positivity criterion (Proposition 9.2.2) that we prove from a “stable Fourier inversion” result (Theorem 9.1.2).
We obtain this stable Fourier inversion result by showing that the Plancherel measure for each $G_{v}$ descends to a formally-defined “stable unitary dual” in a key Proposition 9.1.1. We first interpret the Plancherel measure of a local test function $\widehat{\psi}_{v}$ on the unitary dual as an asymptotic count of global automorphic representations. By ideas from [DGG22], this is then shown to match a stable trace formula asymptotic which necessarily only depends on the stable Fourier transform of $\psi_{v}$. This is the work of §9.1. We conclude the positivity statement in §9.2.
That the stable Fourier transform is somewhere non-vanishing requires number-theoretic input: the existence of local (conjugate) self-dual representations with desired invariants—type, conductor, central character, etc. This is done in §9.3 and is again quite involved in the conjugate self-dual case. Our final statements for $\widetilde{C}^{\infty}_{\mathfrak{n}}$ are in §9.4.
1.5. Conditionality
Since we need to separate (conjugate) symplectic from (conjugate) orthogonal global representations with the trace formula, this result is conditional on the endoscopic classification—[Art13] for the self-dual case and [Mok15] for the conjugate self-dual. These classifications depend on the unpublished weighted, twisted fundamental lemma, three unpublished references A25,26,27 in [Art13], and their unpublished analogues for [Mok15].
1.6. Acknowledgments
The idea for this paper came from watching talks on the IAS YouTube channel about murmurations of families of elliptic curves and how they could be studied with the trace formula—we would like to thank the IAS for making them publicly available. Yiannis Sakellaridis pointed out how higher-rank Rankin-Selberg integrals would be a good way to generalize the classical $\mathrm{GL}_{2}$ argument relating root numbers to Atkin-Lehner involutions. We would also like to thank Alexander Bertoloni-Meli, Peter Dillery, Peter Humphries, Ashwin Iyengar, Arno Kret, Huajie Li, Kimball Martin, Masao Oi, David Schwein, Sug Woo Shin, Joel Specter, and Sandeep Varma for helpful conversations.
The first author was supported by NSF postdoctoral grant 2103149 while working on this project.
1.7. Notation
1.7.1. Groups Considered
Throughout this paper we will be looking at two cases of involution/self-duality on $\mathrm{GL}_{N}$. Summarizing our parallel notational conventions for each case:
The self-dual or orthogonal/symplectic case where we consider:
  • $G_{N}=\mathrm{GL}_{N}$ over a totally real number field $F$ (so that for $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$, the group $G_{\infty}$ has discrete series)
  • Involution $\iota:g\mapsto g^{-T}$
  • Involution $\theta=\operatorname{Ad}_{w_{N}}\circ\iota$ inner to $\iota$ fixing the standard pinning
  • Local groups $G_{N,v}=\mathrm{GL}_{N}(F_{v})$
  • Local conductors $\mathfrak{p}_{v}^{i}$ for non-negative integer $i$
  • Global conductors $\mathfrak{n}$ that are ideals of $\mathcal{O}_{F}$
  • Twisted group $\widetilde{G}_{N}=G_{N}\rtimes\theta$
  • Simple Endoscopic groups $G\in\mathcal{E}_{\mathrm{sim}}(R)$ that are either $\mathrm{SO}_{k}$ or $\mathrm{Sp}_{2k}$’s
The conjugate self-dual or unitary case where we consider:
  • $G_{N}=\operatorname{Res}^{E}_{F}G_{N}$ for a CM quadratic extension $E/F$ of number fields (so that for $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$, $G_{\infty}$ has discrete series)
  • Involution $\bar{\iota}:g\mapsto\bar{g}^{-T}$
  • Involution $\bar{\theta}=\operatorname{Ad}_{w_{N}}\circ\bar{\iota}$ inner to $\bar{\iota}$ fixing the standard pinning
  • Local groups $G_{N,v}=\mathrm{GL}_{N}(F_{v}\otimes_{F}E)$
  • Local conductors $\mathfrak{p}_{v}^{i}$ for $i$ either
    • A pair of non-negative integers if $v$ is split
    • A non-negative integer if $v$ is inert
    • A non-negative half-integer if $v$ is ramified. When speaking locally, we will usually instead index by the exponent $2i$ of $\mathfrak{p}_{w}^{2i}=\mathfrak{p}_{v}^{i}$ for $w$ lying over $v$.
  • Global conductors $\mathfrak{n}$ that are finite products of local levels at different places
  • Twisted group $\widetilde{G}_{N}=G_{N}\rtimes\bar{\theta}$
  • Simple Endoscopic groups $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ that are quasisplit $U_{N}$’s
1.7.2. Other
Basics:
  • $F/\mathbb{Q}$ a totally real number field, $E/F$ an imaginary quadratic extension
  • places of $F$ are denoted $v$, and $q_{S}$ the product of residue field degrees over a finite set $S$ of finite places of $F$
  • $\Gamma_{F}$ the absolute Galois group of $F$
  • $\star^{S}$ and $\star_{S}$ are components at $S$ and away from $S$ of structure $\star$ for $S$ some set of places
  • $\mathrm{spl},\mathrm{ns}$ the sets of finite places of $F$ that are split, non-split in $E$ respectively
  • $v_{1}\boxtimes v_{2}$ is the corresponding representation of $H_{1}\times H_{2}$ if $v_{i}$ is a representation of $H_{i}$
  • $[d]$ is the $d$-dimesional irrep of $\mathrm{SL}_{2}$
  • $w_{N}$ is the antidiagonal matrix of $\mathrm{GL}_{N}$ with alternating signs so that $g\mapsto w_{N}g^{-T}w_{N}$ fixes the standard pinning.
  • $\mathbf{1}_{X}$ is the indicator function of the set $X$ and $\bar{\mathbf{1}}_{X}=\operatorname{vol}(X)^{-1}\mathbf{1}_{X}$
  • $\tau_{v}(A)$ is the eigenvalue of the operator $A$ on the vector $v$
  • $\omega_{\pi}$ is the central character of representation $\pi$
Groups
  • $Z_{G}$ is the center of $G$
  • $G_{\gamma}$ is the centralizer of $\gamma$ in $G$
  • $G^{0}$ is the connected component of the identity
  • $G^{\mathrm{der}}$ is the derived subgroup
  • $G_{\mathrm{ad}}$ is the adjoint quotient
  • $G^{\mathrm{sc}}$ is the simply-connected cover of $G^{\mathrm{der}}$
  • $P_{G}$ is the number of positive roots of $G$
  • $\widehat{G}$ is the complex dual group
  • ${}^{L}\!{G}=\widehat{G}\rtimes\Gamma_{F}$ is the Langlands dual
  • $\check{G}_{v},\check{G}_{S}$ are the unitary duals (with Fell topology and Plancherel measure)
  • $\check{G}_{v}^{\mathrm{temp}},\check{G}_{S}^{\mathrm{temp}}$ are the tempered unitary duals
  • $\widehat{f}$ is the Fourier transform of function $f$ on $G_{v}$ or $G_{S}$. Beware that this is modified from the usual definition at non-tempered $\pi$; see §9.1.1.
With respect to either the self-dual or conjugate self-dual endoscopic classification:
  • ~$\widetilde{\mathcal{E}}_{\mathrm{ell}}(N),\mathcal{E}_{\mathrm{ell}}(G)$ are the elliptic endoscopic groups of $\widetilde{G}_{N},G$ respectively as in §2.2.2
  • ~$\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ are the simple endoscopic groups of $\widetilde{G}_{N}$ as in §1.7.1
  • $f^{H}$ for $f$ a test function on reductive $G/F$ is a choice of endoscopic transfer to some $H\in\mathcal{E}_{\mathrm{ell}}(G)$
  • $f^{N}$ for $f$ a test function on $G\in\widetilde{\mathcal{E}}_{\mathrm{ell}}(N)$ is a test function on $\widetilde{G}_{N}$ that transfers to $f$
  • $\psi=\oplus\tau_{i}[d_{i}]$ is an Arthur parameter with cuspidal components $\tau_{i}$
  • $\Psi(N),\Psi(G)$ are the sets of parameters associated to $\widetilde{G}_{N}$ and $G$ respectively (or for $G=\mathrm{SO}_{2N}^{\eta}$, outer automorphism orbits of such)
  • The “Arthur-$\mathrm{SL}_{2}$” of a parameter is an unordered partition $Q$ representing its restriction to the Arthur-$\mathrm{SL}_{2}$
  • $\pi_{\psi}$ is the automorphic representation of $G_{N}$ corresponding to $\psi$
  • $\tilde{\pi}_{\psi}$ is the extension of $\pi_{\psi}$ to $\widetilde{G}_{N}$
  • $\varphi_{\psi}$ is the $L$-parameter associated to the $A$-parameter $\psi$
  • $\mathcal{S}_{\psi},\mathcal{S}_{\psi_{v}}$ are Arthur’s component groups associated to global or local parameters
  • $s_{\psi},s_{\psi_{v}}$ are the special elements identified in these component groups
  • $\epsilon_{\psi}$ is the identified character on global $\mathcal{S}_{\psi}$
  • $\Pi_{\psi}(G),\Pi_{\psi_{v}}(G_{v})$ are the $A$-packets associated to global or local $A$-parameters on group $G,G_{v}$ (or for $G=\mathrm{SO}_{2N}^{\eta}$, outer automorphism orbits of such).
  • $\eta^{\psi}_{\pi}$ is the character of $\mathcal{S}_{\psi}$ associated to $\pi\in\Pi_{\psi}$
  • $\operatorname{tr}_{\psi_{v}}:=\operatorname{tr}^{G_{v}}_{\psi_{v}}$ is the stable packet trace for parameter $\psi_{v}$ on group $G_{v}$
Shapes
  • $\Delta=((T_{i},d_{i},\eta_{i},\lambda_{i}))_{i}$ is a refined shape as in Definition 5.3.2.
  • $\Box=((T_{i},d_{i},\eta_{i}))_{i}$ is an unrefined shape as in §5.5
  • $\Box(\Delta)$ is the unrefined shape from forgetting information in $\Delta$
  • $\Sigma_{\eta,\lambda},\Sigma_{\eta}=\Box(\Sigma_{\eta,\lambda})$ are the simple (refined) shapes as in Definition 5.3.4
  • $G(\Delta),G(\Box)$ are the groups associated to (refined) shapes $\Delta$ or $\Box$ as in Proposition 5.3.5
  • $G_{F}(\Box)$ is the reduced group associated to shape $\Box$ as in §5.5
  • $\Box_{\lambda}$ is a set of refined shapes $\Delta$ with total infinitesimal character $\lambda$ associated to unrefined shape $\Box$ as in §5.5
  • $\dim\lambda$ is the dimension of the finite-dimensional representation with regular, integral infinitesimal character $\lambda$
  • $m(\lambda)$ is a norm of the infinitesimal character $\lambda$ from §5.6
  • $\dim_{\Box}\lambda$ is a modified dimension based on the unrefined shape $\Box$ from §5.6
Trace Formula Decompositions:
  • $\mathcal{AR}_{\mathrm{disc}}(G)$, $\mathcal{AR}_{\mathrm{cusp}}(G)$ is the set of discrete (resp. cuspidal) automorphic representations of $G/F$
  • $I^{G},S^{G}$ are Arthur’s invariant and stable trace formulas for $G$
  • $I_{\mathrm{disc}}^{G},S_{\mathrm{disc}}^{G}$ are their discrete parts
  • $R^{G}_{\mathrm{disc}}$ is the trace against $\mathcal{AR}_{\mathrm{disc}}(G)$
  • $I^{G}_{\psi},S^{G}_{\psi}$ are the summands of $I^{G}_{\mathrm{disc}},S^{G}_{\mathrm{disc}}$ associated to parameter $\psi$
  • $I^{G}_{\Delta},S^{G}_{\Delta},I^{G}_{\Box},S^{G}_{\Box}$ are the summands of $I^{G}_{\mathrm{disc}},S^{G}_{\mathrm{disc}}$ associated to the refined shape $\Delta$ or shape $\Box$
  • $\star^{N}$ is the version of any of the variants above associated to $\widetilde{G}_{N}$
  • $\mathrm{EP}_{\lambda}$ is the (endoscopically normalized) Euler-Poincaré function for regular integral infinitesimal character $\lambda$ as in §6.1
Conductors:
  • $W(\psi)$ is the space of Whittaker functions on $\mathrm{GL}_{N}(F_{v})$ for character $\psi$
  • $W_{\varphi}:=W^{\psi}_{\varphi}$ is the Whittaker function for vector $\varphi$ in representation $\pi$.
  • $K_{v}$ is the hyperspecial subgroup of the $p$-adic group $G_{v}$
  • $K_{1}(\mathfrak{n})$ is the “mirahoric” level-subgroup for conductor $\mathfrak{n}$ from 3.1.1
  • $K_{1,v}(k):=K_{1}(\mathfrak{p}_{v}^{k})$ is the local factor at $v$ of $K_{1}(\mathfrak{n})$
  • $c(\pi)$ is the conductor of global representation $\pi$.
  • $w$ is place of $E$ lying over a non-split place $v$ of $F$
  • $c(\pi_{v})$ is the $k$ such that local $\pi_{v}$ has conductor $\mathfrak{p}_{v}^{k}$ in the self-dual case and $\mathfrak{p}_{w}^{k}$ in the conjugate self-dual case (beware that this local indexing is twice the global indexing for ramified $v$ in the conjugate self-dual case)
  • $e_{v}$ is the ramification index of $v$ in $E$
  • $\mathfrak{D}_{E/F}$ is the different ideal of $E/F$
  • $j_{v}$ is a “depth-of-ramification” invariant of $v$ from §9.3 in the conjugate self-dual case
  • $b_{v}$ is an invariant of $v$ depending on roots of unity defined in Lemma 9.4.2 in the conjugate self-dual case.
  • $\widetilde{E}^{\infty}_{v,k},\widetilde{E}^{\infty}_{\mathfrak{n}}$ are twisted test functions counting representations weighted by root number defined in §4.3
  • $\widetilde{C}^{\infty}_{v,k},\widetilde{C}^{\infty}_{\mathfrak{n}}$ are twisted test functions counting representations without weighting defined in §4.4
  • $\tilde{f}_{v,k}$, $\tilde{f}_{\mathfrak{n}}$ are indicators of cosets defined in §7.2 and §8.3
  • $\mathcal{AR}_{N}^{\star}(\lambda,\mathfrak{n})$ is the set of (conjugate) $\star$-type cuspidal automorphic representations on $\mathrm{GL}_{N}$ with infinitesimal character $\lambda$ at infinity and conductor $\mathfrak{n}$ for $\star=\mathrm{symp}$ or $\mathrm{orth}$.
2. Self-Dual and Conjugate Self-Dual Representations
We will consider, in parallel, self-dual and conjugate self-dual automorphic representations of $GL_{N}$, and their local constituents.
2.1. Local Notions
Let $F$ be a local field and $\pi$ be a unitary irreducible representation of $GL_{N}(F)$. The dual $\pi^{\vee}$ of $\pi$ is the representation acting on the same space, but such that
$\pi^{\vee}(g)=\pi(g^{-T}).$
2.1.1. Self-dual representations
We will say that $\pi$ is self-dual if $\pi\simeq\pi^{\vee}$. In this case, $\pi$ corresponds under the local Langlands correspondence [HT99, Hen00, Sch13] to an $N$-dimensional representation $\rho$ of the Local Langlands group $L_{F}$, i.e. a representation such that
$\rho(\sigma)=\rho(\sigma)^{-T}\quad\sigma\in L_{F}.$
If $\rho$ is additionally irreducible, then it preserves a nondegenerate bilinear form $B$. We say that $\rho$ is orthogonal or symplectic depending on whether $B$ can be chosen to be. We correspondingly say that $\pi$ is orthogonal-type or symplectic-type. If $\pi$ is symplectic-type, then $N$ is necessarily even, and the central character of $\pi$ is trivial. When $\pi$ is orthogonal-type, its central character is a quadratic character $\omega_{\pi}:F^{\times}\to\pm 1$. Note that being orthogonal and symplectic aren’t mutually exclusive.
If $\rho$ is reducible but semisimple, we say it is orthogonal or symplectic if all its irreducible constituents are.
2.1.2. Conjugate self-dual representations
Let $E/F$ be a (possibly split) extension of local fields, with involution $x\mapsto\bar{x}$. We say that a representation $\pi$ of $GL_{N}(E)$ is conjugate self-dual if $\pi\simeq\bar{\pi}^{\vee}$ where
$\bar{\pi}^{\vee}(g)=\pi(\bar{g}^{-T}).$
Such a representation $\pi$ corresponds under the local Langlands correspondence to an $N$-dimensional conjugate self-dual representation $\rho$ of $\Gamma_{E}$. That is, we have $\rho^{-T}\simeq\rho^{c}$ where $\rho^{c}(\sigma):=\rho(w_{c}^{-1}\sigma w_{c})$ for $w_{c}\in L_{E}\setminus L_{F}$.
If $\rho$ is further irreducible, there is a bilinear form $B$ on $\rho$ such that for all $\sigma\in\Gamma_{E}$
$B(\rho^{c}(\sigma)x,\rho(\sigma)x)=B(x,y),\qquad B(x,y)=\eta B(y,\rho(w_{c}^{2})x)$
for $\eta\pm 1$. If we can choose $\eta=1$, then $\rho$ is conjugate-orthogonal and if we can choose $\eta=-1$, $\rho$ is conjugate-symplectic. Note that these conditions are not mutually exclusive. We accordingly say that $\pi$ is orthogonal-type or symplectic-type.
If $\rho$ is reducible but semisimple, we say it is conjugate orthogonal or conjugate symplectic if all its irreducible constituents are.
2.2. Twisted Groups and Endoscopy
2.2.1. Twisted Groups
Self-dual and conjugate self-dual representations of $\mathrm{GL}_{N}$ are best studied using twisted endoscopy. We first define the key twisted groups.
Over field $F$ or quadratic extension $E/F$, let
$G_{N}=\begin{cases}\mathrm{GL}_{N}&\text{self-dual case}\\ \operatorname{Res}^{E}_{F}\mathrm{GL}_{N}&\text{conjugate self-dual case.}\end{cases}$
In the self-dual case, consider the involution
$\iota:G_{N}:=\mathrm{GL}_{N}\to\mathrm{GL}_{N}:g\mapsto g^{-T}$
over the number field $F$. In the conjugate self-dual case, consider
$\bar{\iota}:G_{N}:=\operatorname{Res}^{E}_{F}\mathrm{GL}_{N}\to\operatorname{Res}^{E}_{F}\mathrm{GL}_{N}:g\mapsto\bar{g}^{-T},$
also defined over $F$.
Next, define the matrix
(2.2.1) $w_{N}:=\begin{pmatrix}&&&(-1)^{N-1}\\ &&\iddots&\\ &-1&&\\ 1&&&\end{pmatrix}$
so that conjugating $\iota,\bar{\iota}$ by $w_{N}$ gives involutions $\theta,\bar{\theta}$ of $\mathrm{GL}_{N}$ and $\operatorname{Res}^{E}_{F}\mathrm{GL}_{N}$ respectively that preserve the standard pinnings.
Note.
As computational reminders, $w_{N}^{-1}=(-1)^{N-1}w_{N}$ and the automorphism $A\mapsto w_{N}A^{T}w_{N}^{-1}$ reflects the entries of an $N\times N$ matrix $A$ along the anti-diagonal and then multiplies the $(i,j)$th entry by the “checkerboard pattern” $(-1)^{i+j}$.
Following [Art13] and [Mok15], define twisted groups
$\widetilde{G}_{N}=\begin{cases}\mathrm{GL}_{N}\rtimes\theta&\text{self-dual case}\\ \operatorname{Res}^{E}_{F}\mathrm{GL}_{N}\rtimes\bar{\theta}&\text{conjugate self-dual case.}\end{cases}$
The (conjugate) self-dual representations of $G_{N}(\mathbb{A})$ or $G_{N}(F_{v})$ are exactly the ones that extend to $\widetilde{G}_{N}$. Any such irrep has exactly two extensions, though in our cases we will choose one as canonical in section 4.1.
2.2.2. Twisted Endoscopic Groups
(Conjugate) self-dual automorphic representations can be studied through the twisted version of Arthur’s trace formula on $\widetilde{G}_{N}$ as in [Art13, §3] or [Mok15, §4]. The trace formula decomposes into terms for elliptic endoscopic groups of $G_{N}$. We will only need to worry about the simple endoscopic groups ~$\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$:
In the self-dual case, the simple twisted endoscopic groups $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ are ([Art13, §1.2]):
  • $N$ even: $\widehat{G}\cong\mathrm{Sp}_{N}$ embedded in one possible way with trivial Galois action. This corresponds to $G\cong\mathrm{SO}_{N+1}$.
  • $N$ even: $\widehat{G}\cong\mathrm{SO}_{N}$ with Galois action by $\operatorname{Out}(G)\cong\mathbb{Z}/2$ through the quadratic character $\eta$. There is a unique embedding for each $\eta$. This corresponds to $G\cong\mathrm{SO}^{\eta}_{N}$, a quasisplit orthogonal group that is not split if $\eta\neq 1$.
  • $N$ odd: $\widehat{G}\cong\mathrm{SO}_{N}$ with trivial Galois action. There is one possible embedding for each quadratic character $\eta$ of $\Gamma_{F}$ (equivalently, of $F^{\times}\backslash\mathbb{A}_{F}^{\times}$). This corresponds to $G\cong\mathrm{Sp}_{N-1}$. We denote the versions for different embeddings $\mathrm{Sp}^{\eta}_{N-1}$.
In the conjugate self dual case, they can be ([Mok15, §2.4.1]):
  • $\widehat{G}\cong\mathrm{GL}_{N}$ with $\operatorname{Gal}(E/F)$ acting through an automorphism inner to $g\mapsto g^{-T}$. This corresponds to $G\cong U^{E/F}_{N}$, the quasisplit unitary group. There are two possible embeddings ${}^{L}\!{G}\to{}^{L}\!{G}_{N}$, and we denote the endoscopic data associated to different embeddings by $U^{\pm}_{E/F}(N)$.
2.3. Global Notions
2.3.1. Self-Dual Case
Consider a number field $F$. A representation $\pi$ of $\mathrm{GL}_{N}(\mathbb{A}_{F})$ is self-dual if $\pi\cong\pi^{\vee}$. The central character $\omega_{\pi}$ of $\pi$ is then necessarily quadratic.
As one of the key outputs of the endoscopic classification [Art13], every self-dual cuspidal automorphic representation $\pi$ of $\mathrm{GL}_{N}/F$ is assigned a group $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$. If the dual group $\widehat{G}$ is symplectic, we call $\pi$ symplectic-type and if $\widehat{G}$ is orthogonal, we call $\pi$ orthogonal-type. As part of the result, whenever $\pi$ is orthogonal/symplectic, then all its local factors $\pi_{v}$ are too since they have parameters factoring through $G$.
Finally, [Art13] gives a criterion for distinguishing orthogonal vs. symplectic which can be taken as a more direct definition (though we will not use it):
Theorem 2.3.1 ([Art13, Thm 1.5.3(a)]).
Let $\pi$ be a self-dual automorphic representation of $\mathrm{GL}_{N}/F$. If $\omega_{\pi}$ is non-trivial, then the assigned $G$ is $\mathrm{SO}_{N}^{\omega_{\pi}}$ or $\mathrm{Sp}_{N-1}^{\omega_{\pi}}$ (through class field theory) and $\pi$ is therefore orthogonal.
Otherwise, $\pi$ is symplectic if an only if its alternating square $L$-function $L(s,\pi,\wedge^{2})$ has a pole at $s=1$ and orthogonal if and only if its symmetric square $L$-function $L(s,\pi,\operatorname{Sym}^{2})$ has a pole at $s=1$.
2.3.2. Conjugate Self-Dual Case
Consider a quadratic extension $E/F$. A representation $\pi$ of $\mathrm{GL}_{N}(\mathbb{A}_{E})$ is conjugate self-dual if $\pi\circ\sigma\cong\pi^{\vee}$ where $\sigma$ is the non-trivial element of $\operatorname{Gal}(E/F)$.
As one of the key outputs of the endoscopic classification [Mok15], every conjugate self-dual cuspidal automorphic representation $\pi$ of $\operatorname{Res}^{E}_{F}\mathrm{GL}_{N}$ is assigned a group $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$. If this group is $U_{N}^{(-1)^{N-1}}$, we call $\pi$ conjugate orthogonal. Otherwise, we call $\pi$ conjugate symplectic. As part of the result, whenever $\pi$ is conjugate orthogonal/symplectic, then all its local factors $\pi_{v}$ are too since they have parameters factoring through $G$.
Similar to the above, [Mok15] gives a criterion for distinguishing conjugate orthogonal vs. conjugate symplectic which can be taken as a more direct definition (which we again will not use):
Theorem 2.3.2 ([Mok15, Thm 2.5.4(a)]).
Let $\pi$ be conjugate self-dual automorphic representation of $\operatorname{Res}^{E}_{F}\mathrm{GL}_{N}$. Then $\pi$ is conjugate symplectic if an only if its Asai $L$-function $L(s,\pi,\mathrm{Asai}^{-})$ has a pole at $s=1$ and conjugate orthogonal if and only if its Asai $L$-function $L(s,\pi,\mathrm{Asai}^{+})$ has a pole at $s=1$.
2.3.3. Restrictions on $F$
Finally, to keep the “discreteness” at infinity analogy with $\mathrm{GL}_{2}/\mathbb{Q}$-modular forms, we will only study cases when the relevant $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ have discrete series at infinity. Therefore, we make the following assumption:
Assumption 2.3.3.
In the conjugate self-dual case assume that $F$ is totally real. In the conjugate self-dual case assume that $E/F$ is CM: a totally imaginary extension of a totally real field.
Outside of the $G=\mathrm{SO}_{2n}^{\eta}$ case, Assumption 2.3.3 implies that $G$ has discrete series at infinity. In the $G=\mathrm{SO}_{2n}^{\eta}$ case, Assumption 2.3.3 gives that $G_{\infty}$ has discrete series if and only if for each $v|\infty$, the following holds: $\eta|_{\operatorname{Gal}_{F_{v}}}$ is trivial if and only if $n$ is even. As we will see in Proposition 5.3.1, only these $G$ will be relevant.
3. New Vectors and Root Numbers
In this section, we exploit integral representations of $L$-functions to realize the root number of a symplectic automorphic $L$-function as the eigenvalue of a certain operator acting on the space of newforms. The main statements 3.4.4, 3.5.4 are local, and in the next section we will consider global applications.
3.1. New Vectors
3.1.1. Basic Version
We recall the theory of newvectors for local representation of $\mathrm{GL}_{N}$. A useful reference is [Hum20, §3], though results there are stated more generally than we need here.
Let $v$ be a place of our number field $F$ and $K_{v}=\mathrm{GL}_{N}(\mathcal{O}_{v})$.
Definition 3.1.1.
For each $i$, the mirahoric subgroup $K_{1}(\mathfrak{p}_{v}^{i})=K_{1,v}(i)\subseteq K_{v}$ is
$K_{1}(\mathfrak{p}_{v}^{i}):=K_{1,v}(i):=\left\{A\in K_{v}:A\cong\begin{pmatrix}*&\cdots&*&*\\ \vdots&\ddots&\vdots&\vdots\\ *&\cdots&*&*\\ 0&\cdots&0&1\end{pmatrix}\pmod{\mathfrak{p}_{v}^{i}}\right\}.$
Given an ideal $\mathfrak{n}=\prod_{v}\mathfrak{p}_{v}^{e_{v}}$ of $\mathcal{O}_{F}$, then
$K_{1}(\mathfrak{n}):=\prod_{v}K_{1}(\mathfrak{p}_{v}^{e_{v}})\subseteq\prod_{v}\mathrm{GL}_{N}(\mathcal{O}_{v}).$
Theorem 3.1.2 ([JPSS81, Thm. 5], [GH11, §13]).
Let $\pi_{v}$ be a smooth admissible irrep of $\mathrm{GL}_{N}(F_{v})$. Then there is $i_{\pi_{v}}\geq 0$ such that
$\dim\left((\pi_{v})^{K_{1}(\mathfrak{p}^{i})}\right)=\begin{cases}0&i<i_{\pi_{v}}\\ 1&i=i_{\pi_{v}}\end{cases}.$
Proof.
Note that [JPSS81, Thm. 5] only gives the case of $\pi_{v}$ generic. The full argument is sketched out in [GH11, §13]. ∎
Definition 3.1.3.
We call $c(\pi_{v})=\mathfrak{p}_{v}^{i_{\pi_{v}}}$ the conductor of $\pi_{v}$. The conductor $\mathfrak{n}_{\pi}$ of a global automorphic representation $\pi=\bigotimes_{v}^{\prime}\pi_{v}$ is defined as
$c(\pi):=\mathfrak{n}=\prod_{v}c(\pi_{v}).$
As local shorthand, we will often say $c(\pi_{v})=i_{\pi_{v}}$.
Definition 3.1.4.
If $\pi$ is a global automorphic representation of $\mathrm{GL}_{N}$, its newvector, defined only up to constant factor, is an element of the one-dimensional space $\pi^{K_{1}(\mathfrak{n}_{\pi})}$.
3.1.2. $\operatorname{Res}^{E}_{F}\mathrm{GL}_{N}$ version
In the unitary case, $G_{N,v}$ is one of the following:
  • If $v$ is split, $\mathrm{GL}_{N}(F_{v}\otimes_{F}E)\cong\mathrm{GL}_{N}(F_{v})^{2}$.
  • If $v$ is inert or ramified, $\mathrm{GL}_{N}(F_{v}\otimes_{F}E)\cong\mathrm{GL}_{N}(E_{w})$ for $w$ lying over $v$.
Applying Theorem 3.1.2 for $GL_{N}(F_{v})^{2}$ and $GL_{N}(E_{w})$ respectively lets us define local conductors as formal powers of $\mathfrak{p}_{v}$ as follows:
  • If $v$ is split: a pair $\mathfrak{p}_{w}^{i},\mathfrak{p}_{w^{\prime}}^{j}$ for integral $i,j$ and $w$, $w^{\prime}$ above $v$. We represent this as $\mathfrak{p}_{w}^{(i,j)}$ for ordered pair of integers $(i,j)$. If $\pi_{v}$ is conjugate self-dual, then $i=j$ so we can formally represent this as $\mathfrak{p}_{v}^{i}$ for a single integer $i$. We use shorthand $c(\pi_{v})=i$.
  • If $v$ is inert: a single $\mathfrak{p}_{w}^{i}$ which we represent formally by $\mathfrak{p}_{v}^{i}$ for integral $i$. We use shorthand $c(\pi_{v})=i$
  • If $v$ is ramified: a single $\mathfrak{p}_{w}^{2i}$ which we represent formally as $\mathfrak{p}_{v}^{i}$ for half-integral $i$. We use shorthand $c(\pi_{v})=2i$.
These correspond to analogous subgroups $K_{1}(\mathfrak{p}_{w}^{i})=K_{1,v}(i)\subseteq G_{N,v}$ that satisfy the dimension-one property of Theorem 3.1.2. We similarly define newvectors.
The global conductor $c(\pi)$ can therefore be thought of as an ideal of $\mathcal{O}_{E}$. If $\pi$ is further conjugate self-dual, then $c(\pi)$ can be represented as a formal product
$c(\pi)=\mathfrak{n}=\prod_{v}\mathfrak{p}_{v}^{k_{v}}$
where $v$ ranges over finite places of $\mathcal{O}_{F}$ and the $k_{v}$ are positive integers except when $v$ is ramified in which case they can also be half-integers.
Note.
Beware that there is a difference of a factor of $2$ between our local and global indexing at ramified primes. These can be distinguished by local indexing always being by an integer and global indexing always being by an ideal.
3.2. Rankin-Selberg Integral Representations
Our construction of test functions that pick out root numbers relies on the Rankin-Selberg integral for $\pi$ and $\pi^{\prime}$. We recall this now as presented in [Cog04], see also [Hum20, §§2,3].
3.2.1. Whittaker Functions
Consider the upper-triangular unipotent subgroup $U\subset GL_{N}$. Pick a nontrivial character $\psi$ of $F^{\times}\backslash\mathbb{A}^{\times}$ and extend it to a character of $U(k)\backslash U(\mathbb{A})$ through
$\psi\begin{pmatrix}1&x_{1}&\cdots&\\ &\ddots&\ddots&\vdots\\ &&1&x_{n-1}\\ &&&1\end{pmatrix}=\psi(x_{1}+\cdots+x_{n-1}).$
Then, for any automorphic function $\varphi$ on $\mathrm{GL}_{N}$, we get a Whittaker function:
[…]
in the Gelfand-Graev space $W(\psi)$ of functions satisfying $W_{\varphi}(xg)=\psi(x)W_{\varphi}(g)$ for all $x\in U(\mathbb{A})$. Through this transformation property, we think of $\varphi\mapsto W_{\varphi}$ as attempting to embed $\pi$ as a $G(\mathbb{A})$-rep into a “twisted quotient” $(U(\mathbb{A}),\psi)\backslash G(\mathbb{A})$. If this map $\pi\to W(\psi)$ is non-zero, $\pi$ is called generic. It is well-known that all cuspidal automorphic representations of $\mathrm{GL}_{N}$ are generic, and that the embedding $\pi\hookrightarrow W(\psi)$ is unique if it exists.
Therefore, If $\varphi=\prod_{v}\varphi_{v}$ is a pure tensor in $\pi=\bigotimes_{v}^{\prime}\pi_{v}$, we have a factorization
$W_{\varphi}=:\prod_{v}W_{\varphi_{v}}$
through a corresponding factorization of the unique $\pi\hookrightarrow W(\psi)$.
We similarly define $\widehat{W}_{\varphi}$ by integration against $\psi$ to get an element of $W(\psi^{-1})$.
3.2.2. Integral Representations
Let $\pi,\pi^{\prime}$ be cuspidal automorphic representations of $\mathrm{GL}_{N}/F$ and $GL_{N-1}/F$ respectively. For any $\varphi\in\pi$ and $\varphi^{\prime}\in\pi^{\prime}$, define the Rankin-Selberg integral:
[…]
Rankin-Selberg theory provides a factorization of (3.2.1): if both $\varphi$, $\varphi^{\prime}$ are pure tensors, then
[…]
Then as a special case of a theorem of Jacquet–Piatetskii-Shapiro–Shalika corrected by Matringe:
Theorem 3.2.1 ([JPSS81, Thm (4)], corrected in [Mat13, Cor 3.3]).
Let $\pi_{v}$ be a generic representation of $\mathrm{GL}_{N,v}$ and $\varphi_{v}$ its newvector. Let $\pi^{\prime}_{v}$ be an unramified representation of $\mathrm{GL}_{N-1,v}$ and $\varphi^{\prime}_{v}$ its unramified vector. Normalize both so that $W_{\varphi_{v}}(1)=W_{\varphi^{\prime}_{v}}(1)=1$. Then $I(s,W_{\varphi_{v}},W_{\varphi^{\prime}_{v}})=L(s,\pi_{v}\times\pi_{v}^{\prime})$.
Moreover, we have a global functional equation
(3.2.3) $I(s,\varphi,\varphi^{\prime})=I(1-s,\varphi^{\iota},(\varphi^{\prime})^{\iota}),$
where $\varphi^{\iota}(g):=\varphi(g^{-T})=\varphi(w_{N}g^{-T})$, since $\varphi$ is automorphic.
We also have a local functional equation ([Cog04, Thm 6.2] specialized to $N\times N-1$): define $W_{\varphi}^{\iota}(g):=W_{\varphi}(w_{N}g^{-T})$. Then,
(3.2.4) $I_{v}(1-s,W^{\iota}_{\varphi_{v}},\widehat{W}^{\iota}_{\varphi^{\prime}_{v}})=\omega_{\pi^{\prime}_{v}}(-1)^{N-1}\gamma(s,\pi_{v}\times\pi^{\prime}_{v},\psi)I_{v}(s,W_{\varphi_{v}},\widehat{W}_{\varphi^{\prime}_{v}})$
defining a $\gamma$-factor $\gamma(s,\pi_{v}\times\pi^{\prime}_{v},\psi)$ that is necessarily a rational function in $q_{v}^{-s}$ and doesn’t depend on $\varphi_{v},\varphi_{v}^{\prime}$.
3.3. Root Numbers
3.3.1. $\epsilon$-factors and definition
The local functional equation (3.2.4) lets us define $\epsilon$-factors from the local $L$-functions:
(3.3.1) $\epsilon(s,\pi\times\pi^{\prime},\psi):=\frac{\gamma(s,\pi_{v}\times\pi^{\prime}_{v},\psi)L(s,\pi_{v}\times\pi^{\prime}_{v})}{L(1-s,\pi_{v}^{\vee}\times(\pi^{\prime}_{v})^{\vee})}.$
These are the factors whose product defines the global functional equation for $L$-functions:
$L(s,\pi\times\pi^{\prime})=\epsilon(s,\pi\times\pi^{\prime})L(1-s,\pi^{\vee}\times(\pi^{\prime})^{\vee}).$
Here
$\epsilon(s,\pi\times\pi^{\prime})=\prod_{v}\epsilon(s,\pi_{v}\times\pi^{\prime}_{v},\psi)$
is therefore necessarily independent of $\psi$.
Factors $\epsilon(s,\pi):=\epsilon(s,\pi,\mathrm{triv})$ for a single representation $\pi$ can be defined through a similar $\mathrm{GL}_{N}\times\mathrm{GL}_{1}$-theory. Then
Theorem 3.3.1 ([Cog04, Prop 6.5, Thm 6.3]).
Let $\pi_{v}$ have conductor $\mathfrak{p}_{v}^{k}$. Then
$\epsilon(s,\pi_{v},\psi)=\epsilon(1/2,\pi_{v},\psi)q_{v}^{-k(s-1/2)}.$
Globally,
(3.3.2) $\epsilon(s,\pi)=\epsilon(1/2,\pi)|c(\pi)|^{1/2-s},$
which motivates:
Definition 3.3.2.
If $\pi$ is an automorphic representation of $\mathrm{GL}_{N}$, let the root number of $\pi$ be $\epsilon(1/2,\pi)$.
As a final useful formula, if $\omega_{\pi_{v}}$ is the central character of $\pi_{v}$,
$\epsilon(1/2,\pi_{v},\psi)\epsilon(1/2,\pi_{v}^{\vee},\psi)=\omega_{\pi_{v}}(-1).$
3.3.2. The (conjugate) self-dual case
We recall results about root numbers of (conjugate) self-dual representations (to translate statements from [GGP11] to the automorphic side, note that that the determinant of the parameter of a representation of $\mathrm{GL}_{N,v}$ the parameter of its central character).
Proposition 3.3.3 ([GGP11, Thm 5.1(1)]).
Let the representation $\pi_{v}$ of $\mathrm{GL}_{N}(F_{v})$ be self-dual with trivial central character. Then $\epsilon(1/2,\pi_{v},\psi)=\pm 1$ and is independent of $\psi$.
Proposition 3.3.4 ([GGP11, Thm 5.1(2)]).
Let the representation $\pi_{v}$ of $\mathrm{GL}_{N}(F_{v}\otimes_{F}E)$ be conjugate self-dual. Then, if $\psi\circ\sigma=\psi^{-1}$, $\epsilon(1/2,\pi_{v},\psi)=\pm 1$. This value is independent of $\psi$ if either $N$ is even or $\pi_{v}$ is conjugate orthogonal.
Proof.
This first statement follows from [GGP11, Thm 5.1(2)]. By the paragraph just afterwards, independence follows whenever the parameter $\varphi_{v}$ for $\pi_{v}$ has conjugate orthogonal determinant. This is equivalent to the stated conditions by the discussion after Lemma 5.2.3. ∎
Definition 3.3.5.
If $N$ is even or $\pi_{v}$ is conjugate orthogonal, call $\epsilon(1/2,\pi_{v})$ for $\psi\circ\sigma=\psi^{-1}$ as in 3.3.4 the standard root number of $\pi_{v}$.
Next, we recall much more difficult consequences of the endoscopic classification:
Theorem 3.3.6 ([Art13, 1.5.3(b)]).
Let $\pi$ be a self-dual, orthogonal-type automorphic representation of $\mathrm{GL}_{N}$. Then $\epsilon(1/2,\pi)=1$.
Proof.
Note that the trivial representation on $\mathrm{GL}_{1}$ is orthogonal-type. ∎
Theorem 3.3.7 ([Mok15, 2.5.4(b)]).
Let $\pi$ be a conjugate self-dual, conjugate orthogonal-type automorphic representation of $\operatorname{Res}^{E}_{F}\mathrm{GL}_{N}$. Then $\epsilon(1/2,\pi)=1$.
Proof.
The trivial representation on $\operatorname{Res}^{E}_{F}\mathrm{GL}_{1}$ is conjugate orthogonal-type. ∎
3.4. Involutions for Newforms : Self-Dual Case
We want to understand the functional equation (3.2.3) and the variant we need to compute epsilon factors for $\varphi$ a newvector. Reformulating a note by Toma [Tom21] in adelic language, we define:
  • $\mathfrak{n},\mathfrak{p}^{k}$ will as shorthand also be some chosen lifts of the corresponding ideals to $F_{v}^{\times}$ or $\mathbb{A}^{\times}$ determined by our choices of uniformizers $\varpi_{v}$. All terms subscripted with these ideals will depend on this choice.
  • $J_{\mathfrak{p}_{v}^{k}}:=J_{v,k}:=\operatorname{diag}(\varpi_{v}^{-k},\cdots,\varpi_{v}^{-k},1)\in\mathrm{GL}_{N}(F_{v})$.
  • $J_{\mathfrak{n}}\in\mathrm{GL}_{N}(\mathbb{A})$ for $\mathfrak{n}$ an ideal of $\mathcal{O}_{F}$ the analogous global product of $J_{\mathfrak{p}_{v}^{k}}$.
  • $\iota_{\mathfrak{p}_{v}^{k}}=\iota_{v,k}:\mathrm{GL}_{N}(F_{v})\to\mathrm{GL}_{n}(F_{v}):g\mapsto g^{-T}J_{\mathfrak{p}_{v}^{k}}$,
  • $\iota_{\mathfrak{n}}:\mathrm{GL}_{N}(\mathbb{A})\to\mathrm{GL}_{N}(\mathbb{A})$ the global version of $\iota_{\mathfrak{p}_{v}^{k}}$,
  • $\iota:=\iota_{\mathcal{O}_{F}}$ (consistent with the previous definition).
  • $\varphi^{\iota_{\mathfrak{n}}}:=\varphi\circ\iota_{\mathfrak{n}}$ for global test vectors $\varphi$.
We also define an involution on local and global Whittaker models: if $w_{N}$ is the long Weyl element as in (2.2.1), then
  • $W^{\iota_{\mathfrak{n}}}_{\varphi}(g):=W_{\varphi}(w_{N}g^{\iota_{\mathfrak{n}}})$,
  • $W^{\iota_{v,k}}_{\varphi_{v}}(g):=W_{\varphi_{v}}(w_{N}g^{\iota_{v,k}})$.
We justify this notation since globally, the left-invariance of $\varphi$ under $G(F)$ gives:
[…]
$=W^{\iota_{\mathfrak{n}}}_{\varphi}(g).$
We use here that the map $x\mapsto w_{N}x^{-T}w_{N}^{-1}$ preserves $\psi$ since
$w_{N}\begin{pmatrix}1&x_{1}&\cdots&\\ &\ddots&\ddots&\vdots\\ &&1&x_{n-1}\\ &&&1\end{pmatrix}^{-T}w_{N}^{-1}$ $=w_{N}\begin{pmatrix}1&-x_{1}&\cdots&\\ &\ddots&\ddots&\vdots\\ &&1&-x_{n-1}\\ &&&1\end{pmatrix}^{T}w_{N}^{-1}$
$=\begin{pmatrix}1&x_{n-1}&\cdots&\\ &\ddots&\ddots&\vdots\\ &&1&x_{1}\\ &&&1\end{pmatrix}.$
It also preserves the Haar measure on $U$ since it has finite order.
The same invariance of $\psi$ gives that locally, if $x\in U(F_{v})$,
(3.4.2) $W^{\iota_{v,k}}_{\varphi_{v}}(xg)=W_{\varphi_{v}}(w_{N}x^{-T}g^{-T}J_{v,k})=W_{\varphi_{v}}((w_{N}x^{-T}w_{N}^{-1})w_{N}g^{-T}J_{v,k})\\ =\psi(w_{N}x^{-T}w_{N}^{-1})W_{\varphi_{v}}(w_{N}g^{-T}J_{v,k})=\psi(x)W^{\iota_{v,k}}_{\varphi_{v}}(g).$
so that $\iota_{v,k}$ preserves the Gelfand-Graev representation. Next, we show that local Whittaker-newforms are actually eigenvectors of this involution:
Lemma 3.4.1.
Let $\pi_{v}$ be a self-dual generic irrep of $\mathrm{GL}_{N,v}$ of conductor $k$ and $\varphi_{v}\in\pi_{v}$ a newvector. Then $W^{\iota_{v,k}}_{\varphi_{v}}=\pm W_{\varphi_{v}}$ where both are in $W(\psi)$.
Proof.
By (3.4.2), $W^{\iota_{v,k}}_{\varphi_{v}}$ is in $W(\psi)$. Since $\iota_{v,k}$ stabilizes $K_{1,v}(k)$, $W^{\iota_{v,k}}_{\varphi_{v}}$ is also fixed by $K_{1,v}(k)$. It generates a space isomorphic to $\pi_{v}^{\vee}\cong\pi_{v}$ since for $h\in GL_{N,v}$,
$W^{\iota_{v,k}}_{h\varphi_{v}}(g)=W_{\varphi_{v}}(w_{N}g^{-T}J_{v,k}h^{-1})=W_{\varphi_{v}}(w_{N}g^{-T}J_{v,k}h^{-1}J_{v,k}^{-1}J_{v,k})=\\ W_{\varphi_{v}}(w_{N}(gJ^{-1}_{v,k}h^{T}J_{v,k})^{-T}J_{v,k})=W^{\iota_{v,k}}_{\varphi_{v}}(gJ_{v,k}^{-1}h^{T}J_{v,k})=(J_{v,k}h^{-T}J_{v,k}^{-1})W^{\iota_{v,k}}_{\varphi_{v}}(g).$
In total, by uniqueness of Whittaker models and newvectors, $W_{\varphi_{v}}^{\iota_{v,k}}$ has to be constant multiple of $W_{\varphi_{v}}$. Since $g\mapsto w_{N}g^{-T}J_{v,k}$ is an involution, this multiple has to be ±$\pm 1$. ∎
Classically for $\mathrm{GL}_{2}$, there is a link between the the local eigenvalues from Lemma 3.4.1 and the root number coming from the use of global $\iota_{\mathfrak{n}}$ in the proof of the functional equation for Hecke $L$-functions (e.g. [DS05, 5.10.2]). The correct generalization to $\mathrm{GL}_{N}$ involves $N\times(N-1)$ Rankin-Selberg $L$-functions. The standard functional equation only involves $\iota$; we give a variant for $\iota_{\mathfrak{n}}$:
Lemma 3.4.2.
Let $\pi_{v},\pi^{\prime}_{v}$ be generic unirreps of $\mathrm{GL}_{N,v}$, $\mathrm{GL}_{N-1,v}$ with central characters $\omega,\omega^{\prime}$ respectively. Let $\varphi_{v},\varphi_{v}^{\prime}\in\pi_{v},\pi_{v}^{\prime}$ respectively. Then
$I_{v}(1-s,W^{\iota_{v,k}}_{\varphi_{v}},\widehat{W}^{\iota}_{\varphi^{\prime}_{v}})\\ =\omega^{\prime}(-1)^{N-1}\omega^{\prime}(\varpi_{v}^{k})q_{v}^{k(N-1)(s-1/2)}\gamma(s,\pi_{v}\times\pi^{\prime}_{v},\psi)I_{v}(s,W_{\varphi_{v}},\widehat{W}_{\varphi^{\prime}_{v}}).$
Proof.
The local functional equation (3.2.4) gives:
$I_{v}(1-s,W^{\iota}_{\varphi_{v}},\widehat{W}^{\iota}_{\varphi^{\prime}_{v}})=\omega^{\prime}(-1)^{N-1}\gamma(s,\pi_{v}\times\pi^{\prime}_{v},\psi)I_{v}(s,W_{\varphi_{v}},\widehat{W}_{\varphi^{\prime}_{v}}).$
However, using that $\varphi^{\iota}$ has central character $\omega^{-1}$,
$I_{v}(s,W^{\iota_{v,k}}_{\varphi_{v}},\widehat{W}^{\iota}_{\varphi^{\prime}_{v}})$
[…]
$=\omega^{\prime}(\varpi_{v}^{k})q_{v}^{k(N-1)(1/2-s)}$
$\qquad\qquad\int_{U_{v}\backslash\mathrm{GL}_{N-1,v}}W^{\iota}_{\varphi_{v}}\begin{pmatrix}\varpi^{k}\mathbf{a}_{v}&\\ &1\end{pmatrix}\widehat{W}^{\iota}_{\varphi^{\prime}_{v}}(\varpi^{k}\mathbf{a}_{v})|\det\varpi^{k}\mathbf{a}_{v}|^{s-1/2}\,d(\varpi^{k}\mathbf{a}_{v})$
[…]
$=\omega^{\prime}(\varpi^{k})q_{v}^{k(N-1)(1/2-s)}I_{v}(s,W^{\iota}_{\varphi_{v}},\widehat{W}^{\iota}_{\varphi^{\prime}_{v}}),$
so the result follows. ∎
Next, we show that for $\pi_{v}$ unramified the eigenvalue of $\iota$ is always $1$:
Proposition 3.4.3.
Let $\pi_{v}$ be an unramified, self-dual generic irrep of $\mathrm{GL}_{N,v}$ and $\varphi_{v}$ an unramified vector. Then $W_{\varphi_{v}}=W^{\iota}_{\varphi_{v}}$ independently of $\psi$.
Proof.
By the Iwasawa decomposition and since both sides of the equality are unramified, it suffices to check equality on elements of the form $u\lambda(\varpi)$ for $u\in U_{N,v}$ and $\lambda$ a cocharacter of the diagonal torus. Here:
$W^{\iota}_{\varphi_{v}}(u\lambda(\varpi))=W_{\varphi_{v}}(w_{N}u^{-T}\lambda(\varpi)^{-T})=W_{\varphi_{v}}(w_{N}u^{-T}w_{N}^{-1}\lambda^{\prime}(\varpi)w_{N})\\ =W_{\varphi_{v}}(w_{N}u^{-T}w_{N}^{-1}\lambda^{\prime}(\varpi)),$
where if $\lambda=(\lambda_{1},\dotsc,\lambda_{n})$, then $\lambda^{\prime}:=-w_{N}\lambda w_{N}^{-1}=(-\lambda_{n},\dotsc,-\lambda_{1})$.
Factoring out the $u$ terms, it therefore suffices to show that
$\psi(u)W_{\varphi_{v}}(\lambda(\varpi))=\psi(w_{N}u^{-T}w_{N}^{-1})W_{\varphi_{v}}(\lambda^{\prime}(\varpi))\quad(=\psi(u)W_{\varphi_{v}}(\lambda^{\prime}(\varpi)).)$
Since $\pi_{v}$ is self-dual, its Satake parameter $(q_{v}^{\alpha_{1}},\dotsc,q_{v}^{\alpha_{N}})$ is the same as the inverse $(q_{v}^{-\alpha_{N}},\dotsc,q_{v}^{-\alpha_{1}})$. The result therefore follows from the Casselman-Shalika formula (e.g. [HKP10, Thm 6.5.1]). ∎
Bootstrapping off this gives our key result:
Theorem 3.4.4.
Let $\pi_{v}$ be a generic self-dual representation of $\mathrm{GL}_{N,v}$ of conductor $k$. Let $\varphi_{v}\in\pi_{v}$ be a newvector and let $\tau^{\psi}_{v,k}(\varphi_{v})=\pm 1$ be defined by $W_{\varphi_{v}}=\tau^{\psi}_{v,k}(\varphi_{v})W^{\iota_{v,k}}_{\varphi_{v}}$ as in Lemma 3.4.1. Then
$\tau^{\psi}_{v,k}(\varphi_{v})=\epsilon(1/2,\pi_{v},\psi)^{N-1}.$
If we further assume that $\pi_{v}$ has trivial central character and $N$ is even, then
$\tau^{\psi}_{v,k}(\varphi_{v})=\epsilon(1/2,\pi_{v})$
and neither depend on the choice of $\psi$.
Proof.
Pick an auxilliary tempered, unramified, self-dual representation $\pi^{\prime}_{v}$ of $\mathrm{GL}_{N-1,v}$ with Satake parameter $(q_{v}^{\alpha_{1}},\dotsc,q_{v}^{\alpha_{N-1}})$. Let $\varphi^{\prime}_{v}\in\pi^{\prime}_{v}$ be the unramified vector. Normalize $\varphi_{v},\varphi^{\prime}_{v}$ so that $W_{\varphi_{v}}(1)=W_{\varphi^{\prime}_{v}}(1)=1$. Then by Theorem 3.2.1,
$I_{v}(s,W_{\varphi},W_{\varphi^{\prime}})=L_{v}(s,\pi_{v}\times\pi_{v}^{\prime}).$
We can further assume that $\pi^{\prime}_{v}$ has trivial central character. Then the definition of epsilon factors (3.3.1) and Lemma 3.4.2 give
$I_{v}(s,W^{\iota_{\mathfrak{n}}}_{\varphi_{v}},W^{\iota}_{\varphi^{\prime}_{v}})=\epsilon(s,\pi_{v}\times\pi^{\prime}_{v},\psi)q_{v}^{k(N-1)(s-1/2)}L_{v}(s,\pi_{v}\times\pi^{\prime}_{v}).$
Since $\pi^{\prime}_{v}$ is self-dual, $W^{\iota}_{\varphi^{\prime}_{v}}=W_{\varphi^{\prime}_{v}}$ by Proposition 3.4.3. Together with Lemma 3.4.1, this shows that
$\tau^{\psi}_{v,k}(\varphi_{v})I_{v}(s,W_{\varphi_{v}},W_{\varphi^{\prime}_{v}})=\epsilon(s,\pi_{v}\times\pi^{\prime}_{v},\psi)q_{v}^{k(N-1)(s-1/2)}L_{v}(s,\pi_{v}\times\pi^{\prime}_{v})\\ \implies\epsilon(s,\pi_{v}\times\pi^{\prime}_{v},\psi)=\tau^{\psi}_{v,k}|\mathfrak{n}|^{(N-1)(1/2-s)}.$
However, since $\pi^{\prime}$ is unramified, we also have
$L_{v}(s,\pi_{v}\times\pi^{\prime}_{v})=\prod_{i=1}^{N-1}L_{v}(s+\alpha_{i},\pi_{v}).$
The definition of $\epsilon$-factors and multiplicativity of $\gamma$-factors (e.g. [Cog04, §3.1.6]) then gives:
$\epsilon(s,\pi_{v}\times\pi^{\prime}_{v},\psi)=\prod_{i=1}^{N-1}\epsilon(s+\alpha_{i},\pi_{v},\psi)=\epsilon(s,\pi_{v},\psi)^{N-1}$
using at the end that $\sum_{i}\alpha_{i}=0$ since $\pi^{\prime}_{v}$ is self-dual. This is the first claim.
If $\pi_{v}$ is self-dual with trivial central character, by 3.3.3 and (3.3.2),
$\epsilon(s,\pi_{v})=\epsilon(1/2,\pi_{v})q_{v}^{k(1/2-s)}=\pm q_{v}^{k(1/2-s)}.$
This implies that $\epsilon(1/2,s)=\tau^{\psi}_{v,k}(\varphi_{v})$ if $N-1$ is odd. ∎
Note.
The above theorem does not compute the local root number when $N$ is odd or when $\pi_{v}$ has central character. However, in those cases, any global $\pi$ is orthogonal-type and therefore satisfies $\epsilon(1/2,\pi)=1$ by Theorem 3.3.6. Therefore this doesn’t matter for our eventual global application.
3.5. Involutions for Newforms : Conjugate Self-Dual Case
We perform a similar argument for conjugate self-dual $\pi$. Let $E/F$ be a quadratic extension with $\sigma:\alpha\mapsto\bar{\alpha}$ the corresponding Galois conjugation action.
We fix our character $\psi:E^{\times}\to\mathbb{C}^{\times}$ defining Whittaker functions to be self-conjugate, i.e. $\psi\circ\sigma=:\psi^{\sigma}=\psi$. It is very important to note that this is not the standard choice $\psi^{\sigma}=\psi^{-1}$ used in [GGP11]. This annoying technical detail will add some extra central character factors that do not appear in the self-dual case.
For appropriate $v,\mathfrak{p}_{v},\mathfrak{n}$, we similarly define:
  • $J_{\mathfrak{p}_{v}^{k}}:=\operatorname{diag}(\varpi_{v}^{-k},\cdots,\varpi_{v}^{-k},1)\in\mathrm{GL}_{N}(F_{v})\subseteq\mathrm{GL}_{N}(F_{v}\otimes_{F}E)$ (note that if $v$ is ramified, we need to instead choose a uniformizer $\varpi_{v}^{1/2}=\varpi_{w}$ for $E_{w}$),
  • $J_{\mathfrak{n}}\in\mathrm{GL}_{N}(\mathbb{A}_{F})\subseteq\mathrm{GL}_{N}(\mathbb{A}_{E})$ for $\mathfrak{n}$ an ideal of $\mathcal{O}_{F}$, the analogous global product of $J_{\mathfrak{p}_{v}^{k}}$,
  • $J_{v,k}$ the locally-indexed version of $J_{\mathfrak{p}_{v}^{k}}$, as in §3.1.2.
  • $\bar{\iota}_{\mathfrak{p}_{v}^{k}}:\mathrm{GL}_{N}(F_{v}\otimes_{F}E)\to\mathrm{GL}_{N}(F_{v}\otimes_{F}E):g\mapsto\bar{g}^{-T}J_{\mathfrak{p}_{v}^{k}}$, and $\bar{\iota}_{v,k}$ the locally-indexed version.
  • $\bar{\iota}_{\mathfrak{n}}:\mathrm{GL}_{N}(\mathbb{A}_{E})\to\mathrm{GL}_{N}(\mathbb{A}_{E})$ the global version of $\bar{\iota}_{\mathfrak{p}_{v}^{k}}$.
  • $\bar{\iota}:=\bar{\iota}_{\mathcal{O}_{F}}$
  • $\varphi^{\bar{\iota}_{\mathfrak{n}}}:=\varphi\circ\bar{\iota}_{\mathfrak{n}}$ for global test vectors $\varphi$.
We can also define for Whittaker functions:
  • $W^{\bar{\iota}_{\mathfrak{n}}}_{\varphi}(g):=W_{\varphi}(w_{N}g^{\bar{\iota}_{\mathfrak{n}}})$,
  • $W^{\bar{\iota}_{\mathfrak{n}}}_{\varphi_{v}}(g):=W_{\varphi_{v}}(w_{N}g^{\bar{\iota}_{\mathfrak{n}}})$.
Following our choice of $\psi^{\sigma}=\psi$, this satisfies analogues of (3.4.1) and (3.4.2):
$W_{\varphi^{\bar{\iota}_{\mathfrak{n}}}_{v}}=W^{\bar{\iota}_{\mathfrak{n}}}_{\varphi_{v}},\qquad W^{\bar{\iota}_{v,i}}_{\varphi_{v}}(xg)=\psi(x)W^{\bar{\iota}_{v,k}}_{\varphi_{v}}(g).$
We also warn that if $\varpi\neq\bar{\varpi}$, then $\bar{\iota}_{v,i}$ isn’t necessarily an involution:
$\bar{\iota}_{v,k}\bar{\iota}_{v,k}g=g\operatorname{diag}((\varpi/\overline{\varpi})^{-k},\dotsc,(\varpi/\overline{\varpi})^{-k},1).$
However, this extra factor belongs to all level subgroups $K_{1,v}(k)$, so precomposition with $\bar{\iota}_{v,k}$ does act as an involution on all new- and oldvector spaces.
Using all this, we continue the analogy:
Lemma 3.5.1.
Let $\pi_{v}$ be a generic conjugate self-dual irrep of $\mathrm{GL}_{N}(F_{v}\otimes_{F}E)$ of conductor $k$ and $\varphi_{v}\in\pi_{v}$ a newvector. Then, provided that $\psi^{\sigma}=\psi$, we have $W^{\bar{\iota}_{v,k}}_{\psi,\varphi_{v}}=\pm W_{\psi,\varphi_{v}}$.
Proof.
This is a very similar argument to Lemma 3.4.1, after noting that precomposition with $\bar{\iota}_{v,i}$ does act as an involution on newvectors. ∎
For the functional equation:
Lemma 3.5.2.
Assume $\psi^{\sigma}=\psi$. Let $\pi_{v},\pi^{\prime}_{v}$ be generic unirreps of $\mathrm{GL}_{N}(F_{v}\otimes_{F}E),\mathrm{GL}_{N-1}(F_{v}\otimes_{F}E)$ with central characters $\omega,\omega^{\prime}$ respectively. Let $\varphi_{v},\varphi_{v}^{\prime}\in\pi_{v},\pi_{v}^{\prime}$ respectively. Then
$I_{v}(1-s,W^{\bar{\iota}_{v,k}}_{\varphi_{v}},\widehat{W}^{\bar{\iota}}_{\varphi^{\prime}_{v}})\\ =\omega^{\prime}(-1)^{N-1}\bar{\omega}^{\prime}(\overline{\varpi}_{v}^{k}){q_{v}}^{k(N-1)(1/2-s)}\gamma(s,\pi_{v}\times\pi^{\prime}_{v},\psi)I_{v}(s,W_{\varphi_{v}},\widehat{W}_{\varphi^{\prime}_{v}}).$
Proof.
As in the proof of Lemma 3.4.2, we have that
$I_{v}(s,W^{\bar{\iota}_{v,k}}_{\varphi_{v}},\widehat{W}^{\bar{\iota}}_{\varphi^{\prime}_{v}})=\bar{\omega}^{\prime}(\overline{\varpi}^{k})q_{v}^{k(N-1)(1/2-s)}I_{v}(s,W^{\bar{\iota}}_{\varphi_{v}},\widehat{W}^{\bar{\iota}}_{\varphi^{\prime}_{v}})$
noting that the $J_{v,k}$ gets conjugated in the $\bar{\iota}_{v,k}$-version of the computation (Note also that this conjugation does nothing unless $v$ is ramified and $k$ is a half-integer). However, conjugation preserves Haar measures since it is an involution, so we also have that
$I_{v}(s,W^{\bar{\iota}}_{\varphi_{v}},\widehat{W}^{\bar{\iota}}_{\varphi^{\prime}_{v}})=I_{v}(s,W^{\iota}_{\varphi_{v}},\widehat{W}^{\iota}_{\varphi^{\prime}_{v}}).$
The result follows as in Lemma 3.4.2. ∎
Continuing with the analogous unramified input to 3.4.3:
Proposition 3.5.3.
Let $v$ be non-split and $\pi_{v}$ be an unramified, conjugate self-dual irrep of $G_{N,v}=\mathrm{GL}_{N}(F_{v}\otimes_{F}E)$ and $\varphi_{v}$ an unramified vector. Then $W_{\varphi_{v}}=W^{\bar{\iota}}_{\varphi_{v}}$ whenever $\psi\circ\sigma=\psi$.
Proof.
This follows from 3.4.3 applied to $\mathrm{GL}_{N}(E_{w})$: for all $u\in U$, $\psi(\bar{u})=\psi(u)$ and for all cocharacters $\lambda$, $W_{\varphi_{v}}(\lambda(\varpi))=W_{\varphi_{v}}(\lambda(\overline{\varpi}))$ since $\varpi/\overline{\varpi}$ has norm $1$. ∎
We now arrive at the conjugate self-dual theorem identifying $\bar{\tau}_{n}(\varphi_{v})$ with the appropriate root number. This will enable us to construct test functions counting automorphic forms of conductor $\mathfrak{n}$ weighted by their root number.
Theorem 3.5.4.
Assume $\psi\circ\sigma=\psi$. Let $\pi_{v}$ be a generic conjugate self-dual representation of $\mathrm{GL}_{N}(F_{v}\otimes_{F}E)$ of conductor $c(\pi_{v})=k$ (as defined in §3.1.2). Let $\varphi_{v}\in\pi_{v}$ be a newvector and let $\bar{\tau}^{\psi}_{v,k}(\varphi_{v})=\pm 1$ be defined by $W_{\varphi_{v}}=\bar{\tau}^{\psi}_{v,k}(\varphi_{v})W^{\iota_{v,k}}_{\varphi_{v}}$ as in Lemma 3.5.1. Then
$\bar{\tau}^{\psi}_{v,k}(\varphi_{v})=\begin{cases}\epsilon(1/2,\pi_{v},\psi)^{N-1}&v\text{ non-split}\\ 1&v\text{ split.}\end{cases}$
Furthermore, in the non-split case, if either $N$ is even or $\pi_{v}$ is conjugate-orthogonal, let $\epsilon(1/2,\pi_{v})$ be the standard root number as in 3.3.5. Then:
$\bar{\tau}^{\psi}_{v,k}(\varphi_{v})=\begin{cases}\omega(a)\epsilon(1/2,\pi_{v})&N\text{ even}\\ 1&N\text{ odd},\,\pi_{v}\text{ conjugate-orthogonal}\end{cases}$
where $\omega$ is the central character of $\pi$ and $a\in F_{v}\otimes_{F}E$ is any element such that $\sigma(a)=-a$.
Proof.
If $v$ is split, then $\pi_{v}$ is a representation of the form $\rho_{v}\boxtimes\rho_{v}^{\vee}$ of $\mathrm{GL}_{N}(F_{v})^{2}$. If $W_{f}$ is a newvector in the $\psi$-Whittaker model for $\rho_{v}$, then $W_{f}\boxtimes W^{\iota_{\mathfrak{n}}}_{f}$ is a newvector for $\pi_{v}$. Since $\iota_{\mathfrak{n}}$ acts as an involution, $\bar{\iota}_{\mathfrak{n}}$ acts as the identity on this newvector.
The non-split case follows very similarly to the proof of Theorem 3.4.4. Explaining the details that change: We pick an auxiliary tempered, unramified, conjugate self-dual representation $\pi^{\prime}$ of $\mathrm{GL}_{N-1}(F_{v}\otimes_{F}E)$ with trivial central character and $\varphi^{\prime}\in\pi^{\prime}$. Using 3.5.2/3.5.3 instead of 3.4.2/3.4.3, we get that
$\epsilon(s,\pi_{v}\times\pi^{\prime}_{v},\psi)=\bar{\tau}^{\psi}_{v,k}(\varphi_{v})q_{w}^{k(N-1)(1/2-s)}.$
When $v$ is non-split, we can also show by the same calculations as in 3.4.4 that
$\epsilon(s,\pi_{v}\times\pi_{v}^{\prime},\psi)=\epsilon(s,\pi_{v},\psi)^{N-1}.$
Note that for unramified representations at non-split places $v$, conjugate self-dual is equivalent to self-dual since the valuation is unchanged by conjugation.
The second statement follows since
$\epsilon(1/2,\pi_{v},\pi,\psi)=\omega(a)\epsilon(1/2,\pi_{v},\psi_{0})$
and that we always have that $\epsilon(1/2,\pi_{v},\psi_{0})=\pm 1$ and is independent of $\psi_{0}$ under our conditions by 3.3.4. We also note that $\omega(a)=\pm 1$ when $\pi_{v}$ is conjugate-orthogonal or $N$ is even. ∎
Note.
In the split case, let $\pi_{v}=\rho_{v}\boxtimes\rho_{v}^{\vee}$ of $\mathrm{GL}_{N}(F_{v})^{2}$. Then
$\epsilon(1/2,\pi_{v},\psi)=\epsilon(1/2,\rho_{v},\psi)\epsilon(1/2,\rho_{v}^{\vee},\psi)=\omega_{\rho_{v}}(-1).$
If $a\in F_{v}^{2}$ such that $a=-\bar{a}$, then $a$ is of the form $(x,-x)$ and
$\omega_{\pi}(a)=\omega_{\rho_{v}}(x)\omega_{\rho_{v}}(-x)^{-1}=\omega_{\rho_{v}}(-1).$
In total $\epsilon(1/2,\pi_{v},\psi)=\omega_{\pi_{v}}(a)$. In other words, the value of $\bar{\tau}^{\psi}_{v,k}(\varphi_{v})$ for $v$ split is not $\epsilon(1/2,\pi_{v},\psi)^{N-1}$, but rather $\epsilon(1/2,\pi_{v})^{N-1}:=\epsilon(1/2,\pi_{v},\psi_{0})^{N-1}$.
In particular, the split case differs from the non-split case by a factor of $\omega(a)^{N-1}$.
Note.
It is at this point that the restriction to $N$ even in our main Theorem 10.3.2 appears. In the self-dual case, we get lucky that when $N$ is odd, all representations are orthogonal-type and therefore have global root number $1$. Here, however, there are conjugate-symplectic-type representations when $N$ is odd which can have arbitrary root number. Our techniques are completely unable to address this case.
4. Test Functions
Using the results of the previous section, we now construct test functions that let us compute (twisted) counts of representations.
4.1. Twisted Action
Recall from (2.2.1) the matrix $w_{N}$ which conjugated $\iota,\bar{\iota}$ into the involutions $\theta,\bar{\theta}$ preserving the standard pinning. We define twisted groups:
$\widetilde{G}_{N}=\begin{cases}G_{N}\rtimes\theta&\text{self-dual or orthogonal/symplectic case}\\ G_{N}\rtimes\bar{\theta}&\text{conjugate self-dual or unitary case.}\end{cases}$
We first realize our involutions $\iota_{\mathfrak{n}},\bar{\iota}_{\mathfrak{n}}$ through conjugation in the twisted group $\widetilde{G}_{N}$.
If $\pi$ is a (conjugate) self-dual generic automorphic representation of $G_{N}$, the twisted group $\widetilde{G}_{N}$ acts on $\pi$ since $\theta$ (or $\bar{\theta}$) acts on its Whittaker model through precomposition. Define $\widetilde{\pi}$ to be this extended representation of $\widetilde{G}_{N}$. This lets us define twisted traces $\operatorname{tr}_{\widetilde{\pi}}\tilde{f}$ for test functions $\tilde{f}$ on $\widetilde{G}_{N}$. Note that this all depends on the choice of $\psi$ used to define the Whittaker model.
Note.
In the conjugate self-dual case, this choice of $\psi$ is a priori different than in the previous section. Specifically: before, we were thinking of the local $G_{N,v}$ as $\mathrm{GL}_{N}(F_{v}\otimes_{F}E)$ and defining Whittaker models through characters $\psi$ of $F_{v}\otimes_{F}E$. Now, we are working with the Whittaker models of the algebraic group $G_{N,v}/F_{v}$ depending on a choice of character $\psi$ of $F_{v}$. This choice in turn determines a choice of character of $F_{v}\otimes_{F}E$ through composition with the trace map and only produces characters of $F_{v}\otimes_{F}E$ satisfying $\psi^{\sigma}=\psi$, as was required in §3.5.
Since each of the $\pi_{v}$ are also generic, each $\widetilde{G}_{N,v}$ also acts on $\pi_{v}$ through precomposition in the Whittaker model. This defines local twisted traces $\operatorname{tr}_{\tilde{\pi}_{v}}\tilde{f}_{v}$ for test functions $\tilde{f}_{v}$ on $\widetilde{G}_{N,v}$. Then, as is well-known:
Proposition 4.1.1.
Let $\tilde{f}=\prod_{v}\tilde{f}_{v}$ be a test function on $\widetilde{G}_{N}$. Then for all generic automorphic representations $\pi$ of $G_{N}$
$\operatorname{tr}_{\widetilde{\pi}}\tilde{f}=\prod_{v}\operatorname{tr}_{\widetilde{\pi}_{v}}\tilde{f}_{v}.$
Proof.
Through the isomorphism of $G(\mathbb{A})$-reps from $\pi$ to its global Whittaker model and computation (3.4.1), we can calculate $\operatorname{tr}_{\tilde{\pi}}\tilde{f}$ as the trace of a convolution operator on the global Whittaker model and then factor. ∎
Our consideration of twisted traces is motivated by the multiplication formulas:
(4.1.1) $\begin{cases}(J^{-1}_{v,k}w_{N}\rtimes\theta)W_{\varphi}=W_{\varphi_{v}}^{\iota_{v,k}}&\text{self-dual},\\ (J^{-1}_{v,k}w_{N}\rtimes\bar{\theta})W_{\varphi}=W_{\varphi_{v}}^{\bar{\iota}_{v,k}}&\text{conjugate self-dual.}\end{cases}$
—i.e. we can relate the eigenvalue of the appropriate $\iota$ on a newvector to traces against $Jw_{N}\rtimes\theta$.
In particular, if we define
$\tilde{f}_{v,k}=(J^{-1}_{v,k}w_{N}\rtimes\theta)\bar{\mathbf{1}}_{K_{1,v(k)}},$
and assume that $\pi_{v}$ has conductor $\mathfrak{p}_{v}^{k}$ with newvector $\varphi_{v}$, we get:
$\operatorname{tr}_{\tilde{\pi}_{v}}\tilde{f}_{v,k}=\operatorname{tr}_{\tilde{\pi}_{v}^{K_{1,v}(k)}}(\iota_{v,k})=\tau^{\psi}_{v,k}(\varphi_{v}).$
from 3.4.4. Therefore, $\tilde{f}_{v,k}$ works as our test function to probe $\epsilon(1/2,\pi_{v})$ if $\pi_{v}$ has conductor exactly $\mathfrak{p}_{v}^{k}$. However, it also picks up the traces against oldform spaces for $\pi_{v}$ of conductor less than $k$.
4.2. Twisted Traces on Spaces of Oldforms
We now study the extra oldform contributions.
4.2.1. Description of the Oldform Space
We first recall results from [Ree91] describing the space of $K_{1,v}(k)$-fixed vectors in any generic representation $\pi_{v}$ of $\mathrm{GL}_{N,v}$ in terms of a Hecke algebra for $GL_{N-1,v}$. This lets us compute the action of $J_{v,i}w_{N}\rtimes\theta$ on all oldforms instead of just the newvectors.
Let $\mathscr{H}_{N-1}$ be the unramified (with respect to $K=\mathrm{GL}_{N-1}(\mathcal{O}_{v})$) Hecke algebra for $\mathrm{GL}_{N-1,v}$. Let $T\subseteq\mathrm{GL}_{N,v}$ be the diagonal torus so that $\mathscr{H}_{N-1}$ has a Cartan basis of double coset indicator functions $\mathbf{1}_{K\lambda(\varpi_{v})K}$ for $\lambda$ a dominant cocharacter of $T$.
Every such $\lambda$ is a unique linear combination of the fundamental coweights
$\lambda_{i}:=(\overbrace{1,\dotsc,1}^{i},\overbrace{0,\dotsc,0}^{N-1-i})\qquad 1\leq i\leq N-1$
with all coefficients positive except for that of $\lambda_{N-1}$. For $a_{1},\dotsc,a_{N-1}\in\mathbb{Z}^{n}$, define
$T(a_{1},\dotsc,a_{N-1}):=\mathbf{1}_{K(a_{1}\lambda_{1}+\cdots+a_{N-1}\lambda_{N-1})(\varpi_{v})K}\in\mathscr{H}_{N-1}$
and
$\mathscr{H}_{N-1}^{\leq k}:=\left\langle T(a_{1},\dotsc,a_{N-1}):a_{i}\geq 0,\sum a_{i}\leq k\right\rangle\subseteq\mathscr{H}_{N-1}.$
There is a normalized action of $\mathscr{H}_{N-1}$ on $\pi_{v}$ given by
[…]
Theorem 4.2.1 (Slight modification of [Ree91, Thm. 1]).
Let $\pi_{v}$ be a generic irrep of $\mathrm{GL}_{N}(F_{v})$ with conductor $c$ and newvector $\varphi_{v}$. Then for $k\geq 0$, the map
$h\mapsto h*\varphi_{v}$
induces a linear isomorphism
$\Psi_{\pi_{v}}:\mathscr{H}_{N-1}^{\leq k}\to\pi_{v}^{K_{1}(\mathfrak{p}_{v}^{c+k})}.$
Proof.
The Satake transform lets us change between the basis for $\mathscr{H}_{N-1}^{\leq k}$ described here and the one in [Ree91]. ∎
4.2.2. Twisted Action on Oldvectors: Self-Dual Case
Using the isomorphism $\Psi_{\pi_{v}}$, we can compute a key formula the self-dual case:
Proposition 4.2.2.
Fix a generic self-dual representation $\pi_{v}$ of $\mathrm{GL}_{N}$ with conductor $c$ and newvector $\varphi_{v}\in\pi_{v}$. Thewe have
$(J_{v,c+k}^{-1}w_{N}\rtimes\theta)\Psi_{\pi_{v}}(h|\det|^{-1/2})=\tau_{c}(v_{0})\Psi_{\pi_{v}}(h(\varpi_{v}^{-k}(*)^{-T})|\det|^{-1/2})$
for all $h\in\mathscr{H}_{N-1}$ and where we recall $\tau_{c}(\varphi_{v})=\pm 1$ is the eigenvalue of $(J_{v,c}^{-1}w_{N}\rtimes\theta)$ acting on $\varphi_{v}$.
Proof.
Computing and recalling that $\Psi_{\pi_{v}}(\varphi)=h*\varphi_{v}$:
[…]
$=\tau_{c}(\varphi_{v})\Psi_{\pi_{v}}(h(\varpi_{v}^{-k}(*)^{-T})|\det|^{-1/2}),$
which finishes the argument. ∎
Proposition 4.2.3.
Fix a generic self-dual representation $\pi_{v}$ of $\mathrm{GL}_{N}$ with conductor $\mathfrak{p}_{v}^{c}$, and newvector $\varphi_{v}$. Then,
  • If $N$ is even:
    $\operatorname{tr}_{\widetilde{\pi}_{v}^{K_{1}(c+k)}}(J_{v,c+k}^{-1}w_{N}\rtimes\theta)=\tau_{c}(\varphi_{v})\begin{dcases}\binom{k/2+N/2-1}{N/2-1}&k\text{ even}\\ 0&k\text{ odd.}\end{dcases}$
  • If $N$ is odd:
    $\operatorname{tr}_{\widetilde{\pi}_{v}^{K_{1}(c+k)}}(J_{v,c+k}^{-1}w_{N}\rtimes\theta)=\tau_{c}(\varphi_{v})\begin{dcases}\binom{k/2+(N-1)/2}{(N-1)/2}&k\text{ even}\\ \binom{(k-1)/2+(N-1)/2}{(N-1)/2}&k\text{ odd.}\end{dcases}$
We recall $\tau_{c}(\varphi_{v})=\pm 1$ is the eigenvalue of $(J_{v,c}^{-1}w_{N}\rtimes\theta)$ acting on $\varphi_{v}$.
Proof.
By Proposition 4.2.2 and Theorem 4.2.1, it suffices to compute the trace of
$h|\det|^{-1/2}\mapsto h(\varpi_{v}^{-k}(*)^{-T})|\det|^{-1/2}$
on $\mathscr{H}_{N-1}^{\leq k}$. We use the basis of $T(a_{1},\dotsc,a_{N-1})|\det|^{-1/2}$.
Since $-\lambda_{i}$ is Weyl-conjugate to $\lambda_{N-1-i}-\lambda_{N-1}$ for $21\leq i\leq N-$ 2, we get that if $\sum a_{i}=d$,
(4.2.1) $T(a_{1},\dotsc,a_{N-1})(\varpi_{v}^{-k}g^{-T})=T(a_{N-2},a_{N-3},\cdots,a_{1},k-d)(g).$
The trace is then exactly the number fixed points of this involution: indices $(a_{i})_{i}$ such that
$(a_{1},\dotsc,a_{N-1})=(a_{N-2},a_{N-3},\cdots,a_{1},k-d).$
If $N-2$ is even, there are only fixed points when $k$ is even. In this case, the choices for $a_{1},\dotsc,a_{N/2-1}$ determine the fixed point and can be any values whose sum is $\leq k/2$. This gives the claimed formula.
If $N-2$ is odd, the choices for $a_{1},\dotsc,a_{(N-1)/2}$ still determine the fixed point. If $k$ is even, then $a_{(N-1)/2}$ is even and the rest can be any values whose sum is $\leq k-a_{(N-1)/2}$. If $k$ is odd, then the situation is the same except $a_{(N-1)/2}$ needs to be odd. Both case give the claimed formula. ∎
4.2.3. Twisted Action on Oldvectors: Conjugate Self-Dual Case
Similarly:
Lemma 4.2.4.
Fix a generic conjugate self-dual representation $\pi_{v}$ of $\mathrm{GL}_{N}(F_{v}\otimes_{F}E)$ with conductor $\mathfrak{p}_{v}^{c}$ and newvector $\varphi_{v}$. Then we have
$(J_{v,c+k}^{-1}w_{N}\rtimes\bar{\theta})\Psi_{\pi_{v}}(h|\det|^{-1/2})=\bar{\tau}_{c}(\varphi_{v})\Psi_{\pi_{v}}(h(\varpi_{v}^{-k}\bar{(*)}^{-T})|\det|^{-1/2})$
for all $h\in\mathscr{H}_{N-1}$ and where we recall $\bar{\tau}_{c}(\varphi_{v})=\pm 1$ is the eigenvalue of $(J_{v,c}^{-1}w_{N}\rtimes\bar{\theta})$ acting on $\varphi_{v}$.
Proof.
This is the same argument as Proposition 4.2.2 after noting that conjugation sends any $K_{N-1,v}$-double coset to itself. ∎
Proposition 4.2.5.
Fix a generic conjugate self-dual representation $\pi_{v}$ of $\mathrm{GL}_{N}(F_{v}\otimes_{F}E)$ with (locally-indexed as in §3.1.2) conductor $c(\pi_{v})=k$ and newvector $\varphi_{v}$. Then,
  • If $v$ is non-split and $N$ is even:
    $\operatorname{tr}_{\widetilde{\pi}_{v}^{K_{1}(c+k)}}(J_{v,c+k}^{-1}w_{N}\rtimes\bar{\theta})=\bar{\tau}_{c}(\varphi_{v})\begin{dcases}\binom{k/2+N/2-1}{N/2-1}&k\text{ even}\\ 0&k\text{ odd}\end{dcases}$
  • If $v$ non-split and $N$ is odd:
    $\operatorname{tr}_{\widetilde{\pi}_{v}^{K_{1}(c+k)}}(J_{v,c+k}^{-1}w_{N}\rtimes\bar{\theta})=\bar{\tau}_{c}(\varphi_{v})\begin{dcases}\binom{k/2+(N-1)/2}{(N-1)/2}&k\text{ even}\\ \binom{(k-1)/2+(N-1)/2}{(N-1)/2}&k\text{ odd}\end{dcases}$
  • If $v$ is split:
    $\operatorname{tr}_{\widetilde{\pi}_{v}^{K_{1}(c+k)}}(J_{v,c+k}^{-1}w_{N}\rtimes\bar{\theta})=\bar{\tau}_{c}(\varphi_{v})\binom{k+N-1}{N-1}.$
We recall $\tau_{c}(\varphi_{v})=\pm 1$ is the eigenvalue of $(J_{v,c}^{-1}w_{N}\rtimes\bar{\theta})$ acting on $\varphi_{v}$.
Proof.
For the non-split cases, $\mathrm{GL}_{N}(F_{v}\otimes_{F}E)=\mathrm{GL}_{N}(E_{w})$ for some $w|v$. The basis elements $T(a_{1},\dotsc,a_{N-1})$ defined via the analogous uniformizer $\varpi_{w}$ of $\mathcal{O}_{w}$ are fixed under conjugation. Therefore, these two cases follow from Proposition 4.2.3 applied to representations of $GL_{N}(E_{w})$.
For the split case, $\mathrm{GL}_{N}(F_{v}\otimes_{F}E)=\mathrm{GL}_{N}(F_{v})^{2}$ so the relevant version of $\mathscr{H}_{N-1}^{\leq k}$ is the product of two copies of the version from 4.2.3, where conjugation switches the two factors. There is then exactly one fixed point of the relevant action from 4.2.4 for each choice of the first factor. ∎
4.3. $\widetilde{E}^{\infty}_{\mathfrak{n}}$: Counts Weighted by Epsilon Factors
Next, we find a twisted trace that counts representations $\pi$ of level $\mathfrak{n}$ weighted by their root numbers $\epsilon(1/2,\pi)$.
4.3.1. Local Test Function
Working locally, Propositions 4.2.2 and 4.2.4 let us solve an upper-triangular system of linear equations to produce the local factors of the test function giving weighted counts of automorphic forms:
Corollary 4.3.1.
Consider $G_{N,v}$ in either the self-dual or conjugate self-dual case—i.e. $G_{v}=\mathrm{GL}_{N,v}$ or $\operatorname{Res}^{E}_{F}\mathrm{GL}_{N,v}$ for $v$ split, ramified or inert. Then there are coefficients $a_{N}(k,i)$ such that if we define
$\widetilde{E}_{v,k}:=\sum_{i=0}^{k}a_{N}(k,i)(J_{v,k-i}^{-1}w_{N}\rtimes\theta)\bar{\mathbf{1}}_{K_{1,v}(k-i)},$
then for $\pi_{v}$ a generic representation of $G_{v}$ we have
$\operatorname{tr}_{\pi_{v}}(\widetilde{E}_{k,v})=\begin{cases}\tau&c(\pi_{v})=k\\ 0&\text{else},\end{cases}$
where $\tau$ is the eigenvalue of $\iota_{v,k}$ or $\bar{\iota}_{v,k}$ acting on the newvector for $\pi_{v}$.
We can compute $a_{N}(k,i)$ exactly in many cases:
Proposition 4.3.2.
The following formulas hold for $\widetilde{E}_{v,k}$ from Corollary 4.3.1::
  • self-dual case, $N$ even,
    $\widetilde{E}_{v,k}:=\sum_{i=0}^{N/2}(-1)^{i}\binom{N/2}{i}(J_{v,k-2i}^{-1}w_{N}\rtimes\theta)\bar{\mathbf{1}}_{K_{1,v}(k-2i)},$
  • conjugate self-dual case, $N$ even, $v$ non-split:
    $\widetilde{E}_{v,k}:=\sum_{i=0}^{N/2}(-1)^{i}\binom{N/2}{i}(J_{v,k-2i}^{-1}w_{N}\rtimes\bar{\theta})\bar{\mathbf{1}}_{K_{1,v}(k-2i)},$
  • conjugate self-dual case, $v$ split
    $\widetilde{E}_{v,k}:=\sum_{i=0}^{N}(-1)^{i}\binom{N}{i}(J_{v,k-i}^{-1}w_{N}\rtimes\bar{\theta})\bar{\mathbf{1}}_{K_{1,v}(k-i)}.$
Proof.
This follows from the formulas in Propositions 4.2.2 and 4.2.4 and the identity
$\sum_{i=0}^{b+1}(-1)^{i}\binom{b+1}{i}\binom{a-i+b}{b}=\begin{cases}1&a=0\\ 0&\text{else}\end{cases},$
for $a\in\mathbb{Z}^{+}$. The combinatorial identity in turn follows from the simpler identity $\sum_{i=0}^{b+1}(-1)^{i}\binom{b+1}{i}i^{k}=0$ when $k<b+1$, as first proved by Euler (see [Gou78]).
See also [Bin17] for a similar version of this result for untwisted traces. ∎
4.3.2. Global Test Function
Putting our local test functions globally:
Definition 4.3.3.
In the self-dual case, if $\mathfrak{n}=\prod_{v}\mathfrak{p}_{v}^{k_{v}}$ is an ideal of $F$, define the test function
$\widetilde{E}^{\infty}_{\mathfrak{n}}=\prod_{v}\widetilde{E}_{\mathfrak{p}_{v}^{k_{v}}}$
on $\widetilde{G}_{N}(\mathbb{A}^{\infty})$ in terms of the $\widetilde{E}_{\mathfrak{p}_{v}^{k_{v}}}$ from Corollary 4.3.1.
In the conjugate-self dual case, we need to twist by a central character at non-split places to deal with the corresponding discrepancy in 3.5.4:
Definition 4.3.4.
In the conjugate self-dual case, let $\mathfrak{n}=\prod_{v}\mathfrak{p}_{v}^{k_{v}}$ be a conductor as in §3.1.2. Choose $a\in E^{\times}$ with $\bar{a}=-a$ (by Hilbert 90) and define the test function
$\widetilde{E}^{\infty}_{\mathfrak{n}}=\prod_{v\text{ split}}\widetilde{E}_{\mathfrak{p}_{v}^{k_{v}}}\prod_{v\text{ ns.}}\delta_{a^{N-1}}\star\widetilde{E}_{\mathfrak{p}_{v}^{k_{v}}}$
on $\widetilde{G}_{N}(\mathbb{A}^{\infty})$ in terms of the $\widetilde{E}_{\mathfrak{p}_{v}^{k_{v}}}$ from Corollary 4.3.1 (but now globally indexed).
Theorem 4.3.5.
Consider the self-dual case $G_{N}=\mathrm{GL}_{N}/F$. Choose $\mathfrak{n}$ an ideal of $\mathcal{O}_{F}$. Then for all generic self-dual automorphic representations $\pi$ of $G_{N}$:
$\operatorname{tr}_{\widetilde{\pi}^{\infty}}(\widetilde{E}^{\infty}_{\mathfrak{n}})=\begin{cases}\epsilon(1/2,\pi_{\infty})^{-1}\epsilon(1/2,\pi)&c(\pi)=\mathfrak{n},\pi\text{ symplectic-type}\\ \epsilon(1/2,\pi_{\infty},\psi)^{-N+1}&c(\pi)=\mathfrak{n},\pi\text{ orthogonal-type}\\ 0&c(\pi)\neq\mathfrak{n}\end{cases}$
(where the second case may depend on the choice of $\psi$ made to define $\widetilde{\pi}$).
Proof.
The $c(\pi)\neq\mathfrak{n}$ case follows from multiplying together the same cases of Corollary 4.3.1.
If $c(\pi)=\mathfrak{n}$, then let $\varphi=\prod_{v\nmid\infty}\varphi_{v}$ be in the space of newvectors for $\pi^{\infty}$. Then by Proposition 4.1.1, Corollary 4.3.1, and (4.1.1),
$\operatorname{tr}_{\widetilde{\pi}^{\infty}}(\widetilde{E}^{\infty}_{\mathfrak{n}})=\prod_{v\nmid\infty}\operatorname{tr}_{\widetilde{\pi}_{v}}(\widetilde{E}_{\mathfrak{p}_{v}^{k}})=\prod_{v\nmid\infty}\tau_{\mathfrak{p}_{v}^{k}}(\varphi_{v}).$
If $\pi$ symplectic-type, then $N$ is even and $\pi$ has trivial central character so Theorem 3.4.4 reduces this to
$\prod_{v\nmid\infty}\epsilon(1/2,\pi_{v})=\epsilon(1/2,\pi_{\infty})^{-1}\epsilon(1/2,\pi)$
which is independent of $\psi$. The result follows.
In general, the trace is $\epsilon(1/2,\pi_{\infty},\psi)^{-N+1}\epsilon(1/2,\pi,\psi)^{N-1}$. However, when $\pi$ is orthogonal-type, [Art13, Thm 1.5.3(b)] gives that $\epsilon(1/2,\pi)=\epsilon(1/2,\pi\times\mathrm{triv})=1$. ∎
For the conjugate self-dual case, we need to choose $\psi=\psi\circ\sigma$ so we can understand the action of $\widetilde{E}^{\infty}_{\mathfrak{n}}$ on newvectors using 3.5.4:
Theorem 4.3.6.
Consider the conjugate self-dual case $G_{N}=\operatorname{Res}^{E}_{F}\mathrm{GL}_{N}$ and define extensions of representations to $\widetilde{G}_{N}$ using $\psi$ such that $\psi\circ\sigma=\psi$. Choose $\mathfrak{n}$ a conjugate self-dual conductor (as in 3.1.2). Then for all generic conjugate self-dual automorphic representations $\pi$ of $G_{N}$ with central character $\omega$:
$\operatorname{tr}_{\widetilde{\pi}^{\infty}}(\widetilde{E}^{\infty}_{\mathfrak{n}})=\begin{cases}\omega_{\infty}(a)^{-1}\epsilon(1/2,\pi_{\infty})^{-1}\epsilon(1/2,\pi)&c(\pi)=\mathfrak{n},N\text{ even},\;\pi\text{ conj.-symp.}\\ \omega_{\infty}(a)^{-1}\epsilon(1/2,\pi_{\infty})^{-1}&c(\pi)=\mathfrak{n},N\text{ even},\;\pi\text{ conj.-orth.}\\ \omega_{\infty}(a)^{-N+1}\epsilon(1/2,\pi_{\infty},\psi)^{-N+1}&c(\pi)=\mathfrak{n},N\text{ odd},\;\pi\text{ conj.-symp.}\\ 1&c(\pi)=\mathfrak{n},N\text{ odd},\;\pi\text{ conj.-orth.}\\ 0&c(\pi)\neq\mathfrak{n}\end{cases}$
where $a$ appears in the definition of $\widetilde{E}^{\infty}_{\mathfrak{n}}$. Note that the $\epsilon$-terms without $\psi$ are standard root numbers with respect to some local $\psi_{0}\circ\sigma=\psi_{0}^{-1}$ as in 3.3.4.
Proof.
Similar to Theorem 4.3.5, the trace is 00 unless $c(\pi)=\mathfrak{n}$ in which case
$\operatorname{tr}_{\widetilde{\pi}^{\infty}}(\widetilde{E}^{\infty}_{\mathfrak{n}})=\prod_{v\nmid\infty}\bar{\tau}_{v,\mathfrak{p}_{v}^{k}}(\varphi_{v}).$
Recall the choice of $a\in E^{\times}$ such that $\bar{a}=-a$ to define $\widetilde{E}^{\infty}_{\mathfrak{n}}$. We use 3.5.4 with this choice of $a$ at all places:
In the first and second cases, the central character is conjugate-orthogonal and therefore valued in ±$\pm 1$ so $\omega^{N-1}=\omega$. Therefore, the trace is
$\omega_{\mathrm{spl}}(a)^{N-1}\prod_{v\nmid\infty}\epsilon(1/2,\pi_{v})\prod_{\begin{subarray}{c}v\nmid\infty\\ v\text{ n.s.}\end{subarray}}\omega_{v}(a)=\omega_{\infty}(a)^{-1}\epsilon(1/2,\pi_{\infty})^{-1}\epsilon(1/2,\pi),$
using the note after 3.5.4 at split places. For the second case, we additionally use Theorem 3.3.7 to compute that $\epsilon(1/2,\pi)=1$ globally.
The third case is similar to the first: the trace is equal to
$\epsilon(1/2,\pi_{\infty},\psi)^{-N+1}\omega_{\infty}(a)^{-N+1}\epsilon(1/2,\pi)^{N-1}.$
Since the global root number is independent of $\psi$, multiplying together 3.3.4 gives that it is ±$\pm 1$ even for our non-standard $\psi$. Therefore, the third factor vanishes.
The fourth case follows immediately from the corresponding case of 3.5.4. ∎
Note.
If $\pi$ is unramified at a split place $v$, then it can be shown that $\omega_{v}(a)=1$. Therefore, if $\pi$ is unramified at all split places, $\omega_{\mathrm{ns}}(a)=\omega_{\infty}(a)^{-1}$.
Beware again that for $N$ odd, the test function $\widetilde{E}^{\infty}_{\mathfrak{n}}$ fails to produce any information about root numbers.
4.4. $\widetilde{C}^{\infty}_{\mathfrak{n}}$: Unweighted Counts through a Twisted Trace
We now use the twisted trace-Paley-Weiner theorem of [Rog88] to modify $\widetilde{E}_{v,k}$ into a function $\widetilde{C}_{v,k}$ on $\widetilde{G}_{N,v}$ counting representations of conductor $\mathfrak{p}_{v}^{k}$ unweighted by their root number.
As the Paley-Weiner input:
Proposition 4.4.1.
Let $\widetilde{G}_{v}=G_{v}\rtimes\theta$ be a twisted group and $\widetilde{f}_{v}$ a test function on $\widetilde{G}_{v}$. Choose a constant $a_{\Theta}$ for each Bernstein component $\Theta$ of $G_{v}$. Then there exists a test function $\widetilde{f}_{v}^{\prime}$ on $\widetilde{G}_{v}$ such that if $\widetilde{\pi}_{v}$ is an extension to $\widetilde{G}_{v}$ of the self-dual irrep $\pi_{v}\in\Theta$, we have
$\operatorname{tr}_{\widetilde{\pi}_{v}}\widetilde{f}_{v}^{\prime}=a_{\Theta}\operatorname{tr}_{\widetilde{\pi}_{v}}\widetilde{f}_{v}.$
Proof.
This is a special case of the main theorem of [Rog88] taking in to account the characterization of twisted Bernstein components in §6 therein. Note that self-dual is the same as $\theta$-stable. ∎
In summary, a local twisted test function can be modified by rescaling it by a different constant $a_{\Theta}$ on each Bernstein component. We next see that root numbers are constant on representations of a given conductor belonging to a specific $\Theta$.
Proposition 4.4.2.
In the self-dual case, let $\Theta$ be a Bernstein component of $\mathrm{GL}_{N,v}$. Choose $k\geq 0$. Then there is a constant $C_{\Theta,k}$ such that for all self-dual irreps $\pi_{v}\in\Theta$ with conductor $c(\pi_{v})=k$, we have $\epsilon(1/2,\pi_{v})=C_{\Theta,k}$.
Proof.
Fix a supercuspidal $\boxtimes_{i\in I}\rho_{i}$ for $\prod_{i\in I}\mathrm{GL}_{n_{i}}$ with $\sum n_{i}=N$ that determines a self-dual infinitesimal character for $\Theta$. Without loss of generality, we assume that for all pairs $(i,j)$, either $\rho_{i}\neq\rho_{j}$ or there is no $s$ such that $\rho_{i}|\det|^{s}=\rho_{j}$.
Consider self-dual $\pi_{v}\in\Theta$ with infinitesimal character $\boxtimes_{i\in I}\rho_{i}|\det|^{s_{i}}$. By the Bernstein-Zelevinsky classification, there is a partition of the sequence $(\rho_{i}|\det|^{s_{i}})_{i}$ into segments $\delta_{j}$ of the form
$\delta_{j}=(\sigma_{j}|\det|^{r_{j}+\frac{1}{2}(t_{j}-1)},\sigma_{j}|\det|^{r_{j}+\frac{1}{2}(t_{j}-3)},\dotsc,\sigma_{j}|\det|^{r_{j}-\frac{1}{2}(t_{j}-1)})$
such that $\pi_{v}$ is the Langlands quotient of $\boxtimes_{j}\Delta_{j}$ where each $\Delta_{j}$ is the discrete representation of $GL_{n_{j}t_{j}}$ associated to the segment $\delta_{j}$. Our condition on the $\rho_{i}$ lets us assume without loss of generality that the multiset of $t_{j}$ copies of each $\sigma_{j}$ is the same as the multiset of $\rho_{i}$’s.
By the formulas from [Hen86, p. 153]
$\epsilon(s,\pi)=\prod_{j}\epsilon(s,\Delta_{j}).$
Furthermore, if $\sigma_{j}$ is ramified, then
(4.4.1) $\epsilon(s,\Delta_{j})=\prod_{k=1}^{t_{j}}\epsilon(s,\sigma_{j}|\det|^{r_{j}+\frac{1}{2}(t_{j}+1)-k})=q_{v}^{r_{j}t_{j}c_{\sigma_{j}}}\epsilon(s,\sigma_{j})^{t_{j}}.$
If $\sigma_{j}$ is unramified, then it is necessarily of the form $|\det|^{s}$ on $\mathrm{GL}_{1}$ and $\Delta_{j}$ is of the form $\mathrm{S}t(t_{j})|\det|^{r_{j}}$ where $\mathrm{S}t(t_{j})$ is the Steinberg representation on $\mathrm{GL}_{t_{j}}$. Then
(4.4.2) $\epsilon(s,\Delta_{j})=(-1)^{t_{j}-1}q_{v}^{r_{j}(t_{j}-1)}q_{v}^{(t_{j}-1)(1/2-s)}$
by standard formulas for the conductor and root number of the Steinberg (e.g. computed again from [Hen86, p. 153]).
Putting all the formulas together, let $J_{0}$ be the set of $j$ such that $\Delta_{j}$ is a Steinberg. Then we can compute the conductor
$c(\pi_{v})=\sum_{j\in J_{0}}(t_{j}-1)+\sum_{j\in J\setminus J_{0}}t_{j}c_{\sigma_{j}}=u-u_{0}+\sum_{i\in I}c_{\rho_{i}}$
for $u$ the total number of unramified $\rho_{i}$ and $u_{0}$ the total number of unramified $\sigma_{i}$.
Also, since $\pi_{v}$ is self-dual, the $j$’s decompose into those such that $r_{j}=0$ and $\sigma_{j}$ is self-dual and pairs $j,j^{\prime}$ such that $t_{j}=t_{j^{\prime}}$, $r_{j}=-r_{j^{\prime}}$, and $\sigma_{j}=\sigma_{j^{\prime}}^{\vee}$. Therefore, when multiplying (4.4.1) and (4.4.2) over all $j$, all the powers of $q_{v}^{r_{j}}$ cancel leaving:
$\epsilon(1/2,\pi)=\prod_{j\in J_{0}}(-1)^{t_{j}-1}\prod_{j\in J\setminus J_{0}}\epsilon(1/2,\sigma_{j})^{t_{J}}=(-1)^{u-u_{0}}\prod_{i\in I}\epsilon(1/2,\rho_{i})\\ =(-1)^{c_{\pi}-\sum_{i\in I}c_{\rho_{i}}}\prod_{i\in I}\epsilon(1/2,\rho_{i}).$
This final result is our $C_{\Theta,c(\pi_{v})}$ which only depends on the $\rho_{i}$ and $c(\pi_{v})$. ∎
Proposition 4.4.3.
Fix a character $\psi$ to define local $\epsilon$-factors. In the conjugate self-dual case, let $\Theta$ be a Bernstein component of $\mathrm{GL}_{N,v}$ for $v$ non-split. Choose $k\geq 0$. Then there is constant $C_{\Theta,k}$ such that for conjugate self-dual irreps $\pi_{v}\in\Theta$ with conductor $c(\pi_{v})=k$, we have $\epsilon(1/2,\pi_{v},\psi)=C_{\Theta,k}$.
Proof.
This is the same argument as Proposition 4.4.2 since invariance of the $v$-adic norm under conjugation means that conjugate-self duality forces the same pairing condition on the $r_{j}$. ∎
Now we are ready to define the local components of our test function that gives an unweighted count of representations:
Corollary 4.4.4.
In either the self-dual or conjugate self dual case, there is a test function $\widetilde{C}_{v,k}$ on $\widetilde{G}_{N,v}$ such that for all generic (conjugate) self-dual $\pi_{v}$
$\operatorname{tr}_{\widetilde{\pi}_{v}}\widetilde{C}_{v,k}=\begin{cases}1&c(\pi_{v})=k\\ 0&\text{else}\end{cases}.$
Proof.
For all $\pi_{v}$ in Bernstein components $\Theta$ such that $\operatorname{tr}_{\widetilde{\pi}_{v}}\widetilde{E}_{v,k}\neq 0$, we have that $C_{\Theta,k}^{-(N-1)}$ from Proposition 4.4.2 (resp. 4.4.3) is the same as the $\tau$ from Corollary 4.3.1 by Proposition 3.4.4 (resp. 3.5.4). Therefore, we just apply Proposition 4.4.1 to $\widetilde{E}_{k,v}$ with constants $a_{\Theta}=C_{\Theta,k}^{-(N-1)}$. ∎
Putting these together globally:
Definition 4.4.5.
If $\mathfrak{n}=\prod_{v}\mathfrak{p}_{v}^{k_{v}}$ is an ideal of $F$, define the test function on $\widetilde{G}_{N}(\mathbb{A}^{\infty})$:
$\widetilde{C}^{\infty}_{\mathfrak{n}}=\prod_{v}\widetilde{C}_{\mathfrak{p}_{v}^{k_{v}}}$
in terms of the $\widetilde{C}_{\mathfrak{p}_{v}^{k_{v}}}$ from Corollary 4.4.4 (but now globally indexed) in either the self-dual or conjugate self dual case.
This of course satisfies:
Corollary 4.4.6.
In either the conjugate self-dual or self-dual case, for all generic self-dual representations $\pi^{\infty}$ of $(G_{N})^{\infty}$,
$\operatorname{tr}_{\widetilde{\pi}^{\infty}}\widetilde{C}^{\infty}_{\mathfrak{n}}=\begin{cases}1&c(\pi^{\infty})=\mathfrak{n}\\ 0&\text{else}.\end{cases}$
Note.
Unlike the case of $\widetilde{E}^{\infty}_{\mathfrak{n}}$ (specifically against orthogonal or conjugate-orthogonal representations), the trace identity characterizing $\widetilde{C}^{\infty}_{\mathfrak{n}}$ applies to more general irreps of $(G_{N})^{\infty}$ than just the ones that are factors of automorphic representations. This will be important for our spectral argument to understand its transfers in Section 9.
5. Asymptotics for the Cuspidal Spectrum: Setup
To compute our counts, we need to isolate certain parts of the automorphic spectrum on $G_{N}$ through the trace formula:
  • In the self-dual or orthogonal/symplectic case, the cuspidal, self-dual representations of either orthogonal or symplectic-type with a given infinitesimal character $\lambda$ at infinity.
  • In the conjugate self-dual or unitary case, the cuspidal, conjugate self-dual representations with a given infinitesimal character $\lambda$ at infinity.
We do this using the endoscopic classification through the formalism of refined shapes of parameters from [DGG22, §4,5]. While the characterization of the discrete spectrum of $\mathrm{GL}_{N}$ in [MW89] might appear to be enough on the spectral side, giving a tractable geometric-side expression requires the full methods of [DGG22].
We freely use notation analogous to [DGG22] throughout. In the case of $G=\mathrm{SO}_{2N}^{\eta}$, we will suppress the bars from [Art13] that represent orbits under the outer automorphism and take the unbarred notations to represent these orbits. The $G=\mathrm{SO}_{2N}^{\eta}$ case is not a main focus of this paper and this change is irrelevant for other $G$. See the notation section §1.7 for more details.
Recall also the list of simple endoscopic groups of $\widetilde{G}_{N}$ from §2.2.2 and assumption 2.3.3 on our number fields.
Note.
We use the full power of the endoscopic classification and Taïbi’s inductive analysis [Taï17] for two purposes. First and most importantly, we need to separate symplectic-type and orthogonal-type automorphic representations. This could potentially be done directly on the classical group side, though we are not aware of any method without some variant of the stable multiplicity formula.
Secondly, we need Taïbi’s analysis to isolate cuspidal representations amongst the discrete ones. This also could be avoided if we just cared about the generic self-dual spectrum. For $\lambda$ regular enough—as in our application where we consider the limit of $\lambda$ further and further away from the walls of the Weyl chamber—the generic spectrum is the same as the full discrete spectrum by its classification in [MW89] essentially in terms of shapes and formula (5.1.1).
5.1. Infinitesimal characters
Recall that the infinitesimal character $\lambda$ is an invariant of irreducible representations of a reductive Lie group $G(\mathbb{R})$. We have $\lambda\in\Omega_{G}\backslash\operatorname{Hom}(\mathfrak{t},\mathbb{C})$, where $\mathfrak{t}$ is a Cartan subalgebra of $G(\mathbb{R})$. This data is the same as a map from Weyl orbits of $X^{*}(\widehat{\mathfrak{t}})=X_{*}(\mathfrak{t})\to\mathbb{C}$, which is further the same as a semisimple conjugacy class in $\widehat{\mathfrak{g}}$.
5.1.1. Infinitesimal characters in the classification
We need a uniform way to talk about infinitesimal characters at infinity for all $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$.
In the self-dual case, the embedding $\widehat{G}\hookrightarrow\mathrm{GL}_{N}(\mathbb{C})$ induces a map from semisimple conjugacy classes in $\widehat{\mathfrak{g}}_{\infty}$ to those in $M_{N}(F_{\infty}\otimes_{\mathbb{R}}\mathbb{C})$. This map is injective except when $G=\mathrm{SO}^{\eta}_{2n}$, where its fibers are orbits under the outer automorphism. We will elide this technicality since everything in the $\mathrm{SO}_{2n}^{\eta}$ case is up to outer automorphism and therefore consider an infinitesimal character of $G$ as a semisimple matrix up to conjugacy: in other words an unordered sequence
$\lambda=(\lambda_{1},\lambda_{2},\dotsc,\lambda_{N})$
with $\lambda_{i}\in F_{\infty}\otimes_{\mathbb{R}}\mathbb{C}$.
In the conjugate self-dual case, since the map $\widehat{G}\hookrightarrow\mathrm{GL}_{N}(\mathbb{C})\times\mathrm{GL}_{N}(\mathbb{C})$ is a restriction from ${}^{L}\!{G}\hookrightarrow{}^{L}\!{G}_{N}$, it is determined by its first coordinate. Therefore, the infinitesimal character can also be represented by a semisimple conjugacy class in $M_{N}(F_{\infty}\otimes_{\mathbb{R}}\mathbb{C})$ and therefore a similar unordered sequence.
It is also useful to package the tuple $\lambda_{v}=(\lambda_{1,v},...,\lambda_{n,v})$ as the generating function $\sum_{j}X^{\lambda_{j,v}}$, which, by abuse of notation, we will also denote $\lambda_{v}$. In this way, if $\bigoplus_{i}\tau_{i,v}[d_{i}]$ is a local Arthur parameter such that each $\tau_{i,v}$ has infinitesimal character $\lambda_{v}^{(i)}=\sum_{j=1}^{n_{i}}X^{\lambda_{j,v}^{(i)}}$, then we have the infinitesimal character assignment
(5.1.1) $\left(\bigoplus_{i}\tau_{i,v}[d_{i}]\right)_{\infty}\mapsto\sum_{i}\lambda_{v}^{(i)}\sum_{l=1}^{d_{i}}X^{\frac{d_{i}+1}{2}-l}.$
It can be seen from this that the character of $\tau[d]$ determines that of $\tau$.
5.1.2. Integrality
The infinitesimal characters of finite-dimensional representations of the $G_{v}$ for $v|\infty$ will play a special role. We express their classification in terms of generating functions.
For the self-dual case:
  • If $G=\mathrm{Sp}_{2n}$, then $\widehat{G}=\mathrm{SO}_{2n+1}$ and the $\lambda_{v}$ are elements of $1+\mathbb{Z}[X,X^{-1}]$ invariant under $X\mapsto X^{-1}$ and with each coefficient 00 or $1$.
  • If $G=\mathrm{SO}_{2n+1}$, then $\widehat{G}=\mathrm{Sp}_{2n}$ and the $\lambda_{v}$ are elements of $X^{1/2}\mathbb{Z}[X,X^{-1}]$ invariant under $X\mapsto X^{-1}$ and with each coefficient 00 or $1$.
  • If $G=\mathrm{SO}_{2n}$, then $\widehat{G}=\mathrm{SO}_{2n}$ and the $\lambda_{v}$ are elements $\mathbb{Z}[X,X^{-1}]$ invariant under $X\mapsto X^{-1}$ and with each coefficient 00 or $1$ except the constant coefficient which is 00 or $2$.
For the conjugate self-dual case:
  • If $N$ is odd, then the $\lambda_{v}$ are elements of $X^{1/2}\mathbb{Z}[X,X^{-1}]$ with each coefficient 00 or $1$.
  • If $N$ is even, then the $\lambda_{v}$ are elements of $\mathbb{Z}[X,X^{-1}]$ with each coefficient 00 or $1$.
Definition 5.1.1.
Call $\lambda_{v}$ integral if it is one of the cases above.
Note that this is the same as the $C$-algebraic condition of [BG14].
5.2. Central Characters
We collect some useful lemmas on central characters:
5.2.1. Central Characters and Shapes
Every parameter $\psi$ has a central character $\eta_{\psi}$, of order $\leq 2$ in the self-dual case. Through class field theory, we will abuse notation and use $\eta_{\psi}$ to denote the corresponding character of $\Gamma_{F}$. Note that
(5.2.1) $\eta_{\bigoplus_{i}\tau_{i}[d_{i}]}=\prod_{i}\eta_{\tau_{i}}^{d_{i}}.$
5.2.2. Central Characters and Transfer
We need the central character results described in [GGP11, §10]. We use throughout as in [KS99, §5.1] that any $G\in\widetilde{\mathcal{E}}_{\mathrm{ell}}(N)$ comes with an injection $(Z_{G_{N}})_{\theta}\hookrightarrow Z_{G}$ from the algebraic group of coinvariants.
First, as is known to experts:
Theorem 5.2.1.
In the self-dual case, let $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ be either $\mathrm{Sp}^{\eta}_{2n}$ or $\mathrm{SO}_{2n}^{\eta}$ and $\varphi_{v}$ be an $L$-parameter for $G_{v}$. Then the central character $\omega_{v}$ of $\pi_{v}\in\Pi_{\varphi_{v}}$ is determined by $\omega_{v}(-1)=\epsilon(1/2,\varphi_{v})/\epsilon(1/2,\det\varphi_{v})$.
Proof.
This is claimed as an expected property in [GGP11, §10]. A proof follows from the main result of [LR05] together with a consistency statement between Langlands-Shahidi epsilon factors and those from Arthur’s classification (see e.g. [AHKO23, B.2] for a discussion). ∎
Lemma 5.2.2.
Let $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ and $\varphi_{v}$ an $L$-parameter for $G_{v}$. Then the central character $\omega_{v}$ of $\pi_{v}\in\Pi_{\varphi_{v}}$ is constant on $\Pi_{\varphi_{v}}$ (i.e. determined by $\varphi_{v}$).
Proof.
As explained in [Mok15, Rmk 3.2.3], the endoscopic character identities [Art13, Thm 2.2.1(a)] and [Mok15, Thm 3.2.1(a)] together with the transformation property for transfer factors given by the generalization of 5.1.C in [KS99, p.53] can be used to show this for the central character restricted to the image of the injection $(Z_{G_{N,v}})_{\theta}\hookrightarrow Z_{G_{v}}$. In the conjugate self-dual case, this is all of $Z_{G_{v}}$.
In the self-dual case, $G_{v}$ is abelian if $N=1,2$. When $N>2$, $Z_{G_{v}}$ is trivial or ±$\pm 1$. It is trivial when $\widehat{G}\cong\mathrm{Sp}_{2n}$. Otherwise, $\omega(-1)=\epsilon(1/2,\varphi_{v})/\epsilon(1/2,\det\varphi_{v})$ by Theorem 5.2.1. ∎
We emphasize another consequence of the transfer factor transformation property after [KS99, 5.1.C]:
Lemma 5.2.3.
Let $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ and $\varphi_{v}$ an $L$-parameter for $G_{v}$. Let $\omega_{v}$ be the common central character of $\pi_{v}\in\Pi_{\varphi_{v}}$ as in Lemma 5.2.2.
Then, $\omega_{v}$ restricted to the image of $(Z_{G_{N,v}})_{\theta}\hookrightarrow Z_{G_{v}}$ uniquely determines the (conjugate) self-dual central character $\omega^{\prime}_{v}$ of the representation $\tilde{\pi}_{\varphi_{v}}$ of $\widetilde{G}_{N,v}$ corresponding to $\varphi_{v}$.
With notation as in the statement of Lemma 5.2.3, we will call this $\omega^{\prime}$ the transfer of $\omega$. Note that because of the possible non-triviality of $\lambda_{C}$ in the generalization after [KS99, 5.1.C], the transfer $\omega^{\prime}$ is not necessarily pullback to $Z_{G,v}$. In fact, in the self-dual case, the image of $(Z_{G_{N,v}})_{\theta}\hookrightarrow Z_{G_{v}}$ is trivial, so $\omega^{\prime}_{v}$ is actually completely determined by this $\lambda_{C}$ and therefore $G$:
  • If $N$ is odd, $\omega^{\prime}_{v}=\eta$ for $G=\mathrm{Sp}_{N-1}^{\eta}$.
  • If $N$ is even, $\omega^{\prime}_{v}=\eta$ for $G=\mathrm{SO}_{N}^{\eta}$ and $\omega^{\prime}_{v}=1$ for $G=\mathrm{SO}_{N+1}$.
In the conjugate self-dual case, it can be checked by working over the algebraic closure that $(Z_{G_{N}})_{\theta}\hookrightarrow Z_{G}$ is given on points111we ignore the difference between this map and its inverse which depends on the specific choice of transfer factors. In our eventual application, the distinction will be moot due to arbitrary choices of infinite-place isomorphisms $\overline{E}_{v}\xrightarrow{\sim}\mathbb{C}$ needed to parameterize infinitesimal characters. by $\beta\mapsto\beta/\bar{\beta}$. By Hilbert 90, this is a bijection over local fields, so the map $\omega_{v}\mapsto\omega^{\prime}_{v}$ is an injection.
  • If $N$ is odd, the determinant of a conjugate symplectic/orthogonal parameter is the same type. Therefore, the image of $\omega_{v}\mapsto\omega^{\prime}_{v}$ is all characters coming from conjugate symplectic or orthogonalparameters respectively depending on the types of parameters of $G$.
  • If $N$ is even, the determinant of a conjugate self-dual parameter is always conjugate-orthogonal. Therefore, the image of $\omega_{v}\mapsto\omega^{\prime}_{v}$ is all characters coming from conjugate orthogonal parameters and doesn’t depend on $G$.
In the case of $U^{+}_{N}$ we can be even more precise:
Lemma 5.2.4.
Consider $G=U_{N}^{+}\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$. Then the transfer $\omega_{v}\mapsto\omega^{\prime}_{v}$ is pullback through $Z_{G_{N,v}}\twoheadrightarrow(Z_{G_{N,v}})_{\theta}\xrightarrow{\sim}Z_{G_{v}}$.
Proof.
This follows by the endoscopic transfer identity [Mok15, Thm 3.2.1(a)] and [Mok15, Thm 2.5.1(c)]: if $\pi_{v}$ has parameter $\varphi_{v}$, then $\omega_{v}$ has parameter $\det\varphi_{v}$ which necessarily lands in ${}^{L}\!{Z}_{G_{v}}$. ∎
5.3. Shapes for the Self-Dual Case
We need a version of the theory of refined shapes for the self-dual case. Informally, a shape is a coarsening of the notion of an Arthur parameter, and corresponds to the part of discrete spectrum which we will isolate in our arguments. We use notation analogous to [DGG22, §2] and refer readers to [Art13] for more background. The key difference with [DGG22] is that the auxiliary data $\eta$ will now be a central character instead of a parity.
5.3.1. Cuspidal Shapes
One of the central results of the endoscopic classification, [Art13, Theorem 1.4.1], is that each self-dual cuspidal parameter $\tau\in\Psi(N)$ factors through a unique simple endoscopic datum $H$. We first show that for the parameters with which we are concerned, this group $H$ is determined by the central and infinitesimal characters of $\tau$.
Proposition 5.3.1.
Let $\tau\in\Psi(N)$ be self-dual, cuspidal parameter with integral infinitesimal character $\lambda$ at infinity and central character $\eta$. Then the endoscopic datum $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ through which $\tau$ factors is entirely determined by $\lambda$ and $\eta$.
Furthermore:
  • $\lambda$ is the infinitesimal character of a finite-dimensional representation of $H$.
  • If $G=\mathrm{SO}_{2n+1}$, then $\eta$ is trivial.
  • If $G$ is a form of $\mathrm{SO}_{2n}$, then $G=\mathrm{SO}_{2n}^{\eta}$. In this case, for each $v|\infty$, $\eta$ is trivial on $\operatorname{Gal}_{F_{v}}$ if and only if $n$ is even. In particular, $G$ necessarily has discrete series.
Proof.
This is the argument of [CR15, §3.11], but we allow for non-trivial central character.
Using $\lambda$
We will see that $\lambda$ determines $\widehat{G}$. First, $\lambda$ is integral so by the version of Clozel’s purity lemma stated in [CR15, Lemma 3.13], for each $v|\infty$, the $L$-parameter $\varphi_{v}:W_{\mathbb{R}}\to\mathrm{GL}_{n}\mathbb{C}$ associated to $\tau_{v}$ is a direct sum of copies of the trivial representation $\mathbf{1}$, the character $\epsilon_{\mathbb{C}/\mathbb{R}}:W_{\mathbb{R}}\twoheadrightarrow\mathbb{R}^{\times}\to\pm 1$, and, for $w\in\mathbb{Z}$, the representations
$I_{w}=\operatorname{Ind}_{\mathbb{C}^{\times}}^{W_{\mathbb{R}}}z^{-w/2}\bar{z}^{w/2}.$
Note that $I_{w}=I_{-w}$ and that $I_{0}=\mathbf{1}\oplus\epsilon_{\mathbb{C}/\mathbb{R}}$.
Then $\lambda_{v}$ is a list of numbers: a 00 for each $\mathbf{1}$ or $\epsilon_{\mathbb{C}/\mathbb{R}}$ and a pair $(w/2,-w/2)$ for each $I_{w}$. In addition $I_{w}$ can be conjugated into an orthogonal/symplectic group if and only if $w$ is even/odd. Since $\tau$ comes from a simple endoscopic group that has either orthogonal or symplectic dual, it follows that all the $w$ have the same parity.
In total, the non-zero numbers in $\lambda_{v}$ are either be all integers or all half-integers for each $\lambda_{v}$, depending on whether $\tau_{v}$ has image in an orthogonal or symplectic group, thereby determining $\widehat{G}$. In addition, it follows from our definition of integral that $\lambda_{v}$ is the infinitesimal character of a finite-dimensional representation on $G$.
Using $\eta$
Next, $\eta$ determines the remaining part of the datum corresponding to $G$. By the discussion on [Art13, p. 10], we have three cases:
  • If $\widehat{G}\cong\mathrm{Sp}_{2n}$, then $G=\mathrm{SO}_{2n+1}$ and $\eta$ is necessarily trivial.
  • If $\widehat{G}\cong\mathrm{SO}_{2n+1}$, then $G=\mathrm{Sp}^{\eta}_{2n}\cong\mathrm{Sp}_{2n}$.
  • If $\widehat{G}\cong\mathrm{SO}_{2n}$, then $G=\mathrm{SO}^{\eta}_{2n}$.
This completely determines $G$. We can also compute $\eta|_{\operatorname{Gal}_{F_{v}}}$ by computing $\det\tau_{v}$. In the $\widehat{G}\cong\mathrm{SO}_{2n}$ case, $\tau_{v}$ is a sum of $n$ different summands $I_{w_{i}}$ for $w_{i}$ even and therefore acts trivially on $\operatorname{Gal}_{F_{v}}$ if and only if $n$ is even. Therefore, $G$ necessarily has discrete series under our Assumption 2.3.3. ∎
5.3.2. Refined Shapes in the Self-Dual Case
Motivated by the above, we define:
Definition 5.3.2.
An refined shape is a sequence
$\Delta=(T_{i},d_{i},\lambda_{i},\eta_{i})_{i}$
up to permutation and where $(T_{i},d_{i})$ are positive integers, $\lambda_{i}$ is an infinitesimal character of rank $T_{i}$, and $\eta_{i}$ is a central character.
We say that $\psi\in\Delta$, or that $\psi$ has refined global shape $\Delta$ if $\psi=\bigoplus_{i}\tau_{i}[d_{i}]$ with each $\tau_{i}$ of rank $T_{i}$ and such that each $\tau_{i}$ has central character $\eta_{i}$ and infinitesimal character $\lambda_{i}$ at infinity.
Let $\Psi_{\mathrm{ell}}^{\Delta}$ be the set of all elliptic, self-dual parameters on $\mathrm{GL}_{n}$ of refined shape $\Delta$.
Beware that $\Psi_{\mathrm{ell}}^{\Delta}$ is empty unless all the bulleted conditions on $\lambda$ and $\eta$ of Proposition 5.3.1 hold.
Definition 5.3.3.
The refined shape $\Delta$ is integral if the total infinitesimal character of $\Delta$ as in (5.1.1) is integral, in the sense of Definition 5.1.1.
We also pick out a special refined shape:
Definition 5.3.4.
Let $\lambda$ be an integral infinitesimal character for a group of rank $N$ and $\eta$ a central character that is trivial if $\lambda$ has half-integral entries. Then $\Sigma_{\eta,\lambda}$ is the trivial refined shape $(N,1,\lambda,\eta)$.
5.3.3. Properties
Proposition 5.3.5.
Let $\Delta$ be an integral refined shape of rank $N$ such that all the $\lambda_{i}$ are integral (in the sense of 5.1.1) and $\eta_{i}$ is trivial whenever $\lambda_{i}$ has half-integer entries. Then there is a group $G(\Delta)\in\widetilde{\mathcal{E}}_{\mathrm{ell}}(N)$ such that all $\psi\in\Delta$ are parameters for $G$. Furthermore, $G(\Delta)$ has discrete series at infinity.
Proof.
Consider $\psi=\bigoplus_{i}\tau_{i}[d_{i}]\in\Psi_{\mathrm{ell}}^{\Delta}$. By the discussion on [Art13, pp. 33-34], it suffices to show that $\Delta$ determines two things:
  • whether each $\tau_{i}[d_{i}]$ has image in an orthogonal or symplectic group. Let the total ranks of the orthogonaland symplecticpieces be $N_{O}$ and $N_{S}$ respectively,
[…]
Then we will have
$G(\Delta)=\begin{cases}\mathrm{SO}_{N_{S}+1}\times\mathrm{Sp}^{\eta}_{N_{O}-1}&N_{O}\text{ odd}\\ \mathrm{SO}_{N_{S}+1}\times\mathrm{SO}^{\eta}_{N_{O}}&N_{O}\text{ even}\end{cases}.$
To show these two facts, apply Proposition 5.3.1 and note that if $d_{i}$ is odd, whether $\tau_{i}[d_{i}]$ has orthogonal or symplectic image is the opposite of that for $\tau_{i}$. If $d_{i}$ is even, then it is the same. In addition, the central character of the sum of the orthogonal-image $\tau_{i}[d_{i}]$ is determined through (5.2.1) by the central characters of the $\tau_{i}$ and which $\tau_{i}[d_{i}]$ have orthogonal image. ∎
This gives us two corollaries:
Corollary 5.3.6.
Let $G\in\widetilde{\mathcal{E}}_{\mathrm{ell}}(N)$. Then
$\Psi_{\mathrm{disc}}(G)=\bigsqcup_{\Delta:H(\Delta)=G}\Psi_{\mathrm{ell}}^{\Delta}.$
Proof.
This follows from the above and the same discussion in [Art13, pp. 33-34]. ∎
Corollary 5.3.7.
Let $\lambda$ be an integral infinitesimal character of rank $N$. Then
$G(\Sigma_{\eta,\lambda})=\begin{cases}\mathrm{Sp}^{\eta}_{2n}&N\text{ odd}\\ \mathrm{SO}_{2n+1}&N\text{ even, }\lambda\text{ has half-integer entries}\\ \mathrm{SO}^{\eta}_{2n}&N\text{ even, }\lambda\text{ has integer entries.}\end{cases}$
In particular, for every $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ and infinitesimal character $\lambda$ of a finite-dimensional representation of $G_{\infty}$, there is $\eta$ such that $H(\Sigma_{\eta,\lambda})=G$ (beware that $\Psi^{\Sigma_{\eta,\lambda}}_{\mathrm{ell}}$ is empty when the conditions from Proposition 5.3.1 don’t hold).
We collect some more properties of refined shapes $\Delta$:
  • By [Art13, (1.4.9)], All $\psi\in\Delta$ correspond to the same pair $(\mathcal{S}_{\psi},s_{\psi})$ up to conjugation. Call the common values $S_{\Delta}$ and $s_{\Delta}$.
  • All $\psi\in\Delta$ have the same Arthur-$\mathrm{SL}_{2}$.
  • The infinitesimal character at infinity and central character of $\psi\in\Delta$ is determined by formulas (5.1.1) and (5.2.1).
Finally, we relate integral refined shapes $\Delta$ to a class of Archimedean parameters defined in [AJ87], which we call AJ-parameters; see also [DGG22, §4.3]:
Lemma 5.3.8.
Let $\Delta$ be an integral refined shape. Then there exists an AJ-parameter $\psi_{\infty}^{\Delta}$ such that for all $\psi\in\Delta$, $\psi_{\infty}=\psi_{\infty}^{\Delta}$. Furthermore, the induced localization map $\mathcal{S}_{\Delta}\to\mathcal{S}_{\psi_{\infty}^{\Delta}}$ is also determined by $\Delta$.
Proof.
This is more-or-less the same proof as [DGG22, Lem. 4.3.4]
We have that $\psi_{\infty}$ is determined by the $\tau_{i,\infty}$. However, checking the cases from [CR15, 3.13], $\tau_{i,\infty}$ is a discrete parameter with infinitesimal character matching that of a finite-dimensional representation so it is the unique discrete series parameter of its infinitesimal character $\lambda_{i}$. This forces $\psi_{\infty}$ to be an AJ-parameter.
The localization statement comes similarly as in [DGG22, Lem. 4.3.4] by inspecting the formula [Art13, (1.4.9)] which holds both locally and globally. ∎
5.4. Shapes in the Conjugate Self-Dual Case
For easy reference, we recall the definition of refined shapes in the conjugate self-dual case from [DGG22, §4]:
Definition 5.4.1.
In the conjugate self-dual case, a refined shape is a sequence
$\Delta=(T_{i},d_{i},\lambda_{i},\eta_{i})_{i}$
up to permutation and where the $T_{i}$ and $d_{i}$ are positive integers, $\lambda_{i}$ is an infinitesimal character of rank $T_{i}$, and $\eta_{i}=\pm 1$.
As the in the self-dual case, the $\eta_{i}$ (which are now just a simple signs) are needed to determine $G(\Delta)$.
Definition 5.4.2.
In the conjugate self-dual case, the refined shape $\Delta$ is integral if the total infinitesimal character of $\Delta$ as in equation (5.1.1) is integral.
As explained in [DGG22, §4], all the analogous properties to §5.3.3 for integral refined shapes hold in the conjugate self-dual case. The arguments are in fact simpler since all $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ have discrete series at infinity.
5.5. Unrefined Shapes
Given a refined shape $\Delta$, we can forget some data to give an unrefined shape $\Box:=\Box(\Delta)=((T_{i},d_{i},\eta_{i}))_{i}$:
  • In the self-dual case, we forget all information about the infinitesimal character except whether it takes integral or half-integral values. For indexing consistency, we fold this information into $\eta$ which now becomes a pair $(\omega,\pm 1)$ of a quadratic central character and a sign.
  • In the conjugate self-dual case, we forget the infinitesimal character.
This still uniquely determines the group $G(\Delta)=G(\Box)$ as in [DGG22, Prop 4.3.1] and Proposition 5.3.5. We also define trivial unrefined shape
$\Sigma_{\eta}:=\Box(\Sigma_{\lambda,\eta}).$
Given infinitesimal character $\lambda$ for $G(\Delta)_{\infty}$, Let $\Box_{\lambda}$ be the set of $\Delta$ with total infinitesimal character $\lambda$ such that $\Box(\Delta)=\Box$. We will sometimes shorthand
$(\lambda_{1},\dotsc,\lambda_{i})\in\Box_{\lambda}\implies(T_{i},d_{i},\lambda_{i},\eta_{i})_{i}\in\Box_{\lambda}.$
since $\Box$ already contains the data of the $T_{i},d_{i},\eta_{i}$.
We define $\lambda[d]$ to be the total infinitesimal character of the simple shape $(T,d,\lambda,\eta)$ as in (5.1.1). Similarly, let $\boxtimes_{i}\lambda_{i}$ be the total infinitesimal character of $(T_{i},1,\lambda_{i},\eta_{i})_{i}$—i.e. concatenating all the coordinates of each $\lambda_{i}$ and ordering them increasingly.
Finally, to an unrefined shape $\Box=(T_{i},d_{i},\eta_{i})_{i}$ we associate a “reduced group”
$G_{F}(\Box):=\prod_{i}G((T_{i},d_{i},\eta_{i})).$
This is not in general an element of ~$\widetilde{\mathcal{E}}_{\mathrm{ell}}(N)$ and can be thought of as the smallest group through which $\psi\in\Box$ functorially factor. It will appear in bounds on growth rates of counts of automorphic reps $\pi$ associated to $\psi\in\Box$.
5.6. Norms of Infinitesimal Characters
To prove our bounds, we will compare two different norms of infinitesimal characters $\lambda$ of some $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$:
  • the dimension of the finite-dimensional representation of $G$ corresponding to $\lambda$:
    $\dim\lambda:=\dim^{G}\lambda:=C_{G}\prod_{\alpha\in\Phi_{+}(G)}\langle\alpha,\lambda\rangle$
    where the constant $C_{G}$ only depends on $G$,
  • a minimum over positive coroots of $G$:
    $m(\lambda):=m^{G}(\lambda):=\min_{\alpha\in\Phi_{+}(G)}\langle\alpha,\lambda\rangle.$
Given an unrefined shape $\Box$, we can also define
$\dim_{\Box}\lambda:=\dim^{G}_{\Box}\lambda:=\max_{(\lambda_{i})_{i}\in\Box_{\lambda}}\prod_{i}\dim\lambda_{i}$
Note that this can be 00 if $\Box_{\lambda}$ is empty.
Given a group $G$, we can compute the number of positive roots:
$P_{G}:=\frac{1}{2}(\dim G-\operatorname{rank}G).$
This gives us a bound:
Lemma 5.6.1.
$\dim_{\Box}\lambda\leq(\dim\lambda)m(\lambda)^{P_{G_{F}(\Box)}-P_{G}}.$
Proof.
The factors $\langle\alpha,\lambda\rangle$ in the Weyl dimension formula for $\dim_{\Box}\lambda$ are always a subset of those in $\dim\lambda$. However, $\dim_{\Box}\lambda$ has $P_{G}-P_{G_{F}(\Box)}$ fewer factors. ∎
6. Weight Asymptotics for the Cuspidal Spectrum
6.1. Setup and Trace Formula Decompositions
We now apply the techniques of [DGG22, §5] to compute our asymptotics. Fix either the self-dual or conjugate self-dual case. Let $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$. Choose:
  • a regular, integral infinitesimal character $\lambda$ for $G_{\infty}$ and $\mathrm{EP}_{\lambda}$ the associated Euler-Poincaré function, endoscopically normalized as in [DGG22, §3.3],
  • a set of unramified places $S$ and a test function $f_{S}$ in the truncated Hecke algebra $\mathscr{H}^{\mathrm{ur}}(G_{S})^{\leq\kappa}$ (see [ST16]),
  • a test function $f^{S,\infty}$ at other places.
Choose $\Box=(T_{i},d_{i},\eta_{i})$ a shape for $G$. We use the strategy of [DGG22, §9] to understand the asymptotics in $\lambda$ of
$I^{G}_{\Box}(\mathrm{EP}_{\lambda}f^{S,\infty}f_{S}):=\sum_{\Delta\in\Box_{\lambda}}I^{G}_{\Delta}(\mathrm{EP}_{\lambda}f^{S,\infty}f_{S}):=\sum_{\Delta\in\Box_{\lambda}}\sum_{\psi\in\Delta}I^{G}_{\psi}(\mathrm{EP}_{\lambda}f^{S,\infty}f_{S})\\ =\sum_{\Delta\in\Box_{\lambda}}\sum_{\psi\in\Delta}\sum_{\pi\in\Pi^{G}_{\psi}}m_{\pi}\operatorname{tr}_{\pi}(\mathrm{EP}_{\lambda}f^{S,\infty}f_{S}),$
where $I^{G}_{\psi}$ is the trace against the part of the discrete spectrum of $G$ associated to $\psi$ as in [DGG22, §2.6] through [Art13, Thm 1.5.2] or [Mok15, Thm 2.5.2].
In actuality, we will instead use the stable versions:
$S^{G}_{\Box}(\mathrm{EP}_{\lambda}f^{S,\infty}f_{S}):=\sum_{\Delta\in\Box_{\lambda}}S^{G}_{\Delta}(\mathrm{EP}_{\lambda}f^{S,\infty}f_{S})\\ :=\sum_{\Delta\in\Box_{\lambda}}\sum_{\psi\in\Delta}m_{\psi}\epsilon_{\psi}(s_{\psi})|\mathcal{S}_{\psi}|\operatorname{tr}_{\Pi^{G}_{\psi}}(\mathrm{EP}_{\lambda}f^{S,\infty}f_{S}),$
which are expressed in terms of traces against $A$-packets and certain signs and constants that appear in the stable multiplicity formulas [Art13, 4.1.2], [Mok15, 5.6.1] (note that we ignore the $\sigma$ terms since all our parameters are elliptic).
The key formula for our final application is, the following, where we recall from §5.5 that in the self-dual case $\Sigma_{\eta}$ is the unrefined shape determined by a rank, a character $\eta$, and a choice of parity of entries of $\lambda$:
Proposition 6.1.1.
In either the self-dual or conjugate self-dual case, let $\eta$ be such that $G(\Sigma_{\eta})=G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$. Then for any infinitesimal character $\lambda$ for $G$,
$\sum_{\psi\in\Sigma_{\eta,\lambda}}\operatorname{tr}_{\widetilde{\pi}^{\infty}_{\psi}}(f^{\infty})=\begin{cases}(1/2)S^{G}_{\Sigma_{\eta}}(\mathrm{EP}_{\lambda}(f^{\infty})^{G})&G=\mathrm{SO}_{2n}^{\eta}\\ S^{G}_{\Sigma_{\eta}}(\mathrm{EP}_{\lambda}(f^{\infty})^{G})&\text{else.}\end{cases}$
Proof.
This is an easier version of the argument in [DGG22, Prop 5.3.1]. We note that for simple shapes, the constants in the stable multiplicity formula are all trivial except for $m_{\psi}$ which is $2$ in the even orthogonal case. ∎
Following the above proposition, we wish to compute asymptotics in $\lambda$ of $S^{G}_{\Sigma_{\eta}}(\mathrm{EP}_{\lambda}(f^{\infty})^{G})$. In the rest of this section, we show inductively that the main term in these asymptotics is a geometric term consisting of evaluation at central elements. This mirrors our previous strategy to compute level asymptotics in [DGG22, §5]. In Sections 7 and 9, we will compute the relevant evaluations at central elements for the functions $\widetilde{E}_{\mathfrak{n}}^{\infty}$ and $\widetilde{C}_{\mathfrak{n}}^{\infty}$ previously constructed.
6.2. Geometric Estimation of Traces of $\mathrm{EP}_{\lambda}$
As input, we use the weight aspect bounds of [ST16] generalized to deal with nontrivial center as in [Dal22].
6.2.1. Computation of Constants
For $G/F$ a reductive group, $\lambda$ a regular integral infinitesimal character of $G_{\infty}$, $\omega_{\lambda}$ the central character of the corresponding finite-dimensional representation of $G_{\infty}$, and $f^{\infty}$ a test function on $G(\mathbb{A}^{\infty})$, the main term in our asymptotics is given by a linear combination of evaluation of $f^{\infty}$ at central elements:
(6.2.1) $\Lambda(G,\omega_{\lambda},f^{\infty}):=\tau^{\prime}(G)\sum_{\gamma\in Z_{G}(F)}\omega^{-1}_{\lambda}(\gamma)f^{\infty}(\gamma).$
Here, $\tau^{\prime}(G)$ is a modified Tamagawa number from [ST16, §6.6] under canonical times Euler-Poincaré measure.
Note that in our cases, $Z_{G}(F)$ always intersects the support of $f^{\infty}$ in a finite set.
6.2.2. Statement
Theorem 6.2.1 (Slight variant/Special case of [Dal22, Thm. 9.1.1]).
With notation as above, there are constants $A,B,C$ with $C\geq 1$ and depending only on $G$ such that,
$(\dim\lambda)^{-1}S^{G}(\mathrm{EP}_{\lambda}f_{S}f^{S,\infty})=\Lambda(G,\omega_{\lambda},f^{\infty})+O(m(\lambda)^{-C}q_{S}^{A+B\kappa}f_{S}(1)),$
Where $\Lambda$ is as in (6.2.1).
Proof.
Hyperendoscopy techniques as in the proof of [Dal22, Thm. 9.1.1] let us write $S^{G}$ in terms of $I^{H}$ for smaller groups. Then this follows from the exact same argument. ∎
6.3. Preliminary Bounds for the Induction
Since we do not need to keep track of functions at split places, the bounds of [DGG22, §7] simplify. Recall the definitions of $\mathcal{T}:=\mathcal{T}_{\Delta},\mathcal{T}_{i}:=\mathcal{T}_{\Delta,i}$ from [DGG22, §6.2.4].
Proposition 6.3.1.
Let $\Delta=(T,d,\lambda,\eta)$ be a simple shape and $f_{S}f^{S,\infty}$ a trace-positive test function on $G(\Delta)^{\infty}$ with $f_{S}$ unramified. Then there is a test function $(f^{S,\infty})^{\prime}$ on $G(1,d,\lambda,\eta)^{S,\infty}$ such that
$\left|S^{Td}_{\Delta}\left((f_{S}f^{S,\infty})^{Td}\right)\right|\leq S^{T}_{\Sigma_{\eta,\lambda}}\left((\mathcal{T}f_{S}(f^{S,\infty})^{\prime})^{T}\right).$
Proof.
This is the same argument as [DGG22, Prop. 7.3.2] with empty ${S_{s}}$. We use a slightly modified version of [DGG22, Lem. 6.3.7], noting that the proof does not actually require the exponent $d$. ∎
Proposition 6.3.2.
Let $\Delta=(T_{i},d_{i},\lambda_{i},\eta_{i})_{i}$ be a shape of rank $N$ and $f_{S}f^{S,\infty}$ a trace-positive test function on $G(\Delta)^{\infty}$ with $f_{S}$ unramified. Then there are test functions $f^{S,\infty}_{i}$ on each $G(\Sigma_{\eta_{i},\lambda_{i}})^{S,\infty}$ such that
[…]
i
SΣηi,λiTi
[…]
product
[…]
Proof.
This is like the proof of [DGG22, Prop 7.3.3]: apply [DGG22, Cor 7.2.2] with empty $S_{s}$ and then Proposition 6.3.1. ∎
6.4. Inductive Result
Now we can induct to get asymptotics of the contribution of $\Sigma_{\eta}$ to the stable trace:
Theorem 6.4.1.
Let $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ in either the self-dual or conjugate self-dual case and let $\lambda$ be an infinitesimal character for $G$. Then there are $A,B,C$ with $C\geq 1$ such that
$(\dim\lambda)^{-1}S_{\Sigma_{\eta}}^{G}(\mathrm{EP}_{\lambda}f_{S}f^{S,\infty})\\ =\begin{cases}\Lambda(G,\omega_{\lambda},f^{\infty})+O(m(\lambda)^{-C}q_{S}^{A+B\kappa}f_{S}(1))&G\neq\mathrm{SO}_{2,\mathrm{triv}}\\ 0&G=\mathrm{SO}_{2,\mathrm{triv}},\end{cases}$
where $\Sigma_{\eta}$ is the unrefined trivial shape such that $G(\Sigma_{\eta})=G$ and $\Lambda(\cdot)$ is given by (6.2.1).
In addition, if $\Box$ is an unrefined shape and $G=G(\Box)$,
(6.4.1) $(\dim_{\Box}\lambda)^{-1}|S_{\Box}^{G}(\mathrm{EP}_{\lambda}f_{S}f^{S,\infty})|\leq C(G,\Box,f^{\infty})+O(m(\lambda)^{-C}q_{S}^{A+B\kappa}f_{S}(1))$
for some inexplicit constant $C(G,\Box,f^{\infty})$ depending only on the arguments.
Proof.
This is the the same induction-on-classification-rank argument as in [DGG22, Thm. 9.3.2]. First, [DGG22, Lem. 9.3.1] reduces to the case where $f^{\infty}$ is trace-positive. In the self-dual case, there are three bases cases (those when $G$ is abelian):
  • $N=1$. Then $\Sigma_{\eta}$ is trivially the only shape so the bound follows from a very special case of Theorem 6.2.1.
  • $N=2$, $G=\mathrm{SO}_{2}^{\mathrm{triv}}$. Then there are no simple parameters since ${}^{L}\!{G}$ embeds into a split torus.
  • $N=2$, $G=\mathrm{SO}_{2}^{\eta}$, $\eta\neq\mathrm{triv}$: then the only parameters with non-trivial shape are those for $\eta_{1}\boxtimes\eta_{2}$ with $\eta=\eta_{1}\eta_{2}$. Since each component $(\eta_{i})_{v}$ for $v|\infty$ is valued in ±$\pm 1$, these only appear in our traces for finitely many $\lambda$. Therefore, we can bound $S^{G}=S^{G}_{\Sigma_{\eta}}+E$ for an error $E$ that is constant in $\lambda$ and apply Theorem 6.2.1.
In the conjugate self-dual case, the bases cases are $N=1$ and $G=U^{\pm}_{1}$ which follows as in the $N=1$ case above.
For the inductive step, first note that by [DGG22, 5.3.1] or an analogous statement in self-dual setting:
$S_{\Sigma_{\eta}}^{G}(\mathrm{EP}_{\lambda}f_{S}f^{S,\infty})=S^{G}(\mathrm{EP}_{\lambda}f_{S}f^{S,\infty})-\sum_{\Box\neq\Sigma_{\lambda}}S_{\Box}^{G}(\mathrm{EP}_{\lambda}f_{S}f^{S,\infty}).$
If we could show that the terms in the rightmost sum the have smaller growth in $m(\lambda)$ than $\dim(\lambda)$, then Theorem 6.2.1 would prove the bound we want for $\Sigma_{\eta}$. For this, outside of our base cases, we always have $P_{G_{F}(\Box)}<P_{G}$ so $\dim(\lambda)^{-1}\dim_{\Box}(\lambda)=O(m(\lambda)^{-1})$ by Lemma 5.6.1. In particular, knowing the second bound would imply that all the terms in the sum are lower order in $m(\lambda)$ and still exponential in $q_{v}^{\kappa}$ by [DGG22, Lem. 6.2.3].
Therefore, it suffices to show the second bound (6.4.1). For this,
$|S_{\Box}^{G}(\mathrm{EP}_{\lambda}f_{S}f^{S,\infty})|\leq\sum_{\Box_{\Delta}=\Box}|S^{N}_{\Delta}((f_{S}f^{S,\infty})^{N})|$
Each summand can be bounded by Proposition 6.3.2, again using [DGG22, Lem.6.2.3] to bound $\mathcal{T}_{i}f_{S}(1)$:
[…]
i
SΣηi,λiTi
[…]
i
Λ
[…]
+O(m(λ)CqSA+BκfS(1))
[…]
i
Λ
[…]
+O(m(λ)CqSA+BκfS(1))
[…]
product
[…]
product
[…]
Λ
[…]
product
[…]
Λ
[…]
by the inductive hypothesis.
The result then follows by noting that $|\Box_{\lambda}|$ is bounded above by a constant only depending on $\Box$ and taking a maximum over all constants that are in the exponent of $q_{v}$. Note that we can remove the dependence of the $\Lambda(G_{i},\omega_{\lambda_{i}},f^{S,\infty}_{i}\mathcal{T}_{i}f_{S})$ on $\omega_{\lambda_{i}}$ through upper-bounding the sum defining it by a sum of absolute values. ∎
6.5. From $S$ to $I$
For future use, we note that the bound also applies to $I$-terms, even though we will not use it in this specific paper.
Theorem 6.5.1.
Choose $G^{*}\in\widetilde{\mathcal{E}}_{\mathrm{ell}}(N)$ and let $G=G^{*}$ in the self-dual case or an extended pure inner form of $G^{*}$ in the conjugate self-dual case. If $\lambda\in\Box_{\lambda}$, the bounds in Theorem 6.4.1 hold with the $S^{G}_{\Box}(\mathrm{EP}_{\lambda}f_{S}f^{S,\infty})$ replaced by $I^{G}_{\Box}(\mathrm{EP}_{\lambda}f_{S}f^{S,\infty})$ after also replacing all $f^{\infty}$ terms on the right with $(f^{\infty})^{G^{*}}$.
Proof.
First, note that the case of non-quasisplit $G$ is conditional on [KMSW14] and its publicly unavailable reference “KMSb”.
For the proof, we apply [DGG22, Cor 10.2.2] (or its exactly analogous self-dual version), referring to the notation therein and noting that
$I^{G}_{\Box}(\mathrm{EP}_{\lambda}f_{S}f^{S,\infty})=\sum_{\Delta\in\Box_{\lambda}}\sum_{\pi_{0}\in\psi^{\Delta}_{\infty}}\operatorname{tr}_{\pi_{0}}(\mathrm{EP}_{\lambda})m^{G}(\pi_{0},\Delta,f_{S}f^{S,\infty}).$
If $\Box$ is stable (in particular when it is $\Sigma_{\lambda}$), the expansion in [DGG22, Cor 10.2.2] only has one term with $s=1$ so $I^{G}_{\Box}(\mathrm{EP}_{\lambda}f_{S}f^{S,\infty})=S^{G^{*}}_{\Box}(\mathrm{EP}_{\lambda}(f_{S}f^{S,\infty})^{G^{*}})$.
Otherwise, this follows from applying Theorem 6.4.1 to each term of [DGG22, Cor 10.2.2] or its self-dual analog. The dominant contribution comes from $s=1$. ∎
Note.
For a shape $\Box$ with some $d_{i}>1$, there is a subset $\Phi_{\Box}\subseteq\Phi_{+}(G)$ such that $\langle\alpha,\lambda\rangle$ is constant over $\alpha\in\Phi_{\Box}$. The $\alpha\in\Phi_{\Box}$ are exactly those not appearing in the computation of $\dim_{\Box}\lambda$.
An inspection of the argument above shows that the error term $O(m(\lambda)^{-1})$ in Theorems 6.4.1 and 6.5.1 can be improved to an analogous $O(m_{\Box}(\lambda)^{-1})$ where $m_{\Box}(\lambda)$ only minimizes over $\alpha\notin\Phi_{\Box}$. This is irrelevant to the current paper where we only consider $\Box=\Sigma$.
7. Shalika Germs and $(\widetilde{E}^{\infty}_{\mathfrak{n}})^{G}$
To apply Theorem 6.4.1, we need to understand the main term (6.2.1) given by a sum over central elements for $(\tilde{f}^{\infty})^{G}=(\widetilde{E}^{\infty}_{\mathfrak{n}})^{G}$, the function whose trace is a count of automorphic forms weighted by their root number. In this case, we wish to show that the main term (6.2.1) vanishes.
7.1. Transfers at Non-Regular Elements through Shalika Germs
In this section, we consider the general problem of understanding values of transfers at central elements. Consider a (possibly twisted) reductive group $\widetilde{G}$ and $H\in\mathcal{E}_{\mathrm{ell}}(\widetilde{G})$. We will eventually apply this to $\widetilde{G}=\widetilde{G}_{N}$ and $H$ a classical group as in §2.2.2.
We first note that for central $\gamma\in H$,
$(\widetilde{f}^{\infty})^{H}(\gamma)=\mathcal{O}_{\gamma}((\widetilde{f}^{\infty})^{H})=\mathcal{SO}_{\gamma}((\widetilde{f}^{\infty})^{H}).$
If $\gamma$ were strongly regular, then we could understand this through the standard orbital integral transfer identities as in [KS99]:
(7.1.1) $\mathcal{SO}^{H}_{\gamma}((\widetilde{f}^{\infty})^{H})=\sum_{i}\Delta^{\widetilde{G}}_{H}(\gamma,\widetilde{\gamma}_{i})\mathcal{O}^{\widetilde{G}}_{\widetilde{\gamma}_{i}}(\widetilde{f}^{\infty}),$
where $\gamma$ is a norm of the stable class containing the rational-class representatives $\widetilde{\gamma}_{i}$ and the $\Delta^{\widetilde{G}}_{H}$ are a choice of transfer factors. These would allow to evaluate $\mathcal{SO}_{\gamma}((\widetilde{f}^{\infty})^{H})$ in terms of orbital integrals of $f^{\infty}$. For arbitrary semisimple elements, doing so is much more complicated.
The standard technique (e.g. [Kot88, §3]) to resolve this issue is comparing Shalika germ expansions. This was implemented in full generality [LS07, Thm 2.4.A] for non-twisted endoscopy. We mimic the argument in the twisted case using a comparison of transfer factors which was the main technical result of [Wal08]. This argument is known to experts and seems implicit in the stabilization of the twisted trace formula—we just reconstruct it to deduce some bounds from the details.
We first recall the definition of non-regular stable orbital integrals:
Definition 7.1.1.
Let $H/F_{v}$ be quasiplit reductive and $\gamma\in H$ an arbitrary semisimple element. If the $\gamma_{i}$ are representatives for the rational classes in the stable class of $\gamma$, then
$\mathcal{SO}_{\gamma}:=\sum_{i}e(H_{\gamma_{i}})\mathcal{O}_{\gamma_{i}},$
where $e(H_{\gamma_{i}})$ is the Kottwitz sign of the necessarily connected $H_{\gamma_{i}}$ (i.e. $(-1)^{r}$ where $r$ is the difference between the split ranks of $H_{\gamma_{i}}$ and its quasisplit inner form).
The main goal of what follows is Proposition 7.1.6. For this, we compare two Shalika germ expansions of (7.1.1): with respect to $H$ on the left-hand side, and with respect to $\widetilde{G}$ on the right-hand side. The latter relies on a descent property for twisted transfer factors, which we begin by recalling below.
7.1.1. Descent for Twisted Transfer Factors
The descent property we need is the main result of [Wal08]; an English-language summary of a special case can be found in [Var09]. We begin with a very general setup: $G/F_{v}$ is an arbitrary reductive group, and we let $\widetilde{G}:=G\rtimes\theta$ for an automorphism $\theta$ fixing a pinning. Let $H\in\mathcal{E}_{\mathrm{ell}}(\widetilde{G})$ be such that there is a transfer map defined to $H$ instead of to a $z$-extension (this will be satisfied for all $(\widetilde{G},H)$ appearing in our self-dual and conjugate self-dual cases).
Fix semisimple $\tilde{\gamma}\in\widetilde{G}_{v}$ and $\gamma\in H_{v}$ such that $\gamma$ is the norm of $\tilde{\gamma}$ (i.e. $\tilde{\gamma}$ is the the stable class that is the transfer of $\gamma$). As explained in [LS07, §1.3-4], since $H$ is quasisplit, we can without loss of generality replace $\gamma$ by a stable conjugate so that $H_{\gamma}=H^{0}_{\gamma}$ is quasisplit. Then [Wal08, §3.5-6] constructs an $\bar{H}$ such that:
  • $\bar{H}$ is an endoscopic group of $(G_{\tilde{\gamma}}^{0})^{\mathrm{sc}}$ (Waldspurger more precisely constructs the the entire endoscopic quadruple).
  • $(H^{0}_{\gamma})^{\mathrm{sc}}$ is a quasisplit “non-standard endoscopic group” for $\bar{H}$ as in [Wal08, §1.7]. In particular, there is transfer of stable conjugacy classes from $\mathfrak{h}_{\gamma}$ to $\bar{\mathfrak{h}}$ constructed from an isomorphism of tori contained in rational Borels (note that this factors through the canonical projection $\mathfrak{h}_{\gamma}\to\mathfrak{h}_{\gamma,\mathrm{sc}}$).
Now, choose strongly $\widetilde{G}$-regular $X_{H}\in\mathfrak{h}_{\gamma}$. Then $X_{H}$ transfers to a stable class $X_{\bar{H}}$ in $\bar{\mathfrak{h}}$ and further to a stable class in $\mathfrak{g}_{\tilde{\gamma}}$. Choose $X_{G}$ in this stable class of $\mathfrak{g}_{\tilde{\gamma}}$. We can decompose it as
$X_{G}=\widetilde{X}_{G,\mathrm{sc}}+\widetilde{X}_{G,z},\qquad\widetilde{X}_{G,\mathrm{sc}}\in\mathfrak{g}_{\tilde{\gamma},\mathrm{sc}},\;\widetilde{X}_{G,z}\in\operatorname{\mathrm{Lie}}(Z_{G_{\widetilde{\gamma}}}).$
Finally, we let $\Delta^{\widetilde{G}}_{H}$ and $\Delta^{\mathfrak{g}_{\widetilde{\gamma},\mathrm{sc}}}_{\mathfrak{h}_{\gamma}}$ be the transfer factors from [KS99] for the groups/Lie algebras determined by the sub/super-scripts (but normalized as in [Wal08] without the $\Delta_{IV}$-factor).
Theorem 7.1.2 ([Wal08, Thm 3.9]).
With notation from above, there is a choice of transfer factors and a neighborhood $0\in U\subset\mathfrak{h}_{\gamma}$ such that for $X_{H}\in U$,
$\Delta^{\widetilde{G}}_{H}(\exp(X_{H})\gamma,\exp(X_{G})\widetilde{\gamma})=\Delta^{\mathfrak{g}_{\widetilde{\gamma},\mathrm{sc}}}_{\bar{\mathfrak{h}}}(X_{\bar{H}},X_{G,\mathrm{sc}}).$
Proof.
The statement of [Wal08, Thm 3.9] is an equality of relative transfer factors which, as a note afterwards points out, can become an equality of transfer factors after making the correct choices. We also comment on notation for the reader’s convenience: our $\widetilde{\gamma}$ is Waldspurger’s $y\eta y^{-1}$ and our $\gamma$ is their $\epsilon$. Our assumptions on $H$ let us ignore all subscripted $1$’s.
We write out Waldspurger’s map $\varphi$ explicitly as it is defined before Remark 3.8. Similarly, we explicitly write out $\bar{G}[y]$ as it is defined at the end of §3.5. ∎
As a nice corollary:
Corollary 7.1.3.
With notation from above, assume that $\dim H_{\gamma}=\dim G_{\widetilde{\gamma}}$. Then $\Delta^{\widetilde{G}}_{H}(\exp(X_{H})\gamma,\exp(X_{G})\widetilde{\gamma})$ is constant for $X_{H}$ close enough to 00.
In particular, we can define $\Delta^{\widetilde{G}}_{H}(\gamma,\widetilde{\gamma})$ to be this constant value.
Proof.
Such corollaries were originally pointed out in [LS07, §2.4]: by the dimension assumption, $\bar{H}$ is necessarily an elliptic endoscopic group of $(G^{0}_{\widetilde{\gamma}})_{\mathrm{sc}}$ of the same dimension, so it is necessarily the quasisplit inner form. Therefore $\Delta^{\mathfrak{g}_{\widetilde{\gamma},\mathrm{sc}}}_{\bar{\mathfrak{h}}}(X_{\bar{H}},X_{G,\mathrm{sc}})=1$ always. The result follows from Theorem 7.1.2 noting that different choices of transfer factors differ by a constant value. ∎
7.1.2. Shalika Germs
We recall some standard material on Shalika germs (see e.g. [MW16, §II.2]). Let $\widetilde{G}=G\rtimes\theta$ be a possibly twisted reductive group over $F_{v}$ and fix semisimple $\tilde{\gamma}\in\widetilde{G}_{v}$. For any test function $f$ on $\widetilde{G}_{v}$, there is a neighborhood $1\in U_{f}\subseteq G_{v,\tilde{\gamma}}$ on which the Shalika germ expansion holds: if $\mathfrak{u}_{f}=\exp^{-1}(U_{f})\subseteq\mathfrak{g}_{\tilde{\gamma}}$, then for all regular semisimple $X\in\mathfrak{u}_{f}$,
(7.1.2) $\mathcal{O}_{\exp(X)\tilde{\gamma}}(f)=\sum_{U\in\mathcal{N}(\mathfrak{g}_{\widetilde{\gamma}})}\Gamma^{\mathfrak{g}_{\widetilde{\gamma}}}_{U}(X)\mathcal{O}_{\exp(U)\widetilde{\gamma}}(f)$
where $\mathcal{N}(\mathfrak{g}_{\tilde{\gamma}})$ is the set of nilpotent $G_{v,X}$-orbits in $\mathfrak{g}_{\tilde{\gamma}}$ and $\Gamma^{\mathfrak{g}_{\tilde{\gamma}}}_{U}(X)$ is the Shalika germ—a smooth function on $\mathfrak{g}_{\tilde{\gamma}}^{\mathrm{rss}}$ independent of $f$ such that for all $\alpha\in F_{v}^{\times}$:
(7.1.3) $\Gamma^{\mathfrak{g}_{\widetilde{\gamma}}}_{U}(\alpha^{2}X)=|\alpha|_{v}^{-\dim\mathcal{O}_{U}}\Gamma^{\mathfrak{g}_{\widetilde{\gamma}}}_{U}(X).$
7.1.3. Comparing Germ Expansions
We can now mimic the argument of [LS07, 2.4.A]. Fix a reductive, possibly twisted, $\widetilde{G}=G\rtimes\theta$ over $F_{v}$, $H_{v}\in\mathcal{E}_{\mathrm{ell}}(\widetilde{G}_{v})$ and semisimple $\gamma\in H_{v}$.
Let $T_{H}$ be a maximal torus of $H_{v}$ that is also an elliptic maximal torus of $H_{\gamma}$. Let $j_{i}$ for $1\leq i\leq l$ be representatives for the the $H_{v}$-conjugacy classes of admissible embeddings $T_{H}\hookrightarrow H_{v}$.
For any test function $f$ on $G_{v}$, there is a neighborhood $U\subseteq T_{H}$ such that the Shalika germ expansion holds for $f^{H}$ on each $j_{i}(U)$. Then for strongly regular $X_{H}\in\mathfrak{u}=\exp^{-1}(U)$,
(7.1.4) $\mathcal{SO}_{\exp(X_{H})\gamma}(f^{H})=\sum_{i}\sum_{U\in\mathcal{N}(\mathfrak{h}_{\gamma_{i}})}\Gamma^{\mathfrak{h}_{\gamma_{i}}}_{U}(j_{i}(X_{H}))\mathcal{O}_{\exp(U)\gamma_{i}}(f^{H})$
where we use the shorthand $\gamma_{i}:=j_{i}(\gamma)$. This gives us:
Lemma 7.1.4.
Choose a test function $f^{H}$ on $H$, semisimple $\gamma\in H$, and strongly regular $X_{H}\in\mathfrak{t}_{H}$ (i.e. the Lie algebra of our elliptic torus $T_{H}$ of $H_{\gamma}$). Then the function
$r(\alpha):=\mathcal{SO}_{\exp(\alpha^{2}X_{H})\gamma}(f^{H})$
for small enough $\alpha$ is of the form
$r(\alpha)=C\mathcal{SO}_{\gamma}(f^{H})+\sum_{d_{i}}C_{d_{i}}|\alpha|_{v}^{-d_{i}}$
for $C\neq 0$, some set of $d_{i}>0$, and arbitrary constants $C_{d_{i}}$.
Proof.
This is the intermediate result of [Kot88, §3] up until the third equation on page 639. Summarizing the argument, Kottwitz computes the homogeneous of degree-00 part of (7.1.4) after applying (7.1.3): first, the main result of [Rog81] computes all the $\Gamma_{1}$ terms on the elliptic torus $T_{H}$. Then, formulas from [Kot86] simplify the remainders into the definition of the non-regular stable orbital integral [^2] We warn that this definition is not simply the sum of orbital integrals and instead involves “Kottwitz signs” of centralizers as coefficients..
Kottwitz uses a clever choice of measure normalization to simultaneously remove the formal degree of the Steinberg from the formula in [Rog81] and to simplify the relation between embeddings $j_{i}$ and the non-regular stable orbital integral. ∎
Fix $X:=X_{H}\in\mathfrak{t}_{H}$ such that $\exp(\alpha^{2}X)\gamma$ is strongly regular for some open set of small enough $\alpha$. Then, setting $\widetilde{\gamma}_{i}=\tilde{\jmath}_{i}(\gamma)$, we get the other side of the comparison from (7.1.1):
(7.1.5) $\mathcal{SO}_{\exp(X)\gamma}(f^{H})\\ =\frac{D^{G}(\widetilde{\gamma}\exp(\tilde{\jmath}(X)))^{1/2}}{D^{H}(\gamma\exp(X))^{1/2}}\sum_{i}\Delta^{\widetilde{G}}_{H}(\exp(X)\gamma,\tilde{\jmath}_{i}(\exp(X))\widetilde{\gamma}_{i})\mathcal{O}_{\tilde{\jmath}_{i}(\exp(X))\widetilde{\gamma}_{i}}(f),$
where we again use Waldspurger’s convention for transfer factors (leaving out the $\Delta_{IV}$ term) and the $D$ terms are Weyl discriminants:
$D^{H}(\gamma)=|\det(1-\operatorname{Ad}(\gamma)|_{\mathfrak{h}/\mathfrak{h}_{\gamma}})|_{v}$
(note that the $D^{G}$-term is stable-conjugacy invariant and therefore independent of the suppressed $i$).
We can further restrict our choice of $X$ so that the quotient of $D$-terms simplifies to an expression
$\frac{D^{G}(\widetilde{\gamma})^{1/2}}{D^{H}(\gamma)^{1/2}}\frac{D^{\mathfrak{g}_{\tilde{\gamma}}}(\tilde{\jmath}(X))^{1/2}}{D^{\mathfrak{h}_{\gamma}}(X)^{1/2}}$
that involves a similar term for Lie algebras:
$D^{\mathfrak{h}}(X)=|\det(\operatorname{ad}(X)|_{\mathfrak{h}/\mathfrak{h}_{X}})|_{v}.$
By Theorem 7.1.2, for small enough $X_{H}$, the transfer factors in (7.1.5) simplify to
$\Delta^{\mathfrak{g}_{\tilde{\gamma}_{i},\mathrm{sc}}}_{\bar{\mathfrak{h}}_{i}}(X_{\bar{H}_{i}},\tilde{\jmath}_{i}(X)_{\mathrm{sc}})$
after constructing appropriate $\bar{H}_{i}$ that is an endoscopic group of $(G_{\tilde{\gamma}_{i}}^{0})_{\mathrm{sc}}$ as in §7.1.1 and where $X_{\bar{H}_{i}}$ is the (non-standard) transfer of $X$ to $\bar{H}_{i}$.
We can then Shalika-germ expand again, this time on $\widetilde{G}$ and around $\tilde{\gamma}_{i}$; for $X$ small enough, the identity (7.1.5) becomes:
(7.1.6) $\mathcal{SO}_{\exp(X)\gamma}(f^{H})=\frac{D^{G}(\widetilde{\gamma})^{1/2}}{D^{H}(\gamma)^{1/2}}\frac{D^{\mathfrak{g}_{\tilde{\gamma}}}(\tilde{\jmath}(X))^{1/2}}{D^{\mathfrak{h}_{\gamma}}(X)^{1/2}}\\ \times\sum_{i}\sum_{U\in\mathcal{N}(\mathfrak{g}_{\widetilde{\gamma}_{i},\mathrm{sc}})}\Delta^{\mathfrak{g}_{\widetilde{\gamma}_{i},\mathrm{sc}}}_{\bar{\mathfrak{h}}_{i}}(X_{\bar{H}_{i}},\tilde{\jmath}_{i}(X)_{\mathrm{sc}})\Gamma^{\mathfrak{g}_{\widetilde{\gamma}_{i}}}_{U}(\tilde{\jmath}_{i}(X))\mathcal{O}_{\exp(U)\widetilde{\gamma}_{i}}(f).$
We recall:
Lemma 7.1.5 ([Fer07, Lem. 3.2.1]).
Let $G/F_{v}$ be a reductive group and $H\in\mathcal{E}_{\mathrm{ell}}(G)$. Then for any choice of Lie algebra transfer factors, there is a quadratic character $\chi_{G,H}:F_{v}^{\times}\to\pm 1$ such that the transfer factors satisfy:
$\Delta(\alpha X_{H},\alpha X_{G})=\chi_{G,H}(\alpha)\Delta(X_{H},X_{G})$
for all $\alpha\in F_{v}^{\times}$.
In particular, since the quadratic character $\chi_{G,H}$ vanishes on $\alpha^{2}$, the summand in (7.1.6) for a single pair $(i,U)$ scales with $\alpha$ like:
(7.1.7) $\frac{D^{\mathfrak{g}_{\widetilde{\gamma}}}(\alpha^{2}\tilde{\jmath}(X))^{1/2}}{D^{\mathfrak{h}_{\gamma}}(\alpha^{2}X)^{1/2}}\Delta^{\mathfrak{g}_{\widetilde{\gamma}_{i},\mathrm{sc}}}_{\bar{\mathfrak{h}}_{i}}(\alpha^{2}X_{\bar{H}_{i}},\alpha^{2}\tilde{\jmath}_{i}(X)_{\mathrm{sc}})\Gamma^{\mathfrak{g}_{\widetilde{\gamma}_{i}}}_{U}(\alpha^{2}\tilde{\jmath}_{i}(X))\\ =C_{i,U}|\alpha|^{-\dim\mathcal{O}_{U}+\dim G_{\tilde{\gamma}}-\dim H_{\gamma}}$
for some constant $C_{i,U}$ and small enough $\alpha$.
7.1.4. Results
As a consequence that is surely known to experts but does not seem to be written down anywhere, we get the following key input for the vanishing of the main term in our global trace formula expressions:
Proposition 7.1.6.
Let $f$ be a test function on $\widetilde{G}_{v}$ and $\gamma,\widetilde{\gamma}_{i}$ as above. Assume that for all $i$ and all nilpotent orbits $U$ of $\mathfrak{g}_{\tilde{\gamma}_{i}}$,
$\dim\mathcal{O}_{U}=\dim G_{\widetilde{\gamma}}-\dim H_{\gamma}\implies\mathcal{O}_{\exp(U)\widetilde{\gamma}_{i}}(f)=0.$
Then $\mathcal{SO}_{\gamma}(f^{H})=0$.
(Note that if $\dim G_{\widetilde{\gamma}}=\dim H_{\gamma}$, we only need to check that $\mathcal{O}_{\widetilde{\gamma}_{i}}(f)=0$.)
Proof.
This comes from comparing the expansion in Lemma 7.1.4 to equations (7.1.6) and (7.1.7) for $X:=X_{H}\in\mathfrak{t}_{H}$. Note that we get much stronger information about the terms in (7.1.7), in particular, that many of them need to vanish. However, we don’t actually need this extra information. ∎
We can also derive a more refined bound which we expect to be useful in a future work studying the level-aspect version of this problem:
Proposition 7.1.7.
Let $f$ be a test function on $\widetilde{G}_{v}$ and $\gamma,\tilde{\gamma}_{i}$ as above. Assume that for all $i$ and all nilpotent orbits $U$ of $\mathfrak{g}_{\tilde{\gamma}_{i}}$ such that
$\dim\mathcal{O}_{U}=\dim G_{\widetilde{\gamma}}-\dim H_{\gamma},$
there is a bound
$|D^{G}(\widetilde{\gamma}_{i})^{1/2}\mathcal{O}_{\exp(U)\widetilde{\gamma}_{i}}(f)|\leq M.$
Then we can bound
$|D^{H}(\gamma)^{1/2}\mathcal{SO}_{\exp(X)\gamma}(f^{H})|\leq MC_{G_{v},H_{v}}$
for some constant $C_{\tilde{G}_{v},H_{v}}$ depending only on $v,\tilde{G},H$.
Proof.
Let $\mathcal{U}$ be the set of such $(i,U)$. Comparing Lemma 7.1.4 and (7.1.6),
$|D^{H}(\gamma)^{1/2}\mathcal{SO}_{\exp(X)\gamma}(f^{H})|\leq M\left|\sum_{(i,U)\in\mathcal{U}}C_{i,U}(f)\right|,$
where the $C_{i,U}$ are as in (7.1.7). The number of terms in the sum can be bounded by a constant depending only on $G$ so it suffices to bound each term. Furthermore, since local fields have finitely many extensions of each degree, there are only finitely many possibilities for the pair $(H_{\gamma},G_{\tilde{\gamma}_{i}})$ so it suffices to allow the $C_{i,U}$ bound to also depend on these centralizers.
For our special $(i,U)\in\mathcal{U}$ where (7.1.7) is locally constant, we compute $C_{i,U}$ by plugging in a small enough fixed $\alpha_{0}^{2}X_{0}$. “Small enough” depends on the neighborhood $U$ in Theorem 7.1.2 and the validity of the Shalika germ expansions. All these only depends on $H_{\gamma},G_{\tilde{\gamma}_{i}},H,G$. In other words, there is a small enough $\alpha_{0}^{2}X_{0}$ depending only on $H_{\gamma},G_{\tilde{\gamma}_{i}},H,G$. This finishes the argument. ∎
Finally, we get an exact formula in special cases—a twisted version of [LS07, Lemma 2.4.A]:
Corollary 7.1.8.
Let $f$ be a test function on $\widetilde{G}_{v}$ and $\gamma,\tilde{\gamma}_{i}$ as above. Assume that $\dim G_{\widetilde{\gamma}}=\dim H_{\gamma}$. Then
$\mathcal{SO}_{\gamma}(f^{H})=\sum_{i}e(G_{\widetilde{\gamma}_{i}})\Delta^{\widetilde{G}}_{H}(\gamma,\tilde{\gamma}_{i})\mathcal{O}_{\widetilde{\gamma}_{i}}(f),$
where $e(\cdot)$ is the Kottwitz sign (and recalling the definition of transfer factors at these potentially non-regular elements from Corollary 7.1.3).
Proof.
We again compare Lemma 7.1.4 to (7.1.6). The argument of [LS07, 2.4.A] exactly carries over after replacing their discussion involving 1.6.A with Corollary 7.1.3. We point out that the signs $e(G_{\widetilde{\gamma}_{i}})$ again come from [Rog81] through another use of the aforementioned argument in [Kot88, §3] and that the right-hand side should be though of as the “correct” definition of a non-regular $\kappa$-orbital integral. ∎
7.2. Application to $(\widetilde{E}^{\infty}_{\mathfrak{n}})^{G}$: Self-Dual Case
We apply the result of 7.1.1 to show the vanishing at central elements of the function $(\widetilde{E}^{\infty}_{\mathfrak{n}})^{G}$ defined in 4.3.
In this section, we will only study the self-dual case with $N$ even and $G=\mathrm{SO}_{N+1}\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ where root numbers are most interesting. The conjugate self-dual case is significantly more computationally involved and will be put off until next section.
Note first that the class $1\rtimes\theta\in\widetilde{G}_{N}$ is the one that transfers to $1\in G$ since $\theta$ preserves the standard pinning. To eventually use Proposition 7.1.6, we note:
$(G_{N})_{1\rtimes\theta}=\begin{cases}\mathrm{Sp}_{N}&N\text{ even}\\ \mathrm{O}_{N}&N\text{ odd}\end{cases}.$
Furthermore, when $N$ is even, the stable class of $1\rtimes\theta$ is the same as its rational class since $\mathrm{Sp}_{N}$ is semisimple and simply connected implying that $H^{1}(F_{v},(G_{N})_{1\rtimes\theta})=0$ by a standard result on $p$-adic groups.
7.2.1. Local Result
We first study the indicator functions of cosets used to build the local factors of $\widetilde{E}^{\infty}_{\mathfrak{n}}$:
Proposition 7.2.1.
Let $N$ be even, and choose $G=\mathrm{SO}_{N+1}\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ (i.e, $\widehat{G}\cong\mathrm{Sp}_{N}$). Let $v$ be a place of $F$, choose $k>0$, and let
$\tilde{f}_{v,k}:=(J_{v,k}^{-1}w_{N}\rtimes\theta)\bar{\mathbf{1}}_{K_{1,v}(k)}.$
Then for all $\gamma\in Z_{G_{v}}$, $\tilde{f}_{v,k}^{G}(\gamma)=0$.
Proof.
First, $\tilde{f}_{v,k}$ is the indicator function of the coset
$K_{1,v}(k)(J_{v,k}^{-1}w_{N}\rtimes\theta)=K_{1,v}(k)\begin{pmatrix}&&&\pm\varpi_{v}^{-k}\\ &&\iddots&\\ &\pm\varpi_{v}^{-k}&&\\ 1&&&\end{pmatrix}\rtimes\theta.$
Elements of this coset have a non-zero entry in the bottom-left corner since elements of $K_{1,v}(k)$ necessarily have a non-zero entry in the bottom-right when $k>0$.
However, the identity on $G$ transfers to the stable class of $1\rtimes\theta$ in $\widetilde{G}_{N}$ and all elements in this class are of the form $A\rtimes\theta$ with
$w_{N}A^{T}w_{N}^{-1}=A.$
In particular, since $N$ is even, the signs in $w_{N}$ force all elements in the stable class of $1\rtimes\theta$ to have bottom left corner $a$ satisfying $a=-a\implies a=0$. Therefore our coset doesn’t intersect the stable class that transfers to the identity so $\mathcal{O}_{1\rtimes\theta}(\tilde{f}_{v,k})=0$.
We then apply Proposition 7.1.6. For the identity element, $\dim G_{1\rtimes\theta}=\dim H$ so we only need to vanishing at $U=0$ which follows by the above argument. Therefore $\mathcal{SO}_{1}(\tilde{f}_{v,k}^{H})=\tilde{f}_{v,k}^{H}(1)=0$. Since $\mathrm{SO}_{N+1}$ has trivial center, we are done. ∎
7.2.2. Global Result
We can then put these together to get a global result:
Corollary 7.2.2.
Choose $G=\mathrm{SO}_{N+1}\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ (i.e, $\widehat{G}\cong\mathrm{Sp}_{N}$). Choose conductor
$\mathfrak{n}=\prod_{v}\mathfrak{p}_{v}^{v(\mathfrak{n})}.$
If there exists finite place $v$ satisfying either:
  • $v(\mathfrak{n})$ is odd,
  • $v(\mathfrak{n})>N$,
then $(\widetilde{E}^{\infty}_{\mathfrak{n}})^{G}(1)=0$. Otherwise,
$(\widetilde{E}^{\infty}_{\mathfrak{n}})^{G}(1)=\prod_{v}(-1)^{v(\mathfrak{n})/2}\binom{N/2}{v(\mathfrak{n})/2}.$
Proof.
First, $\widetilde{E}^{\infty}_{\mathfrak{n}}$ has $\widetilde{E}_{v,v(\mathfrak{n})}$ as its factor at $v$. By Proposition 4.3.2, $\widetilde{E}_{v,v(\mathfrak{n})}$ is a linear combination of the $\tilde{f}_{v,k}$ from Proposition 7.2.1 with $\tilde{f}_{v,0}$ appearing with coefficient 00 if the stated conditions on $v$ hold, and coefficient $(-1)^{v(\mathfrak{n})/2}\binom{N/2}{v(\mathfrak{n})/2}$ otherwise.
Next, Proposition 7.2.1 gives that for $k>0$, $\tilde{f}_{v,k}^{H}(\gamma)=0$. The twisted fundamental lemma gives that $\tilde{f}_{v,0}^{H}(\gamma)=\bar{\mathbf{1}}_{K_{v}\rtimes\theta}^{H}(1)=\bar{\mathbf{1}}_{K_{H,v}}(1)=1$. The result follows. ∎
Note.
When $N$ is odd or $G\cong\mathrm{SO}_{N,\eta}$, $\widetilde{E}^{\infty}_{\mathfrak{n}}$ has positive trace against every cuspidal $A$-packet on $G$. Therefore, by Proposition 9.2.2 and Lemma 9.3.2, we instead have that both
$(\widetilde{E}^{\infty}_{\mathfrak{n}})^{G}(1)\pm(\widetilde{E}^{\infty}_{\mathfrak{n}})^{G}(-1)>0.$
so the above computation needs to somehow fail to not contradict.
When $N$ is odd, we can see this in the failure of the argument of Proposition 7.1.6 due to the negative sign for the bottom-left corner becoming positive. When $N$ is even and $\widehat{G}$ is orthogonal, $G$ no longer has the same dimension as $(G_{N,v})_{\theta}$, so the formula of Corollary 7.1.8 doesn’t apply and non-trivial unipotent orbital integrals in $\widetilde{G}_{N,v}$ contribute to central values of $(\widetilde{E}^{\infty}_{\mathfrak{n}})^{G}$.
8. Application to $(\widetilde{E}^{\infty}_{\mathfrak{n}})^{G}$: Conjugate Self-Dual Case
The conjugate self-dual case is significantly more complicated. As preliminary computations:
8.1. Twisted Conjugacy Classes
8.1.1. Transfers
As in the self-dual case, the stable class of $1\rtimes\bar{\theta}\in\widetilde{G}_{N}$ transfers to the identity in any $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ since $\bar{\theta}$ fixes the standard pinning. Also,
$(G_{N})_{1\rtimes\bar{\theta}}=U^{E/F}_{N}.$
As in the generalization after [KS99, 5.1.C], we can compute transfers from $Z_{G_{N}}(F_{v})\rtimes\bar{\theta}$ to $Z_{G}(F_{v})$:
$\beta\rtimes\bar{\theta}\mapsto\beta/\bar{\beta},$
implicitly using $Z_{G_{N}}(F_{v})=E_{v}^{\times}$ and $Z_{G}(F_{v})=U_{1}(E_{v})$. This map is surjective onto $Z_{G}(F_{v})$ by Hilbert 90.
8.1.2. The Stable Class of the Identity
We compute how the stable “identity class” $1\rtimes\bar{\theta}$ decomposes into rational classes. We recall (e.g. [Lab11, §III.3]) that for $\gamma$ semisimple, the rational classes inside the stable class of $\gamma$ are in bijection with
$\ker(H^{1}(F,G^{0}_{\gamma})\to H^{1}(F,G)).$
Applying this formula, consider the non-trivial case when $v$ is non-split. Then
$\ker(H^{1}(F_{v},U^{E_{w}/F_{v}}_{N})\to H^{1}(F_{v},\operatorname{Res}_{F_{v}}^{E_{w}}\mathrm{GL}_{N}))=H^{1}(F_{v},U^{E_{w}/F_{v}}_{N})=\mathbb{Z}/2\mathbb{Z}$
by Shapiro’s lemma, Hilbert 90, and Kottwitz’s formula from [Kot86] for the unitary cohomology group. Therefore, the stable conjugacy class of $1\rtimes\bar{\theta}$ breaks up into two rational conjugacy classes.
We will need to distinguish the rational classes. Working over $\overline{F}_{v}$, $\widetilde{G}_{N}(\overline{F}_{v})\cong(\mathrm{GL}_{N}\times\mathrm{GL}_{N})(\overline{F}_{v})$ and conjugation acts by swapping the coordinates, so the stable conjugacy class over $F_{v}$ consists of all elements of the form $B\rtimes\bar{\theta}$ with $B\in G_{N}(F_{v})\cong\mathrm{GL}_{N}(E_{w})$ and
$B=\bar{\theta}(B)^{-1}\;(=w_{N}\bar{B}^{T}w_{N}).$
In particular, $\det B=\det\bar{B}$, so $B\rtimes\theta\mapsto\det(B)$ is a map from the stable conjugacy class to $F_{v}^{\times}$. We can also compute that all values of this invariant can be realized by matrices in the stable class (e.g, pick an appropriate anti-diagonal matrix).
Since
$A(B\rtimes\bar{\theta})A^{-1}=ABw_{N}\bar{A}^{T}w_{N}\rtimes\bar{\theta},$
rational conjugacy can only scale the above determinant invariant by elements of $N_{E_{w}/F_{v}}(E_{w}^{\times})$. In total,
$B\rtimes\bar{\theta}\mapsto\det(B)\in F_{v}^{\times}/N_{E_{w}/F_{v}}(E_{w}^{\times})\cong\mathbb{Z}/2\mathbb{Z}$
is a complete invariant determining the rational conjugacy class of $B\rtimes\bar{\theta}$.
For $B\rtimes\bar{\theta}$ in the stable class of arbitrary $\gamma\rtimes\bar{\theta}$ for $\gamma\in Z_{G_{N,v}}$, we can instead use the invariant $\det(B)/\gamma^{N}\in F_{v}^{\times}/N_{E_{w}/F_{v}}(E_{w}^{\times})$.
8.1.3. Transfer Factor Lemmas
We finally need some information about transfer factors for $G=U^{+}_{N}\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$:
Lemma 8.1.1.
Consider $G=U^{+}_{N}\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ and $e\rtimes\bar{\theta}$ a representative for the other rational class in the stable class of $1\rtimes\bar{\theta}$. Then for any choice of transfer factors:
$\Delta^{\widetilde{G}_{N}}_{G}(1,1\rtimes\bar{\theta})=\Delta^{\widetilde{G}_{N}}_{G}(1,e\rtimes\bar{\theta}).$
Proof.
Any strongly regular $\gamma\rtimes\bar{\theta}$ close enough to $1\rtimes\bar{\theta}$ would have stable conjugate $\gamma^{\prime}\rtimes\bar{\theta}$ close to $e\rtimes\bar{\theta}$. By Corollary 7.1.3, it suffices to show the result for all such pairs $\gamma\rtimes\bar{\theta}$ and $\gamma^{\prime}\rtimes\bar{\theta}$ in some small enough neighborhood.
For this, we apply [KS99, 5.1.D]. The hypercocycle $\kappa_{T}$ is determined by the cocycle $a_{T}$ defined on pp. 39-40 in loc. cit. Using notation therein, in the case of $H=U^{+}_{N}$, we have that $\widehat{H}=\widehat{G}_{1}$ exactly. Therefore $a_{T}$ is trivial so $\kappa_{T}$ is as well. The result follows. ∎
Lemma 8.1.2.
Consider $G=U^{+}_{N}\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$, and semisimple $\gamma\rtimes\bar{\theta}\in\widetilde{G}_{N,v}$ and $\delta\in G_{v}$. Then, for all $a\in E^{\times}=Z_{G_{N,v}}$,
$\Delta^{\widetilde{G}_{N}}_{G}(\delta,\gamma)=\Delta^{\widetilde{G}_{N}}_{G}((a/\bar{a})\delta,a\gamma).$
Proof.
This follows from the generalization after [KS99, 5.1.C]. The character $\lambda_{C}$ there must be trivial to be consistent with Lemma 5.2.4, see [Mok15, Rmk 3.2.3].
Note that for the non-strongly regular case, we prove this first for strongly regular classes close to $\gamma$ and $\delta$ and then use Corollary/Definition 7.1.3. ∎
8.2. Number-Theoretic Lemmas
We also need two number-theoretic lemmas. Beware an indexing technicality: at places $v$ of $F$ that ramify in $E$ to places $w$, we state all results in terms of powers of $\mathfrak{p}_{w}$. Let $\mathfrak{D}_{E_{w}/F_{v}}$ be the different ideal.
Lemma 8.2.1.
Define the map
$\phi:\mathcal{O}_{E_{w}}^{\times}\to\mathcal{O}_{E_{w}}^{\times},\quad\phi(x)=\frac{x}{\bar{x}}.$
Then
$\phi(1+\mathfrak{p}_{w}^{k})\subseteq 1+\mathfrak{D}_{E_{w}/F_{v}}\cap\mathfrak{p}_{w}^{k}$
Furthermore, if $E_{w}/F_{v}$ is at most tamely ramified or $k=0$, then
$\phi(1+\mathfrak{p}_{w}^{k})=(1+(\mathfrak{p}_{w}^{k}\cap\mathfrak{D}_{E_{w}/F_{v}}))_{Nm=1}.$
Proof.
For all $x\in\mathcal{O}^{\times}_{E_{w}}$, we have $\phi(x)\equiv 1\mod\mathfrak{D}_{E_{w}/F_{v}}$ since $x-\bar{x}\in\mathfrak{D}_{E_{w}/F_{v}}$ by [Ser13, IV.2]. This gives the inclusion.
For the equality, let $j\geq 0$ be such that $\mathfrak{D}_{E_{w}/F_{v}}=\mathfrak{p}_{w}^{j}$. It suffices to show that $H^{1}(\operatorname{Gal}(E_{w}/F_{v}),1+\mathfrak{p}_{w}^{k})$ vanishes for all $k\geq j$. If $E_{w}$ is unramified, then $\mathfrak{D}_{E_{w}/F_{v}}=\mathcal{O}_{E_{w}}$, and the vanishing for all $k$ follows from [NSW13, VII.7.1.2]. This lemma also gives the desired vanishing tamely ramified case for $k\geq 1$ (since then $\mathfrak{D}_{E_{w}/F_{v}}=\mathfrak{p}_{w}$).
It remains to check the case when $v$ is wildly ramified and $k=0$: that $\phi(\mathcal{O}_{E_{w}}^{\times})=(1+(\mathfrak{p}_{w}^{k}\cap\mathfrak{D}_{E_{w}/F_{v}}))_{Nm=1}$. This is equivalent to checking that in the long exact sequence associated to
$0\to 1+\mathfrak{D}_{E_{w}/F_{v}}\to\mathcal{O}^{\times}_{E_{w}}\to\mathcal{O}^{\times}_{E_{w}}/(1+\mathfrak{D}_{E_{w}/F_{v}})\to 0,$
the natural map $H^{1}(\operatorname{Gal}(E_{w}/F_{v}),1+\mathfrak{D}_{E_{w}/F_{v}})\to H^{1}(\operatorname{Gal}(E_{w}/F_{v}),\mathcal{O}^{\times}_{E_{w}})$ is trivial. We check instead that
(8.2.1) $H^{1}(\operatorname{Gal}(E_{w}/F_{v}),\mathcal{O}^{\times}_{E_{w}})\hookrightarrow H^{1}(\operatorname{Gal}(E_{w}/F_{v}),\mathcal{O}^{\times}_{E_{w}}/(1+\mathfrak{D}_{E_{w}/F_{v}})).$
A computation with the long exact sequence associated to $\mathcal{O}_{E_{w}}^{\times}\hookrightarrow E_{w}^{\times}\twoheadrightarrow\mathbb{Z}$ shows that $H^{1}(\operatorname{Gal}(E_{w}/F_{v}),\mathcal{O}^{\times}_{E_{w}})\simeq\mathbb{Z}/2\mathbb{Z}$, generated by the class of $\varpi/\overline{\varpi}$. To get (8.2.1), it suffices to show that this class survives under reduction modulo $1+\mathfrak{D}_{E_{w}/F_{v}}=:1+\mathfrak{p}_{w}^{j}$. Since $\varpi/\overline{\varpi}\in(1+\mathfrak{p}_{w}^{j-1})\setminus(1+\mathfrak{p}_{w}^{j})$ [Ser13, IV.1-2], the cocycle is nontrivial in the quotient. Furthermore, since $\mathcal{O}^{\times}_{E_{w}}/(1+\mathfrak{D}_{E_{w}/F_{v}})$ is a trivial Galois module, any coboundary is trivial. In total, $\varpi/\overline{\varpi}$ represents a nontrivial class. ∎
Lemma 8.2.2.
Let $N$ be odd. There is $A\in\mathrm{GL}_{N}(\mathcal{O}_{E_{w}})$ such that $w_{N}\bar{A}^{T}w_{N}^{-1}=yA$ if and only if $N_{E_{w}/F_{v}}(y)=1$ and either:
  1. (1)
    $E_{w}/F_{v}$ is unramified,
  2. (2)
    $E_{w}/F_{v}$ is tamely ramified and $y^{N}\not\equiv-1\pmod{\mathfrak{p}_{w}}$,
  3. (3)
    $E_{w}/F_{v}$ is wildly ramified and $y^{N}\notin 1+\mathfrak{p}_{w}$ or $y\in 1+\mathfrak{D}_{E_{w}/F_{v}}$.
Proof.
We instead construct $B=w_{N}A\in\mathrm{GL}_{N}(\mathcal{O}_{E_{w}})$, which would have to satisfy:
$\bar{B}^{T}=\bar{A}^{T}w_{N}^{-1}=w_{N}^{-1}w_{N}\bar{A}^{T}w_{N}^{-1}=(-1)^{N-1}w_{N}yA=yB.$
If $y$ can be written as $\alpha/\overline{\alpha}$ for $\alpha\in\mathcal{O}_{E}^{\times}$, then $B$ can be chosen to be diagonal with $\alpha$’s on the diagonal.
If not, then any such $B$ has only zeros on the diagonal. Let its entries be $b_{ij}$ so that
$\det B=\sum_{\sigma}\operatorname{sgn}(\sigma)\prod_{i}b_{i\sigma(i)},$
where by the assumption on $B$, we can sum over just permutations that don’t have fixed points. Comparing terms for $\sigma$ and $\sigma^{-1}$,
$\operatorname{sgn}(\sigma^{-1})\prod_{i}b_{i\sigma^{-1}(i)}=\operatorname{sgn}(\sigma)\prod_{i}b_{\sigma(i)i}=\operatorname{sgn}(\sigma)y^{N}\prod_{i}\bar{b}_{i\sigma(i)}.$
Since $N$ is odd, all permutations with no fixed points aren’t equal to their inverses so we can therefore pair them up and write $\det B$ in the form $R+y^{N}\bar{R}$ where $R$ can be freely chosen to be any element of $\mathcal{O}_{E_{w}}$.
If $y^{N}\not\equiv-1\pmod{\mathfrak{p}_{v}}$, this can be arranged to be in $\mathcal{O}_{E_{w}}^{\times}$ by choosing $R\in\mathcal{O}_{F_{v}}^{\times}$. Otherwise, there is $R$ such that $R+y^{N}\bar{R}\in\mathcal{O}_{E_{w}}^{\times}$ if and only if there is $R$ such that $R/\bar{R}\not\equiv 1\pmod{\mathfrak{p}_{v}}$. Summarizing, the possible $y$’s are exactly those of norm $1$ for which at least one of the following hold:
  1. (1)
    There is $R\in\mathcal{O}_{E_{w}}^{\times}$ such that $R/\bar{R}\not\equiv 1\pmod{\mathfrak{p}_{w}}$,
  2. (2)
    $y$ can be written in the form $\alpha/\overline{\alpha}$ for $\alpha\in\mathcal{O}_{E_{w}}^{\times}$,
  3. (3)
    $y^{N}\not\equiv-1\pmod{\mathfrak{p}_{w}}$.
As in Lemma 8.2.1, (1) holds if and only if $E_{w}/F_{v}$ is unramified. For $E_{w}/F_{v}$ tamely ramified, (2) implies that $y\equiv 1\pmod{\mathfrak{p}_{w}}$ which together with $\mathfrak{p}_{v}\nmid 2$ implies (3).
If $E_{w}/F_{v}$ is wildly ramified, then $\mathfrak{p}_{w}|2$ so (3) is equivalent to $y^{N}\notin 1+\mathfrak{p}_{w}$. Also (2) in our setting is equivalent to $y\in 1+\mathfrak{D}_{E_{w}/F_{v}}$ by Lemma 8.2.1. ∎
8.3. Local Results
Now we are ready for the main results. We again study coset indicators $\tilde{f}_{v,k}$ used to build up $\widetilde{E}^{\infty}_{\mathfrak{n}}$ in Corollary 4.3.1. Beware that we use local indexing for $\tilde{f}_{v,k}$; see §3.1.2.
8.3.1. Non-Intersection of Support
Proposition 8.3.1.
In the conjugate self-dual case, let $N$ be even and $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$. Let $v$ be a place of $F$ that doesn’t split in $E$ and $w$ the corresponding place of $E$. Let
$\tilde{f}_{v,k}:=(J_{v,k}^{-1}w_{N}\rtimes\bar{\theta})\bar{\mathbf{1}}_{K_{1,v}(k)}.$
Then for all $\gamma\in Z_{G_{v}}\subseteq E_{w}^{\times}$, the support of $\tilde{f}_{v,k}^{G}$ intersects the stable conjugacy class of $\gamma$ only if:
  1. (1)
    $-\gamma\in\phi(1+\mathfrak{p}_{w}^{k})$ (see Lemma 8.2.1), and:
  2. (2)
    either:
    1. (a)
      $E_{w}/F_{v}$ is unramified,
    2. (b)
      $E_{w}/F_{v}$ is ramified and $k$ is even.
Proof.
As explained in 8.1.1, the element $\gamma\in U_{E/F}(1)_{v}=Z_{G_{v}}$ transfers to the stable conjugacy class $\beta\rtimes\bar{\theta}$ of $\widetilde{G}_{N}$ with $\beta/\bar{\beta}=\gamma$. This stable class consists exactly of elements of the form
(8.3.1) $\beta A\rtimes\bar{\theta}\text{ with }w_{N}\overline{A}^{T}w_{N}^{-1}=A.$
We want to compare this stable class with the support of $\tilde{f}_{v,k}$, which consists of all elements of the form
$X(J_{v,k}^{-1}w_{N}\rtimes\bar{\theta})=X\begin{pmatrix}&&&\pm\varpi_{v}^{-k}\\ &&\iddots&\\ &\pm\varpi_{v}^{-k}&&\\ 1&&&\end{pmatrix}\rtimes\bar{\theta}$
for $X\in K_{1,v}(k)$. We write $X$ in block matrix form
(8.3.2) $X=\begin{pmatrix}X_{0}&\vec{v}_{1}\\ \vec{v}_{2}&x\end{pmatrix}\iff X(J_{v,k}^{-1}w_{N}\rtimes\bar{\theta})=\begin{pmatrix}\vec{v}_{1}&\varpi_{v}^{-k}X_{0}w_{N-1}\\ x&\varpi_{v}^{-k}\vec{v}_{2}w_{N-1}\end{pmatrix}\rtimes\bar{\theta}.$
We now assume that there exists such an element $X(J_{v,k}^{-1}w_{N}\rtimes\bar{\theta})$ in the stable class of $\beta\rtimes\bar{\theta}$ and derive restrictions on $\gamma$.
Necessary Condition 1:
As in the proof of Proposition 7.2.1, we have a restriction from the bottom-left matrix entry $x$ of $X(J_{v,k}^{-1}w_{N}\rtimes\theta)$. Since $N$ is even, condition (8.3.1) gives that $x=\beta\alpha$ for $\alpha$ satisfying $\overline{\alpha}=-\alpha$. However, we also have $x\equiv 1\pmod{\mathfrak{p}_{w}^{k}}$. Therefore, the class of $\beta\rtimes\theta$ only intersects the support of $f_{k}$ if there is $\alpha\in E_{w}$ such that:
(8.3.3) $\bar{\alpha}=-\alpha,\quad\alpha\beta\in\mathcal{O}_{E_{w}},\quad\text{and}\quad\alpha\beta\equiv 1\pmod{\mathfrak{p}_{w}^{k}}.$
This is equivalent to something simpler: if $\alpha,\beta$ satisfy conditions (8.3.3), then there is $h=\alpha\beta-1\in\mathfrak{p}_{w}^{k}$ such that
$\gamma=-(1+h)/(1+\bar{h}).$
Conversely, given $h\in\mathfrak{p}_{w}^{k}$, $\gamma$ satisfying the above equation, and $\beta$ such that $\beta/\bar{\beta}=\gamma$, setting $\alpha=(1+h)/\beta$ satisfies conditions (8.3.3). Finally, by Lemma 8.2.1, the $\gamma=-(1+h)/(1+\bar{h})$ condition is further equivalent to
(8.3.4) $-\gamma\in\phi(1+\mathfrak{p}_{w}^{k})$
for the map $\phi$ of Lemma 8.2.1. This is condition (1).
Necessary Condition 2:
We also have a restriction from the top-right $(N-1)\times(N-1)$ block in the right-hand side of (8.3.2). Condition (8.3.1) requires that
$w_{N-1}\overline{(\beta^{-1}\varpi^{-k}X_{0}w_{N-1})}^{T}w_{N-1}^{-1}=-(\beta^{-1}\varpi^{-k}X_{0}w_{N-1}),$
where the minus sign appears since “restricting” the conjugation by $w_{N}$ to the block doesn’t exactly give $w_{N-1}$. In particular $X_{0}w_{N-1}$ is an eigenvector for the conjugate-linear involution on $GL_{N_{1}}(\mathcal{O}_{E_{w}})$ given by $Y\mapsto w_{N-1}\overline{Y}^{T}w_{N-1}^{-1}$, with eigenvalue
$-(\overline{\beta}/\beta)(\overline{\varpi}^{k}/\varpi^{k})=-\gamma^{-1}(\overline{\varpi}^{k}/\varpi^{k}).$
Assume that $k>0$. Since $\vec{v}_{2}\equiv 0\pmod{\mathfrak{p}_{w}^{k}}$ and $X$ is invertible over $\mathcal{O}_{E_{w}}$, the matrix $X_{0}$, and therefore $X_{0}w_{N-1}$ must also be invertible over $\mathcal{O}_{E_{w}}$.
In particular, $-\gamma^{-1}(\overline{\varpi}^{k}/\varpi^{k})$ satisfies the conditions of Lemma 8.2.2.
Sufficiency
In the converse direction, if $\gamma$ and $\beta$ satisfy (8.3.3) and $y=-\gamma^{-1}(\overline{\varpi}/\varpi)$ satisfies the conditions of Lemma 8.2.2 for $\mathrm{GL}_{N-1}(\mathcal{O}_{E_{w}})$, we can choose $X_{0}w_{N-1}\in\mathrm{GL}_{N-1}(\mathcal{O}_{E_{w}})$ satisfying
$w_{N-1}\overline{(X_{0}w_{N-1})^{T}}w_{N-1}^{-1}=-(\overline{\beta}/\beta)(\overline{\varpi}^{k}/\varpi^{k})(X_{0}w_{N-1}).$
Then choosing $\alpha$ as above, the matrix
$A=\begin{pmatrix}&\varpi^{-k}\beta^{-1}X_{0}w_{N-1}\\ \alpha&\end{pmatrix}$
satisfies $w_{N}\bar{A}^{T}w_{N}^{-1}=A$ and $\beta A\rtimes\bar{\theta}$ is in the support of $\tilde{f}_{v,k}$.
Simplification of Condition 2
Finally, (8.3.4) lets us simplify condition 2. By Lemma 8.2.1, condition (8.3.4) gives that $-\gamma^{-1}(\varpi/\overline{\varpi})^{-k}\equiv(\varpi/\overline{\varpi})^{-k}$ mod both $\mathfrak{p}_{w}$ and $\mathfrak{D}_{E_{w}/F_{v}}$. This congruence class mod either ideal is independent of the choice of $\varpi$ since different choices make the quotient vary only by $\alpha/\bar{\alpha}\equiv 1\pmod{\mathfrak{D}_{E_{w}/F_{v}}}$ for $\alpha\in\mathcal{O}_{E_{w}}$. Our desired conditions from Lemma 8.2.2 therefore become
  • If $v$ is unramified: nothing
  • If $v$ is tamely ramified: $(\varpi/\overline{\varpi})^{-(N-1)k}\equiv 1\pmod{\mathfrak{p}_{w}}$
  • If $v$ is wildly ramified: $(\varpi/\overline{\varpi})^{-(N-1)k}\not\equiv 1\pmod{\mathfrak{p}_{w}}$ or $(\varpi/\overline{\varpi})^{-k}\equiv 1\pmod{\mathfrak{D}_{E_{w}/F_{v}}}$.
A tamely ramified quadratic extension is generated by a square root of a uniformizer, so in this case $\varpi/\overline{\varpi}\equiv-1\pmod{\mathfrak{p}_{w}}$. Therefore, since $N-1$ is odd, the condition in Lemma 8.2.2 becomes that $k$ is even.
In the wildly ramified case, we note since $\mathfrak{D}_{E_{w}/F_{v}}$ is exactly generated by $\varpi-\overline{\varpi}$, we have that $\varpi/\overline{\varpi}$ is a non-trivial element of $(1+\varpi^{-1}\mathfrak{D}_{E_{w}/F_{v}})/(1+\mathfrak{D}_{E_{w}/F_{v}})$. In particular $(\varpi/\overline{\varpi})\equiv 1\pmod{\mathfrak{p}_{w}}$ always. On the other hand, since $v|2$, the quotient is a $2$-group so we again have that $(\varpi/\overline{\varpi})^{-k}\in 1+\mathfrak{D}_{E_{w}/F_{v}}$ holds if and only if $k$ is even. ∎
8.3.2. Orbital Integral Vanishing
Our support result immediately gives:
Corollary 8.3.2.
Consider the setup of Proposition 8.3.1. For all $\gamma\in Z_{G_{v}}$ that do not satisfy the conditions therein, we have $f^{G}_{v,k}(\gamma)=0$.
Proof.
Since $\dim(G_{v})=\dim((G_{N,v})_{\gamma\rtimes\theta})$, this follows from Proposition 7.1.6. ∎
In the case of $G=U^{+}_{N}$, we can say more:
Proposition 8.3.3.
Consider the setup of Proposition 8.3.1 and specialize to the case where $G=U_{N}^{+}$. Assume that $\gamma$ satisfies the conditions therein. Then:
  1. (1)
    if $\gamma^{\prime}$ also satisfies the conditions of Proposition 8.3.1, then $\tilde{f}^{G}_{v,k}(\gamma)=\tilde{f}^{G}_{v,k}(\gamma^{\prime})$,
  2. (2)
    if $v$ is unramified, $f_{v,k}^{G}(\gamma)$ is non-zero with sign $(-1)^{k}$,
  3. (3)
    if $v$ is ramified, $f_{v,k}^{G}(\gamma)=0$ whenever $k>0$.
Proof.
We prove the points in order:
Constancy
For $i=1,2$, consider $\gamma_{i}\in Z_{G_{v}}$ satisfying the conditions of Proposition 8.3.1 and let $\gamma_{i}$ be a norm of $\beta_{i}\rtimes\bar{\theta}\in Z_{G_{N,v}}\rtimes\theta$; we can choose $\beta_{i}\equiv a\pmod{\mathfrak{p}_{w}^{k}}$ for some $a\in E^{\times}_{w}$ such that $\bar{a}=-a$. Let $z=\beta_{2}/\beta_{1}$, noting that $z\equiv 1\pmod{\mathfrak{p}_{w}^{k}}$ so that $z\in K_{1,v}(k)$. Let $e=z/\bar{z}=\gamma_{2}/\gamma_{1}$.
First, by Lemma 8.1.2,
$\Delta^{\widetilde{G}}_{G}(e\gamma,z\widetilde{\gamma})=\Delta^{\widetilde{G}}_{G}(\gamma,\widetilde{\gamma}).$
Second, $z\in K_{1,v}(k)$ gives that $\tilde{f}_{v,k}$ is invariant under translation by $z$ so $\mathcal{O}_{\widetilde{\gamma}}(\tilde{f}_{v,k})=\mathcal{O}_{z\widetilde{\gamma}}(\tilde{f}_{v,k})$ for any $\widetilde{\gamma}\in\widetilde{G}_{N,v}$. Let $\beta_{i}^{\prime}\rtimes\bar{\theta}$ be a representative for the other rational class in the stable class of $\beta_{i}\rtimes\bar{\theta}$. Putting everything together, the formula in Corollary 7.1.8 shows that
(8.3.5) $\tilde{f}_{v,k}^{G}(\gamma_{1})\\ =e(G_{\beta_{1}\rtimes\bar{\theta}})\Delta^{\widetilde{G}}_{G}(\gamma_{1},\beta_{1}\rtimes\bar{\theta})\mathcal{O}_{\beta_{1}\rtimes\bar{\theta}}(\tilde{f}_{v,k})+e(G_{\beta_{1}^{\prime}\rtimes\bar{\theta}})\Delta^{\widetilde{G}}_{G}(\gamma_{1},\beta_{1}^{\prime}\rtimes\bar{\theta})\mathcal{O}_{\beta_{1}^{\prime}\rtimes\bar{\theta}}(\tilde{f}_{v,k})\\ =e(G_{\beta_{2}\rtimes\bar{\theta}})\Delta^{\widetilde{G}}_{G}(\gamma_{2},\beta_{2}\rtimes\bar{\theta})\mathcal{O}_{\beta_{2}\rtimes\bar{\theta}}(\tilde{f}_{v,k})+e(G_{\beta_{2}^{\prime}\rtimes\bar{\theta}})\Delta^{\widetilde{G}}_{G}(\gamma_{2},\beta_{2}^{\prime}\rtimes\bar{\theta})\mathcal{O}_{\beta_{2}^{\prime}\rtimes\bar{\theta}}(\tilde{f}_{v,k})\\ =\tilde{f}_{v,k}^{G}(\gamma_{2}).$
This proves the equality we want.
Unramified Calculation
For the unramified calculation, choose again $\gamma\in Z_{G_{v}}$ that is a norm of $\beta\rtimes\bar{\theta}\in Z_{G_{N,v}}\rtimes\theta$ such that the other rational class inside the stable class of $\beta\rtimes\bar{\theta}$ is represented by $\beta^{\prime}\rtimes\bar{\theta}$.
In the unramified case, $N_{E_{w}/F_{v}}(E_{w}^{\times})$ consists of all elements of even valuation. Therefore, by the discussion in §8.1.2, we can without loss of generality choose the conjugation orbit of $\beta\rtimes\bar{\theta}$ to be only supported on elements of even parity while that of $\beta^{\prime}\rtimes\bar{\theta}$ is only on those with odd parity. Note that since $\det\beta$ has even parity, this also implies that $G_{\beta\rtimes\bar{\theta}}$ is the quasisplit inner form of $G_{v}$ while $G_{\beta^{\prime}\rtimes\bar{\theta}}$ is the non-quasisplit one. In particular, since $N$ is even, $e(G_{\beta\rtimes\bar{\theta}})=1$ and $e(G_{\beta^{\prime}\rtimes\bar{\theta}})=-1$.
However, on the support of $\tilde{f}_{v,k}$, the determinant has constant valuation $-k(N-1)$ which has the same parity as $k$. Therefore, only one of the orbital integrals in (8.3.5) is non-zero—the one for $\beta\rtimes\bar{\theta}$ if $k$ is even and that for $\beta^{\prime}\rtimes\bar{\theta}$ if $k$ is odd.
Finally, Lemma 8.1.1 forces the two transfer factors to have the same sign. The consistency of our choice with the twisted fundamental lemma forces this common sign to be positive. Combining this with the above two paragraphs finishes the sign computation.
Ramified Calculation
As in the unramified calculation, choose $\gamma,\beta\rtimes\bar{\theta},\beta^{\prime}\rtimes\bar{\theta}$. This computation is more convenient to do using a different presentation of $\widetilde{G}_{N,v}$. We realize
$\widetilde{G}_{N}(E_{w})\cong\mathrm{GL}_{N}(E_{w})^{2}\rtimes s$
where $s$ is the involution switching the coordinates and $\widetilde{G}_{N,v}$ is embedded through
$g(J^{-1}_{\mathfrak{n}}w_{N}\rtimes\bar{\theta})\mapsto(g,J^{-1}_{\mathfrak{n}}\bar{g}^{-T}J_{\mathfrak{n}})\rtimes s.$
If $E_{w}/F_{v}$ is ramified there is $\epsilon\in\mathcal{O}_{F_{v}}\setminus N_{E_{w}/F_{v}}(E_{w}^{\times})$. Consider the $N\times N$ matrix
$E=\begin{pmatrix}\epsilon&&&\\ &\ddots&&\\ &&\epsilon&\\ &&&1\end{pmatrix}.$
Recall the form of elements in $K_{1,v}(k)(J_{\mathfrak{n}}^{-1}w_{N}\rtimes\bar{\theta})$ given in (8.3.2). If $k\geq 1$, any such element additionally of the form in (8.3.1) must satisfy $\vec{v}_{1}=\vec{v}_{2}=0$ since otherwise their coordinates have different valuations but must also be equal up to sign. In particular, $E$ commutes with all $X$ such that $X(J_{\mathfrak{n}}^{-1}w_{N}\rtimes\bar{\theta})$ is of the form of both (8.3.2) and (8.3.1). Furthermore, $J^{-1}_{\mathfrak{n}}\bar{E}^{-T}J_{\mathfrak{n}}=E$, so in total,
$(E,1)X(J_{\mathfrak{n}}^{-1}w_{N}\rtimes\bar{\theta})(E,1)^{-1}=EX(J_{\mathfrak{n}}^{-1}w_{N}\rtimes\bar{\theta}).$
considering $(E,1)$ as an element of $G_{N}(E_{w})=\mathrm{GL}_{N}(E_{w})\times\mathrm{GL}_{N}(E_{w})$. In particular, this preserves both the conditions of (8.3.2) and (8.3.1)—in other words, conjugation by $(E,1)$ sends the intersection of support of $\tilde{f}_{v,k}$ with the stable orbit of any $\beta\rtimes\bar{\theta}$ for $\beta\in Z_{G,v}$ to itself.
Finally, note that since $N$ is even, the determinants of $XE$ and $X$ descend to the two different classes in $F_{v}^{\times}/N_{E_{w}/F_{v}}(E_{w}^{\times})$ so conjugation by $(E,1)$ also swaps the two rational orbits in the stable orbit. Since the measures on the different centralizers used to define orbital integrals in a stable class are related by conjugation in the algebraic closure, this gives that the two orbital integrals in (8.3.5) are equal. By the same argument as the unramified calculation, this implies that the transfer must vanish. ∎
We need two extra arguments. When $v$ is ramified and $k=0$ we will use the spectral arguments of Section 9:
Lemma 8.3.4.
Consider the same setup as Proposition 8.3.1 and assume additionally that $G=U^{+}_{N}$, $v$ is ramified, and $k=0$. Then for $\gamma\in Z_{G_{v}}$, $\tilde{f}_{v,0}^{G}(\gamma)>0$ whenever $\gamma\equiv-1\pmod{\mathfrak{D}_{E_{w}/F_{v}}}$.
Proof.
By Proposition 8.3.3 (1), it suffices to show that $\tilde{f}_{v,0}^{G}(-1)>0$. However, as in Theorem 4.3.6, the function $\delta_{a}\star\tilde{f}_{v,0}$ is trace-positive since unramified representations have trivial standard root number. Therefore, since conjugate self-dual unramified representations of $\mathrm{GL}_{N}(E_{w})$ exist and are both conjugate-orthogonal and -symplectic, Proposition 9.2.2 shows that $(\delta_{a}\star\tilde{f}_{v,0})^{G}(1)>0$.
Finally, by Lemma 8.1.2, $(\delta_{a}\star\tilde{f}_{v,0})^{G}(1)=\tilde{f}_{v,0}^{G}(a/\bar{a})=\tilde{f}_{v,0}^{G}(-1)$. ∎
Finally, in the case of $v$ split:
Lemma 8.3.5.
Consider the same setup as Proposition 8.3.1 except that $v$ now splits in $E$ and we allow odd $N$. Then we can choose
$(\tilde{f}_{v,k})^{G}=\bar{\mathbf{1}}_{K_{1,v}(k)}.$
In particular
$\tilde{f}_{v,1}^{G}(\gamma)=\begin{cases}\operatorname{vol}(K_{1,v}(k))^{-1}&\gamma\in\mathcal{O}_{E_{w}}^{\times}\text{ and }\gamma\equiv 1\pmod{\mathfrak{p}^{k}_{v}}\\ 0&\text{ else}\end{cases}.$
Proof.
This follows from Lemma 6.1.1(3) in [DGG22] using that $J_{v,k}^{-1}w_{N}\rtimes\theta$ is an involution that normalizes $K_{1,v}(k)$. ∎
8.4. Global Result
We now put these together to get a global vanishing result:
Corollary 8.4.1.
Let $N$ be even and consider $G=U_{N,+}\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$. Fix a conductor
$\mathfrak{n}=\prod_{v}\mathfrak{p}_{v}^{k_{v}}$
globally indexed as in §3.1.2. If there exists a ramified place $v$ of $F$ such that either:
  • $k_{v}$ is half-integral,
  • $k_{v}>N/2$,
then $(\widetilde{E}^{\infty}_{\mathfrak{n}})^{G}(\gamma)=0$ for all $\gamma\in Z_{G}(F)$.
Proof.
Recall the choice of $a\in E^{\times}$ such that $a/\bar{a}=-1$ from Definition 4.3.4. Let $\tilde{f}_{\mathfrak{p}_{v}^{k}}$ be as in Proposition 8.3.1 except globally indexed. Then $\widetilde{E}^{\infty}_{\mathfrak{n}}$ is a linear combination of terms of the form
$\tilde{f}_{\mathfrak{n}^{\prime}}=\prod_{v\text{ split}}\tilde{f}_{\mathfrak{p}_{v}^{k_{v}}}\prod_{v\text{ ns}}\delta_{a^{N-1}}\star\tilde{f}_{\mathfrak{p}_{v}^{k_{v}}}.$
First, note that by Lemma 8.1.2, for all test functions $\widetilde{\varphi}$ on $\widetilde{G}_{N,v}$,
$(\delta_{a^{N-1}}\star\widetilde{\varphi})^{G}(\gamma)=\widetilde{\varphi}^{G}((-1)^{N-1}\gamma).$
Therefore, by Proposition 8.3.3, we have $\tilde{f}_{\mathfrak{n}^{\prime}}^{G}(\gamma)=0$ for all $\gamma\in Z_{G,v}$ if there is a ramified $v$ such that $v(\mathfrak{n}^{\prime})>0$. By Proposition 4.3.2, the bulleted conditions exactly guarantee that all of the $\tilde{f}_{\mathfrak{n}^{\prime}}$ making up $\widetilde{E}^{\infty}_{\mathfrak{n}}$ satisfy this. ∎
Unlike in the self-dual case of Corollary 7.2.2, without the conditions on $\mathfrak{n}$, there can be multiple summands $\tilde{f}_{\mathfrak{n}^{\prime}}$ with non-zero transfers appearing with different signs. We of course expect the terms (8.4.1) for these $\tilde{f}_{\mathfrak{n}^{\prime}}$ to have very different magnitudes and therefore not cancel. However, the techniques established here do not allow us to computing fine enough information on the magnitudes of transfers to show this. We therefore give a partial non-vanishing result which will allow us, in some instances, to describe the bias in signs as the weight varies:
Corollary 8.4.2.
Consider the setup of Corollary 8.4.1 and assume that $\mathfrak{n}$ does not satisfy the conditions therein. Let $\mathfrak{n}_{\mathrm{ur}}$ be the part of $\mathfrak{n}$ without any ramified factors. Then
$\prod_{v\nmid\infty}(-1)^{k_{v}}\sum_{\gamma\in Z_{G}(F)}\omega(\gamma)(\widetilde{E}^{\infty}_{\mathfrak{n}})^{G}(\gamma)>0$
for any character $\omega$ on $Z_{G}\subseteq E^{\times}$ such that $\omega$ is trivial on $(1+\mathfrak{n}_{\mathrm{ur}}\mathfrak{D}_{E/F})\cap\mathcal{O}_{E}^{\times}$ and non-trivial on $(1+\mathfrak{n}^{\prime}\mathfrak{D}_{E/F})\cap\mathcal{O}_{E}^{\times}$ for all proper $\mathfrak{n}^{\prime}|\mathfrak{n}_{\mathrm{ur}}$.
Proof.
Recall the notation in the proof of Corollary 8.4.1, in particular the choice of $a\in E^{\times}$ such that $\bar{a}/a=-1$. As in that proof, by Lemma 8.1.2, all test functions $\widetilde{\varphi}$ on $\widetilde{G}_{N,v}$ satisfy
$(\delta_{a^{N-1}}\star\widetilde{\varphi})^{G}(\gamma)=\widetilde{\varphi}^{G}((-1)^{N-1}\gamma).$
The function $\widetilde{E}^{\infty}_{\mathfrak{n}}$ is a linear combination of $f_{\mathfrak{n}^{\prime}}$ for conductors $\mathfrak{n}^{\prime}|\mathfrak{n}$. If $\mathfrak{n}^{\prime}$ has a ramified divisor, Proposition 8.3.3 (3) and Corollary 8.3.2 give that $f^{G}_{\mathfrak{n}^{\prime}}(\gamma)=0$ for all $\gamma$.
If $\mathfrak{n}^{\prime}\mid\mathfrak{n}$ has no ramified divisors, multiply together the appropriate cases of Corollary 8.3.2, Proposition 8.3.3 (1), and Lemmas 8.3.5 and 8.3.4 over all $v$. This gives that for $\gamma\in Z_{G}$, $\tilde{f}_{\mathfrak{n}^{\prime}}(\gamma)^{G}$ is non-zero and constant for $\gamma\in\mathcal{O}_{E}^{\times}$ with $\gamma\equiv-1\pmod{\mathfrak{n}^{\prime}\mathfrak{D}_{E/F}}$ and zero otherwise (using the $k=0$ case of Lemma 8.2.1 to find the image of $\phi$). Therefore, for such $\mathfrak{n}^{\prime}$
(8.4.1) $\sum_{\gamma\in Z_{G}(F)}\omega(\gamma)\tilde{f}_{\mathfrak{n}^{\prime}}^{G}(\gamma)\neq 0$
whenever $\omega$ is trivial on $(1+\mathfrak{n}^{\prime}\mathfrak{D}_{E/F})\cap\mathcal{O}_{E}^{\times}$ and is 00 otherwise. Furthermore, when this term is non-zero, Proposition 8.3.3(2) shows that it has sign $\prod_{v\text{ ur}}(-1)^{k_{v}}$.
If $\mathfrak{n}$ does not satisfy the conditions of Corollary 8.4.1, then Proposition 4.3.2 says that there is $\tilde{f}_{\mathfrak{n}^{\prime}}$ appearing in the linear combination defining $\widetilde{E}^{\infty}_{\mathfrak{n}}$ with $\mathfrak{n}^{\prime}$ having no ramified divisors. Note that all such $\mathfrak{n}^{\prime}$ divide $\mathfrak{n}_{\mathrm{ur}}$. The conditions on $\omega$ guarantee that (8.4.1) vanishes for all of these $\mathfrak{n}^{\prime}$ except $\mathfrak{n}_{\mathrm{ur}}$ itself. Therefore the result follows after noting that $\tilde{f}_{\mathfrak{n}_{\mathrm{ur}}}$ appears with coefficient $\prod_{v\text{ ram.}}(-1)^{k_{v}}$ in $\widetilde{E}^{\infty}_{\mathfrak{n}}$. ∎
Of course, we expect the full result:
Conjecture 8.4.3.
Corollary 8.4.2 should hold for all $\omega$ that are trivial on $(1+\mathfrak{n}_{\mathrm{ur}}\mathfrak{D}_{E/F})\cap\mathcal{O}_{E}^{\times}$.
9. A Spectral Argument and $(\widetilde{C}^{\infty}_{\mathfrak{n}})^{G}$
We will also need to understand the main term (6.2.1) for $(\widetilde{C}^{\infty}_{\mathfrak{n}})^{G}$. Since $C_{\mathfrak{n}}^{\infty}$ was defined spectrally, this will necessarily happen through a spectral argument:
9.1. A Stable Fourier Inversion Formula
9.1.1. Setup
Specifically, we use a “stable” Fourier inversion result for a formally defined “stable” Fourier transform. Fix $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ in either the self-dual or conjugate self-dual case and a finite set of finite places $S$. Let $\check{G}_{S}^{\mathrm{temp}}$ be the tempered unitary dual of $G_{S}$.
Let $\check{G}_{S}^{\mathrm{st},\mathrm{temp}}$ be the set of tempered $L$-packets/parameters for $G_{S}$ (or in the case of $\mathrm{SO}_{2n}$, outer-automorphism orbits of these). Since tempered $L$-packets partition tempered representations by either [Art13, Thm 1.5.1(b)] or [Mok15, Thm 2.5.1(b)], there is a map
$\rho:\check{G}_{S}^{\mathrm{temp}}\to\check{G}_{S}^{\mathrm{st},\mathrm{temp}}.$
Associated to $\rho$, we have pushforward and normalized pullback maps between functions on $\check{G}_{S}^{\mathrm{temp}}$ and those on $\check{G}_{S}^{\mathrm{st},\mathrm{temp}}$:
$\rho_{*}f(\Pi)=\sum_{\pi\in\rho^{-1}(\Pi)}f(\pi),\qquad\bar{\rho}^{*}f(\pi)=|\rho^{-1}(\rho(\pi))|^{-1}f\circ\rho(\pi),$
where $|\rho^{-1}(\rho(\pi))|$ is the cardinality of the packet $\Pi$ containing $\pi$. Dual to these we also have pullback and normalized pushforward between functionals on both spaces, denoted $\rho^{*},\bar{\rho}_{*}$. Note that for all functions $f$ and functionals $\mu$,
$\rho_{*}\bar{\rho}^{*}f=f,\qquad\bar{\rho}_{*}\rho^{*}\mu=\mu.$
Conversely, $\bar{\rho}^{*}\rho_{*}f$ and $\rho^{*}\bar{\rho}_{*}\mu$ can be though of as averaging over $L$-packets.
Next, for any $U_{S}\subseteq G_{S}$ such that $U_{S}\cap Z_{G_{S}}$ is compact, define
$\mathcal{PW}(G_{S},U_{S})=\{\widehat{f}:\check{G}^{\mathrm{temp}}_{S}\to\mathbb{C}:f\in\mathcal{C}_{c}^{\infty}(U_{S})\},$
the Paley-Wiener space of Fourier transforms of smooth, compactly supported functions with support in $U_{S}$. We can similarly define an unrestricted $\mathcal{PW}(G_{S})$ of all compactly supported functions. By looking at the image under pushforward, we also define $\mathcal{PW}^{\mathrm{st}}(G_{S},U_{S})$ and $\mathcal{PW}^{\mathrm{st}}(G_{S})$.
To avoid the use of the Ramanujan conjecture333this proven for regular integral infinitesimal character at infinity in many of our cases in [Shi11], [KS20], [KS23] through constructing associated Galois representations. However, these results are not yet comprehensive so we will instead use a lower-tech argument., we need to extend these constructions to the non-tempered specturm in way that is consistent with the endoscopic classification. Since non-tempered $L$-packets are not necessarily made up of irreps, we need to work with a modified Fourier transform for non-tempered $\pi$:
$\widehat{f}(\pi):=\operatorname{tr}_{\mathcal{I}_{\pi}}f$
where $\mathcal{I}_{\pi}$ is the unique parabolic induction of a tempered representation tensor a character that has $\pi$ as a Langlands quotient. Then, using the definition of non-tempered $L$-packets (see e.g. [Art13, pp. 44-45]), there is an extension our map
$\rho:\check{G}_{S}\to\check{G}_{S}^{\mathrm{st}}$
to have image in the space of all $L$-parameters such that for $f\in\mathcal{C}_{c}^{\infty}(G_{S})$
$\rho_{*}\widehat{f}(\Pi)=\operatorname{tr}_{\Pi}f$
for even non-tempered $L$-packets $\Pi$. Of course, our modified $\widehat{f}$ still satisfies the Fourier inversion formula since it is unchanged on the support of $\mu^{\mathrm{pl}}$.
9.1.2. Global Input
Now we are ready for the main proof:
Proposition 9.1.1.
Let $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ in either the self-dual ($G\neq\mathrm{SO}_{2})$ or conjugate self-dual case. Let $S$ be a finite set of finite places of $F$ and $U_{S}\subseteq G_{S}$ such that $U_{S}\cap Z_{G_{s}}$ is compact. Then there exists a sequence of measures $\mu_{n}$ on $\check{G}_{S}^{\mathrm{st}}$ such that for all $\widehat{f}_{S}\in\mathcal{PW}(G_{S},U_{S})$, we have
$\lim_{n\to\infty}\rho^{*}\mu_{n}(\widehat{f}_{S})=\mu_{G_{S}}^{\mathrm{pl}}(\widehat{f}_{S}).$
Proof.
Since $Z_{G_{\infty}}$ is compact for the groups $G$ under consideration, there always exists compact open $K^{S,\infty}\subseteq G^{S,\infty}$ such that $U_{S}\times K^{S,\infty}\cap Z_{G}(F)=1$. Let $\widehat{f}_{S}$ be the Fourier transform of $f_{S}\in\mathcal{C}_{c}^{\infty}(U_{S})$.
Let $\lambda$ be a regular, integral infinitesimal character, and let
$f_{\lambda}=(\dim\lambda)^{-1}\mathrm{EP}_{\lambda}f_{S}\bar{\mathbf{1}}_{K^{S,\infty}}.$
Our strategy is to define
$\mu_{\lambda}(\widehat{f}_{S}):=S_{\Sigma_{\eta}}^{G}(f_{\lambda}),$
show that $\mu_{\lambda}$ is indeed a measure on $\check{G}_{S}$ pulled back from $\check{G}_{S}^{\mathrm{st}}$, and to show convergence to the Plancherel measure as $m(\lambda)\to\infty$.
First, the equalities [DGG22, (5.3.1)] together with the twisted endoscopic character identities for simple $A$-parameters shows that
$\mu_{\lambda}(f_{\lambda})=S^{G}_{\Sigma_{\eta}}(f_{\lambda})\\ =(\dim\lambda)^{-1}\sum_{\Pi\in\Sigma_{\eta,\lambda}}\left(\sum_{\pi^{S,\infty}\in\Pi^{S,\infty}}\dim((\pi^{S,\infty})^{K^{S,\infty}})\right)\left(\sum_{\pi_{S}\in\Pi_{S}}\operatorname{tr}_{\pi_{S}}(f_{S})\right),$
which by inspection is of the form $\rho^{*}\mu_{\lambda}^{\mathrm{st}}$ for some $\mu_{\lambda}^{\mathrm{st}}$ on $\check{G}^{\mathrm{st}}_{S}$.
For the convergence, we have from Theorem 6.4.1 that
$S^{G}_{\Sigma_{\eta}}(f_{\lambda})=\tau^{\prime}(G)f_{S}(1)\bar{\mathbf{1}}_{K^{S,\infty}}(1)+O(m(\lambda)^{-C})$
using that our choice of $K^{S,\infty}$ guarantees that all non-identity terms in the sum (6.2.1) defining $\Lambda$ vanish. Evaluating constants and using Fourier inversion shows that
$\lim_{m(\lambda)\to\infty}S^{G}_{\Sigma_{\eta}}(\tau^{\prime}(G)^{-1}\operatorname{vol}(K^{S,\infty})f_{\lambda})=\mu^{\mathrm{pl}}(\widehat{f}_{S}).$
Combining gives that
(9.1.1) $\lim_{m(\lambda)\to\infty}\rho^{*}\mu^{\mathrm{st}}_{\lambda}(\widehat{f}_{S})=\tau^{\prime}(G)\operatorname{vol}(K^{S,\infty})^{-1}\mu^{\mathrm{pl}}(\widehat{f}_{S}).$
which is what we wanted after choosing an appropriate sequence of $\lambda$. ∎
9.1.3. Final Result
Rephrasing this in familiar language:
Theorem 9.1.2 (Stability of Plancherel Measure).
Let $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ in either the self-dual or conjugate self-dual case and $S$ a finite set of finite places of $F$. Then for all $f_{S}\in\mathcal{C}_{c}^{\infty}(G_{S})$,
[…]
Proof.
The first equality is by Fourier inversion. The second is equivalent to proving that for all $U_{S}$ such that $U_{S}\cap Z_{G_{S}}$ is compact and $\widehat{f}_{S}\in\mathcal{PW}(G_{S},U_{S})$,
$\mu^{\mathrm{pl}}(\widehat{f}_{S})=\rho^{*}\bar{\rho}_{*}\mu^{\mathrm{pl}}(\widehat{f}_{S}).$
Let $\mathcal{PW}^{\mathrm{st}}(G_{S},U_{S}):=\rho_{*}\left(\mathcal{PW}(G_{S},U_{S})\right)$. Proposition 9.1.1 gives that, as a linear functional on $\mathcal{PW}(G_{S},U_{S})$, the Plancherel measure $\mu^{\mathrm{pl}}$ is of the form $\rho^{*}\mu$ for some linear functional $\mu$ on $\mathcal{PW}^{\mathrm{st}}(G_{S},U_{S})$.
The result then follows since $\bar{\rho}_{*}\circ\rho^{*}=\mathrm{id}$ on the linear dual of $\mathcal{PW}^{\mathrm{st}}(G_{S},U_{S})$. ∎
We note a corollary that might be of independent interest (though it follows from the formal degree conjecture whenever that is known—e.g. from [BP21]):
Corollary 9.1.3.
Let $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ in either the self-dual or conjugate self-dual case and $v$ a finite place. For any supercuspidal $L$-packet $\Pi_{v}$ of representations of $G_{v}$ (i.e. whose $L$-parameter is irreducible and trivial on the Deligne-$\mathrm{SL}_{2}$), all $\pi_{v}\in\Pi_{v}$ have the same formal degree.
Proof.
All representations $\pi_{v}$ in a supercuspidal $L$-packet are supercuspidal [^4] This is part of the construction of the LLC through the endoscopic classification.. Therefore, their compactly supported matrix coefficients $c_{\pi_{v}}$ satisfy that $\widehat{c}_{\pi_{v}}=\mathbf{1}_{\{\pi_{v}\}}$ so that $c_{\pi_{v}}(1)$ is the formal degree of $\pi_{v}$. The result then follows from applying Theorem 9.1.2. ∎
9.2. Positivity Test
We now derive from Theorem 9.1.2 a positivity test:
9.2.1. Understanding $G_{S}^{\mathrm{st},\mathrm{temp}}$
We first need to understand $\check{G}_{S}^{\mathrm{st},\mathrm{temp}}$ and $\bar{\rho}_{*}\mu^{\mathrm{pl}}$ better through Harish-Chandra’s Plancherel formula. Without loss of generality, we assume $S=\{v\}$ is a singleton. See [Sil96] for a review of the Plancherel formula which we will loosely follow and the proof of 3.5.1 in [Art13] for a very terse summary focused on applications to the endoscopic classification.
First, $\check{G}^{\mathrm{temp}}_{v}$ decomposes into pieces called $\mathfrak{l}$-components. Consider the set of pairs $(M,\sigma)$ of Levis $M$ of $G_{v}$ and discrete series (modulo the center) $\sigma$ of $M$ up to conjugation by $G_{v}$. If we fix a minimal Levi $M_{0}$ and $M\supset M_{0}$, then the set $M^{*}/\Omega^{G}_{M}$ of unramified unitary characters of $M$ mod $\Omega^{G}_{M}:=N_{G_{v}}(M)$ acts on the set of $(M^{\prime},\sigma)$ with $M^{\prime}$ conjugate to $M$.
Let $\mathcal{O}$ be the set of orbits of this action. Each $[M,\sigma]\in\mathcal{O}$ can be given the structure of quotient of a product of circles by a finite group through the corresponding structure on $M^{*}/\Omega^{G}_{M}$. If $[M,\sigma]\in\mathcal{O}$, define
$\check{G}_{v}[M,\sigma]:=\bigsqcup_{(M,\sigma)\in[M,\sigma]}\{\text{irred. subquotients }\pi\text{ of }\mathcal{I}_{P_{M}}^{G}(\sigma)\},$
where $\mathcal{I}_{P_{M}}^{G}$ is the normalized parabolic induction through a parabolic $P_{M}$ with Levi factor $M$ (its irreducible subquotients do not depend on the choice of $P$).
Then
$\check{G}^{\mathrm{temp}}_{v}=\bigsqcup_{[M,\sigma]\in\mathcal{O}}\check{G}_{v}[M,\sigma].$
In addition, for any open $U\subseteq[M,\sigma]$, the set
$\bigsqcup_{(M,\sigma)\in U}\{\text{irred. subq. }\pi\text{ of }\mathcal{I}_{P_{M}}^{G}(\sigma)\}$
is open in the Fell topology (this can be seen for example by the characterization of the restriction of the Fell topology to the infinite-dimensional irreps in $\check{G}_{v}$ in terms of matrix coefficients and the trace Paley-Weiner theorem).
We now establish an analoguous decomposition for $\check{G}_{v}^{\mathrm{st},\mathrm{temp}}$. If $\varphi$ is a tempered parameter of $G$, there exists a Levi $M$ such that $\varphi$ is the unique-up-to-conjugation pushforward of discrete parameter $\varphi_{M}$ for $M$. As part of the construction of LLC through the endoscopic classification,
(9.2.1) $\Pi_{\varphi}=\bigsqcup_{\sigma_{M}\in\Pi_{\varphi_{M}}}\{\text{irred. subq. }\pi\text{ of }\mathcal{I}_{P_{M}}^{G}(\sigma_{M})\}$
(see the discussion before Proposition 6.6.1 in [Art13] and the discussion in §7.6 in [Mok15]).
Therefore, given a Levi $M$ and discrete packet $\Pi_{M}$ of $M$, we can similarly define a stable $\mathfrak{l}$-component $\check{G}^{\mathrm{st}}_{v}[M,\Pi_{M}]$ which also inherits a real manifold-quotient structure from $M^{*}/\Omega^{G}_{M}$ and gives disjoint union over an analogous $\mathcal{O}^{\mathrm{st}}$:
$\check{G}^{\mathrm{st},\mathrm{temp}}_{v}=\bigsqcup_{[M,\Pi_{M}]\in\mathcal{O}^{\mathrm{st}}}\check{G}^{\mathrm{st}}_{v}[M,\Pi_{M}].$
In this way the map $\rho:\check{G}_{v}^{\mathrm{temp}}\to\check{G}_{v}^{\mathrm{st},\mathrm{temp}}$ is continuous. Note also that since $\mu^{\mathrm{pl}}$ is supported on all of $\check{G}_{v}^{\mathrm{temp}}$, its pushforward $\bar{\rho}_{*}\mu^{\mathrm{pl}}$ is supported on all of $\check{G}_{v}^{\mathrm{st},\mathrm{temp}}$.
Our practical punchline for this discussion is that:
Lemma 9.2.1.
Let $f_{v}$ be a test function on $G_{v}$. Then $\rho_{*}\widehat{f}_{v}$ is continuous on $\check{G}_{v}^{\mathrm{st},\mathrm{temp}}$.
Proof.
It suffices to show this on each stable $\mathfrak{l}$-component $\check{G}^{\mathrm{st}}_{v}[M,\Pi_{M}]$. For each $\pi_{M}\in\Pi_{M}$,
$\chi\mapsto\operatorname{tr}_{\mathcal{I}_{P_{M}}^{G}(\pi_{M}\otimes\chi)}f_{v}$
is continuous on $\check{G}_{v}[M,\pi_{M}]$ by the trace Paley-Wiener theorem [BDK86] (which actually gives much stronger regularity). However, by (9.2.1), the function $\rho_{*}\widehat{f}$ on $\check{G}_{v}^{\mathrm{st}}[M,\Pi_{M}]$ is just the sum of these functions over $\pi_{M}$. ∎
9.2.2. Spectral Interpretation of Main Term
As a second input for the positivity test, we import some formalism from [Dal22, §8.3] to interpret (6.2.1) spectrally. For any group $H$, there is a central character map $\Omega:\check{H}\to\check{Z}_{H}$. Given a test function $f$ on $H$ and a central character $\omega$, we can define the conditional Plancherel expectation as the Radon-Nikodym derivative
$E^{\mathrm{pl}}(\widehat{f}|\omega):=\left.\frac{d\Omega_{*}(\widehat{f}\mu^{\mathrm{pl}}_{H})}{d\mu^{\mathrm{pl}}_{Z_{H}}}\right|_{\omega}$
which exists and is necessarily continuous as a function on $\check{Z}_{H}$.
Now, consider $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ in either the self-dual or conjugate self-dual case. Let $f=\prod_{v\nmid\infty}f_{v}$ be a test function on $G^{\infty}$ such that $f_{v}=\bar{\mathbf{1}}_{K_{v}}$, for $K_{v}$ a hyperspecial maximal compact, outside a finite set of places $S$. Then, as a special case of [Dal22, Prop. 8.3.5], we can write
(9.2.2) $\sum_{\gamma\in Z_{G}(F)}\omega^{-1}(\gamma)f(\gamma)=C\sum_{\omega_{S}\in\Omega_{\lambda}(K^{S,\infty})}E^{\mathrm{pl}}(\widehat{f}_{S}|\omega_{S})$
for some constant $C>0$ depending on various choices of measures and where
$\Omega_{\lambda}(K^{S,\infty})=\{\omega_{S}\in\check{Z}_{G_{S}}:\omega_{S}|_{Z_{G}(F)\cap K^{S,\infty}}=\omega_{\lambda}|_{Z_{G}(F)\cap K^{S,\infty}}\}.$
In many cases, $Z_{G_{S}}$ is compact so $\check{Z}_{G_{S}}$ is discrete. Then, $E^{\mathrm{pl}}(\widehat{f}_{S}|\omega_{S})$ simplifies as just the integral of $\widehat{f}\,d\mu^{\mathrm{pl}}_{G_{S}}$ over some union of connected components of $\check{G}_{S}$.
9.2.3. The Positivity Test
Putting everything together lets us show:
Proposition 9.2.2.
Let $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ in either the self-dual or conjugate self-dual case. Let $f$ be a test function on $G(\mathbb{A})$ equal to $\mathbf{1}_{K_{v}}$ outside of a finite set of places $S$. Assume that for all tempered $L$-packets (or in the case of $\mathrm{SO}_{2n}$ the appropriate unions of such) $\Pi_{S}$ of $G_{S}$,
$\operatorname{tr}_{\Pi_{S}}(f)\geq 0.$
Then for any character $\omega$ of $Z_{G,\infty}$
$\sum_{\gamma\in Z_{G}(F)}\omega^{-1}(\gamma)f(\gamma)\geq 0.$
Furthermore, the inequality is strict if there exists a tempered $L$-packet $\Pi_{S}$ such that $\operatorname{tr}_{\Pi_{S}}(f)>0$ and $\omega_{\pi_{S}}|_{Z_{G}(F)\cap K^{S,\infty}}=\omega|_{Z_{G}(F)\cap K^{S,\infty}}$ for all $\pi_{S}\in\Pi_{S}$.
Proof.
Consider any measurable $X\subseteq\check{Z}_{G_{S}}$ and $G_{S,X}:=\Omega^{-1}(X)$. Then $\check{G}_{S,X}$ is the exact preimage of a subset $\check{G}_{S,X}^{\mathrm{st}}$ of $\check{G}_{S}^{\mathrm{st}}$ since central characters are constant on $L$-packets by Lemma 5.2.2. Then by the definition of $E^{\mathrm{pl}}$ and Theorem 9.1.2,
[…]
This forces $E^{\mathrm{pl}}(\widehat{f}_{S}|\omega)\geq 0$ for all $\omega$ which, by (9.2.2), proves the non-negativity statement.
For the strict positivity result, if such a packet $\Pi_{S}$ with appropriate central character $\omega_{0}$ exists, then from Lemma 9.2.1 and the fact that every tempered packet is in the support of $\bar{\rho}_{*}\mu^{\mathrm{pl}}_{G_{S}}$, we get that $\rho_{*}(\widehat{f}_{S})\,d\bar{\rho}_{*}\mu^{\mathrm{pl}}_{G_{S}}$ is supported at $\Pi_{S}$. By an argument similar to the previous paragraph, this implies that $E^{\mathrm{pl}}(\widehat{f}_{S}|\omega)\,d\mu^{\mathrm{pl}}_{Z_{G_{S}}}$ is supported at $\omega=\omega_{0}$. In particular, one summand $E^{\mathrm{pl}}(\widehat{f}_{S}|\omega_{0})$ appearing in (9.2.2) is strictly positive. ∎
9.3. Existence of Representations
Our strict positivity test requires the existence of local (conjugate) self-dual representations of $GL_{N}$ with arbitrary conductor and specified central character. We use the shorthand notation for conductors, see §3.1.
9.3.1. Self-dual case.
First note that the result fails for $N=1$. Self-dual characters of $F_{v}^{\times}$ must be valued in ±$\pm 1$, and so have conductor at most $1$ (though there always does exist a self-dual character of conductor 1.)
Lemma 9.3.1.
Fix a place $v$, an integer $k\geq 0$, and a character $\omega:F_{v}^{\times}\to\pm 1$ with conductor $\leq k$. Then there exists a self-dual tempered representation $\pi_{v}$ of $GL_{2}(F_{v})$ of conductor $c(\pi_{v})=k$ and central character $\eta$. Moreover,
  • If $\eta$ is trivial then $\pi_{v}$ can be chosen to be symplectic for all $k$, or orthogonal for $k$ even.
  • If $\eta$ is nontrivial, then $\pi_{v}$ is necessarily of orthogonal type. If it is cuspidal, then its parameter factors through ${{}^{L}}SO_{2}^{\eta}$.
Proof.
Self-dual representations of $GL_{2}$ are characterized by the property that their central character is valued in ±$\pm 1$.
For $k$ even and $\eta=1$, we can take $\pi_{v}$ to be the principal series associated to characters $\chi\boxtimes\chi^{-1}$ with $\chi$ unitary of conductor $k/2$. This is self-dual with trivial central character and conductor $k$. The images of the non-irreducible parameters associated to these representations preserve both a symmetric and skew-symmetric form and can arise as the localizations of automorphic representations of either orthogonal or symplectic-type.
For $k\leq 2$ and orthogonal type, we can consider the principal series induced from a pair of self-dual quadratic characters $\chi_{1}$, $\chi_{2}$ such that $\chi_{1}\chi_{2}=\eta$ and $c(\chi_{1})+c(\chi_{2})=k$.
For $k=1$ and symplectic-type, consider the principal series induced from a pair of self-dual quadratic characters $\chi_{1}$, $\chi_{2}$ such that $\chi_{1}\chi_{2}=\eta$ and $c(\chi_{1})+c(\chi_{2})=1$ (in the orthogonal case).
Finally, for the remaining cases with $k\geq 2$ we rely on the construction of supercuspidals in [BH06], from characters $\chi$ of $L_{w}^{\times}$, for a quadratic extension $L_{w}/F_{v}$ with ramification index $e$. In this construction, if $\chi$ has conductor $\mathfrak{p}_{w}^{\ell}$, the resulting representation $\pi_{\chi}$ has central character $\chi\mid_{F^{\times}_{v}}$ and conductor $\mathfrak{p}_{v}^{k}$ for $k=\frac{2\ell}{e}+2$, see [BH06, Theorems 20.2 and 25.2]. This reduces the existence question to that of characters of suitably ramified $L^{\times}_{w}$ with a given restriction to $F_{v}^{\times}$ and conductor, which follows from standard results on characters of finite abelian groups.
Supercuspidal representations correspond to irreducible $L$-parameters, which factor through a non-abelian subgroup of ${{}^{L}}G_{N}$. This subgroup is ${}^{L}\!{\mathrm{Sp}}_{2}={}^{L}\!{S}L_{2}$ precisely when the central character $\eta=\chi\mid_{F^{\times}}$ is trivial. Otherwise, the $L$-parameter is orthogonal and factors though ${}^{L}\!{S}O_{2}^{\eta}$. ∎
Lemma 9.3.2.
Let $N>2$ and let $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$. Fix a place $v$ and a conductor $k$, and, if $\widehat{G}$ is orthogonal, a central character $\omega_{v}=\pm 1$, that can only be $-1$ if $k>0$. Assume further that if $G_{v}=\mathrm{SO}_{N}^{\eta_{v}}$, then $k\geq c(\eta_{v})$.
Then there is a self-dual tempered representation $\pi_{v}$ of $G_{N,v}$ with:
  • conductor $k$,
  • $L$-parameter factoring through ${}^{L}\!{G}_{v}$,
  • central character $\omega_{v}$ on $G_{v}$.
Proof.
Let $m\in\{1,2\}$ have the same parity as $N$ and consider the representation
$\pi_{v}=\chi_{1}\boxplus\cdots\boxplus\chi_{\lfloor\frac{N-1}{2}\rfloor}\boxplus\sigma_{m}\boxplus\chi^{-1}_{\lfloor\frac{N-1}{2}\rfloor}\boxplus\cdots\boxplus\chi^{-1}_{1},$
for (not necessarily self-dual) characters $\chi_{i}$, and $\sigma_{m}$ a self-dual unitary representation of $GL_{m}$. Then $\pi_{v}$ is unitary with the same central character as $\sigma_{m}$ and
$c(\pi_{v})=c(\sigma_{m})+2\sum_{i}c(\chi_{i}),\qquad\epsilon(1/2,\pi_{v})=\epsilon(1/2,\sigma_{m})\prod_{i}\chi_{i}(-1).$
In addition, by Lemma 9.3.1, we can choose $\sigma_{m}$ and the $\chi_{i}$ so that $\pi_{v}$ is tempered, has arbitrary conductor, and arbitrary self-dual central character $\eta_{v}$.
If $N$ is odd, this $\pi_{v}$ is necessarily orthogonal type and factors through ${}^{L}\!{G}={}^{L}\!{S}p_{N-1}^{\eta}$ if the central character of $\pi_{v}$ on $G_{N,v}$ is $\eta_{v}$. In addition, if $c>0$ and $N>2$, we can also choose the $\chi_{i}(-1)$ and therefore the root number. Since the root number determines the central character $\omega_{v}$ on $G_{v}$ by Theorem 5.2.1, $\pi_{v}$ can be chosen to satisfy all our desired conditions.
If $N$ is even and $G=\mathrm{SO}_{N+1}$, we can choose $\sigma_{m}$ to symplectic as in Lemma 9.3.1, implying that $\pi_{v}$ is too (as in the discussion on [Art13, p. 34]). Following Lemma 9.3.1, the representation $\sigma_{m}$ can be chosen so that $\pi_{v}$ satisfies all our desired conditions.
Finally, for $G=\mathrm{SO}_{N}^{\eta}$, we choose $\sigma_{m}$ to have parameter factoring through ${}^{L}\!{\mathrm{SO}}_{2}^{\eta}$ as in Lemma 9.3.1. Then, the parameter of $\pi_{v}$ factors through ${}^{L}\!{\mathrm{SO}}_{N}^{\eta}$. We can also choose the $\chi_{i}(-1)$ appropriately to determine the root number of $\pi_{v}$ and therefore its central character. ∎
9.3.2. Conjugate self-dual case
In the following lemmas, we fix a non-split place $v$ of $F$ and a place $w$ of $E$ above $v$. We use local indexing conventions for conductors (where ramified conductors are integers instead of half-integers, see 3.1.2). Let
(9.3.1) $j_{v}:=\min\{1/2,\{j:1+\mathfrak{p}_{w}^{j}\subseteq N_{E/F}(E^{\times})\}\}.$
The $1/2$ is for indexing in order that the the term $12j_{v}-$ 1, which will appear a lot in this section, satisfies
$2j_{v}-1=\begin{cases}0&v\text{ unramified}\\ 1&v\text{ tame}\\ \geq 1&v\text{ wild.}\end{cases}$
Lemma 9.3.3.
Let $\chi:E_{w}\to\mathbb{C}$ be a conjugate self-dual character, so that restriction $\chi\mid_{F_{v}^{\times}}=\kappa\in\pm 1$ is either trivial or the quadratic character corresponding to the extension $E_{w}$. Then the $L$-parameter of $\chi$ factors through the endoscopic datum $U^{\kappa}_{1}$.
Proof.
We write down the explicit embeddings, following [Mok15, §2.1]. Using class field theory, we view the character $\chi$ as a parameter
$\varphi_{\chi}:\Gamma_{E_{w}}\to GL_{1}\times\Gamma_{E_{w}}.$
Recall that
${}^{L}\!{G}(1)=(GL_{1}\times GL_{1})\rtimes\Gamma_{F_{v}}$
where the action of $\Gamma_{F_{v}}$ is through the quotient $\Gamma_{E_{w}/F_{v}}$ and the nontrivial element swaps the two factors. Self-duality enables us to extend $\varphi_{\chi}$ to a parameter
$\varphi_{\chi}:\Gamma_{F_{v}}\to{}^{L}\!{G}(1)$
as follows. Fixing a representative $z_{c}$ of $\Gamma_{F_{v}}\setminus\Gamma_{E_{w}}$, we let:
$\varphi_{\chi}(\sigma)$ $=(\chi(\sigma),\chi^{-1}(\sigma))\rtimes\sigma\quad\sigma\in\Gamma_{E_{w}}$
$\varphi_{\chi}(z_{c})$ $=(\alpha\chi(z_{c}^{2}),\alpha^{-1})\rtimes z_{c}.$
In this way the value of $\varphi_{\chi}(z_{c})$ is determined up to choice of a scalar $\alpha$ (or, equivalently, up to a choice of representative $z_{c}$). Note that since $z_{c}^{2}$ is invariant under conjugation by $z_{c}$, we must have $\chi(z_{c}^{2})=\chi(z_{c}^{2})^{-1}$, i.e. $\chi(z_{c}^{2})\in\{\pm 1\}$. Additionally, the condition that $z_{c}\notin\Gamma_{E_{w}}$ implies by local class field theory that $\chi(z_{c}^{2})=\kappa$.
Next, recall that
${}^{L}\!{U}(1)=GL_{1}\rtimes\Gamma_{F_{v}}$
where again, the action is via the quotient $\Gamma_{E_{w}/F_{v}}$ and conjugation by a nontrivial element sends $g\mapsto g^{-1}.$ Each embedding of ${}^{L}\!{U}(1)$ into ${}^{L}\!{G}(1)$ involves a choice: for $\kappa\in{\pm 1}$ we fix an auxiliary character $\eta$ of $E_{w}$ whose restriction of $F_{v}^{\times}$ is $\kappa$ and we have
$\xi_{\kappa}:{}^{L}\!{U}(1)$ $\to{}^{L}\!{G}(1)$
$\xi_{\kappa}(g)$ $=(g,g^{-1})\rtimes 1,\quad g\in GL_{1}(\mathbb{C})$
$\xi_{\kappa}(\sigma)$ $=(\eta(\sigma),\eta(\sigma^{-1}))\rtimes\sigma,\quad\sigma\in\Gamma_{E_{w}}$
$\xi_{\kappa}(z_{c})$ $=(\kappa,1)\rtimes z_{c}.$
For the parameter $\varphi_{\chi}$ to factor through $\xi_{\kappa}$, the choice of image of $z_{c}$ must be compatible with the embedding, i.e. we must have $\kappa=\chi(z_{c}^{2})$, which gives the result. ∎
Lemma 9.3.4.
Fix a conductor $k\in\mathbb{Z}_{\geq 0}$. Then, for each $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(1)$, there exists a conjugate self-dual character of $E_{w}^{\times}$ of conductor $k$ such that the associated $L$-parameter factors through ${}^{L}\!{G}$ if and only if one the following hold:
  • when $\kappa=1$:
    • $E_{w}$ is unramified,
    • $E_{w}$ is ramified and $k$ is even,
  • when $\kappa=-1$:
    • $E_{w}$ is unramified
    • $E_{w}$ is ramified and either:
        $k=2j_{v}-1$,
        $k\geq 2j_{v}$ is even,
where $j_{v}$ is as in (9.3.1).
Proof.
The condition of self-duality for characters is $\chi(\overline{g})=\chi(g^{-1})$, or equivalently, that $\chi|_{F_{v}^{\times}}\in\pm 1$, where we abuse notation and denote by $-1$ the unique nontrival character of $F^{\times}/N(E^{\times}_{w})$ where $N=N_{E_{w}/F_{v}}$ is the norm map. The sign ±$\pm 1$ determines which element of $\mathcal{E}_{\mathrm{sim}}(1)$ the parameter $\chi$ factors through, as in Lemma 9.3.3.
Next, denote $U^{0}=\mathcal{O}_{E_{w}}^{\times}$ and $U^{k}=1+\mathfrak{p}_{w}^{k}$. For indexing convenience, denote $U^{-1}$ to be some set strictly bigger than $E^{\times}_{w}$. We now consider necessary and sufficient conditions for the existence of a character of sign $\kappa$ and conductor $k$:
$\kappa=1$
There exists a character of sign $\kappa=1$ and conductor $k$ if and only if
(9.3.2) $U^{k-1}\not\subset F^{\times}U^{k}.$
For $k=0$, this condition is empty. For $k>0$, we have
[…]
and, by comparing cardinalities, the latter holds if any only if
  • $E_{w}$ is unramified, or
  • $E_{w}$ is ramified and $k$ is even.
$\kappa=-1$:
To determine a necessary and sufficient condition for existence here, we use that given two characters $\chi_{1}$, $\chi_{2}$ with conductors $c(\chi_{1})>c(\chi_{2})$, then $c(\chi_{1}\chi_{2})=c(\chi_{1})$. Furthermore, the parities satisfy $\kappa(\chi_{1}\chi_{2})=\kappa(\chi_{1})\kappa(\chi_{2})$. In particular, if $k_{0}$ is the smallest possible conductor for which there exists $\chi$ with $\kappa=-1$, then for all $k>k_{0}$, conductor $k$ is possible with $\kappa=-1$ if and only if it is with $\kappa=1$. Therefore, it suffices to find this $k_{0}$.
If $E_{w}$ is unramified, then any character $\chi$ with $\chi(\mathcal{O}_{E}^{\times})\equiv 1$ and such that $\chi$ takes some uniformizer to $-1$ is unramified and has $\kappa=-1$. Therefore, $k_{0}=0$.
If $E_{w}$ is ramified, note the existence of a character $\chi$ with $\kappa=-1$ requires that
$F^{\times}\not\subset N(E^{\times})U^{k}\iff F^{\times}\cap U^{k}\subseteq N(E^{\times}).$
This forces $k\geq 2j-1$. To construct $\chi$ with $k=2j-1$, it suffices to further note that
$1+\mathfrak{p}_{v}^{j-1}\notin N(E^{\times})\implies U^{2j-2}\notin N(E^{\times})U^{2j-1}$
Therefore, $k_{0}=2j-1$. ∎
The full list of possible pairs $(k,l)$ of conductor $c(\pi_{v})=k$ and central character conductor $c(\omega_{\pi_{v}})=l$ is very hard to determine and likely does not have a nice description, particularly in the wildly ramified case. We settle for providing an easy-to-describe subset that covers most relevant possibilities outside of some wild ramification issues.
Lemma 9.3.5.
Let $N\geq 4$ be even, and let $G=U_{N}^{+}\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ in the conjugate self-dual case. Fix a non-split place $v$ and let $e_{v}:=e(E_{w}:F_{v})$ be the ramification index of $E_{w}/F_{v}$. Fix $k\geq 0$ a conductor, and fix a central character $\omega_{v}$ on $G_{v}$ transferring to $\omega^{\prime}_{v}$ on $G_{N,v}$ with conductor $l\leq k$.
Then there is a conjugate self-dual tempered representation $\pi_{v}$ of $G_{N,v}$ with conductor $c(\pi_{v})=k$ such that its corresponding $L$-parameter $\psi_{\pi_{v}}$ factors through ${}^{L}\!{G}_{v}$ and has central character $\omega_{v}$ on $G_{v}$ for the following cases:
  • $l=\max\{k-(2j_{v}-1),0\}$ when $e_{v}=1$ and the largest even number less than this otherwise and:
    • $v$ unramified: all possible $k$,
    • $v$ tamely ramified: all possible $k$,
    • $v$ wildly ramified: $k\leq N/2$ or $k>4j_{v}-2$.
  • $l=0$:
    • $v$ unramified: all possible $k$,
    • $v$ tamely ramified: all possible $k$,
    • $v$ wildly ramified: $k\leq N/2$ or $k\geq 8j_{v}-4$
Proof.
First, note that $k=l=0$ is always achievable by looking at a self-dual Satake parameter.
Otherwise, consider $\pi_{v}$ to be the Langlands quotient of the parabolic induction of blocks
$\rho_{1}\boxtimes\cdots\boxtimes\rho_{t}$
where each $\rho_{i}$ is a conjugate-symplectic representation of some even $G_{N_{i},v}$. It suffices to consider the case where all $\rho_{i}$ are trivial except for one that is a principal series with central character $\omega^{\prime}_{v}$ and possibly some number that are dimension-2 Steinbergs.
Since $c(\omega_{\pi_{v}})=\sum_{i}c(\omega_{\rho_{i}})$ and $c(\pi_{v})=\sum_{i}c(\rho_{i})$, it suffices to describe the possible values of conductor $c(\rho_{i})=k$ and central character conductor $c(\omega_{\rho_{i}})=l$ that can be achieved with our list of possible $\rho_{i}$ and choose appropriately amongst these options.
Rank $2$ Principal Series:
Let $\chi_{0}$ be a choice of conjugate-symplectic character of $E_{v}^{\times}$ with minimum conductor as in the proof of Lemma 9.3.4. Choose conjugate orthogonal $\nu_{1}\nu_{2}=\omega^{\prime}_{v}$ with conductors $k_{1},k_{2}$ respectively. Consider $\rho_{i}$ the parabolic induction of $\chi_{0}\nu_{1}\boxtimes\chi_{0}^{-1}\nu_{2}$ which is conjugate-symplectic since the components are (see the discussion after Definition 2.4.7 in [Mok15]).
There are two possible cases for conductors:
  • $k_{1}=k_{2}\geq l$: Then $k=\max\{2k_{1},4j_{v}-2\}$.
  • $k_{1}=l$ and $k_{1}>k_{2}$: Then $k=\max\{l,2j-1\}+\max\{k_{2},2j_{v}-1\}$.
Note in the above computations that in the ramified cases when $j_{v}\neq 0$, the $k_{i}$ are always even and therefore not equal to $12j_{v}-$ 1. Therefore, we never need to worry about the case where the conductor of a product with $\chi_{0}$ is not the maximum of its factors.
This in particular realizes the following possible pairs $(k,l)$ (with $\omega^{\prime}_{v}$ arbitrary of conductor $l$):
  • $v$ unramified: $k$ even and $l=0$ or $k$ arbitrary and $k=l$
  • $v$ tamely ramified: $k\geq 4$ a multiple of $4$ and $l=0$, $k=2$ and $l=0$, or $k>2$ odd and $l=k-1$.
  • $v$ wildly ramified: $k\geq 4j_{v}$ a multiple of $4$ and $l=0$, $k=4j_{v}-2$ and $l=0$, or $k>4j_{v}-2$ odd and $l=k-(2j_{v}-1)$.
Rank-$4$ Principal Series:
In the ramified cases when $l=0$, we also need to consider dimension-$4$ principal series. Choose $\nu_{1}\nu_{2}\nu_{3}=1$ and $\rho_{i}$ the parabolic induction of
$\chi_{0}\nu_{1}\boxtimes\chi_{0}\nu_{2}\boxtimes\chi_{0}^{-1}\nu_{3}\boxtimes\chi_{0}^{-1}.$
We consider the case when $c(\nu_{0})=c(\nu_{1})=k_{2}\geq c(\nu_{2})=k_{2}$. Then $\rho_{i}$ has conductor
$k=\max\{2k_{1},4j_{v}-2\}+\max\{k_{2},2j_{v}-1\}+(2j_{v}-1)$
which can be any integer larger than $48j_{v}-$ 4.
Steinberg
If $\rho_{i}$ is the Steinberg, then it gives $k=1$ and $l=0$.
Even-Rank Trivial
If $\rho_{i}$ is trivial with even rank, then it is both conjugate-orthogonal and conjugate-symplectic. This gives $k=0$ and $l=0$. ∎
Note.
Constructing representations of odd conductor when $N$ is even crucially depended on choosing $G=U^{+}_{N}$ through the possible choice of characters being induced. This is a good consistency check for Proposition 8.3.1: when $G=U^{-}_{N}$, the function $\widetilde{E}^{\infty}_{v,k}=\widetilde{C}^{\infty}_{v,k}$ transfers to 00 when $k$ is odd. Given Proposition 9.2.2, this is incompatible with the existence of odd-conductor representations and sign $\kappa=-1$.
At split places things are simpler:
Lemma 9.3.6.
Let $N\geq 3$, and let $G=U_{N}^{\pm}\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ in the conjugate self-dual case. Fix a split place $v$ and a conductor $k$. Fix a central character $\omega_{v}$ on $G_{v}$ transferring to $\omega^{\prime}_{v}$ on $G_{N,v}$ with conductor $l\leq k$.
Then there is a conjugate self-dual tempered representation $\pi_{v}$ of $G_{N,v}$ with conductor $c(\pi_{v})=k$ such that its corresponding $L$-parameter $\psi_{\pi_{v}}$ factors through ${}^{L}\!{G}_{v}$ and has central character $\omega_{v}$ on $G_{v}$.
Proof.
Any such $\pi_{v}$ on $G_{N,v}\simeq\mathrm{GL}_{N}(F_{v})\times\mathrm{GL}_{N}(F_{v})$ is of the form $\pi_{0}\boxtimes(\pi_{0})^{\vee}$. In the split place, the $\kappa=\pm$ embeddings are the same, so this has parameter factoring through both. In our conventions, the conductor of $\pi_{v}$ is that of $\pi_{0}$, so it suffices to find $\pi_{0}$ of conductor $k$ and central character $\omega^{\prime}_{v}=\omega_{v}\circ\phi$. This can be done using principal series. ∎
9.4. Application to $(\widetilde{C}^{\infty}_{\mathfrak{n}})^{G}$
Now we are finally ready to show when our main term (6.2.1) for $f^{\infty}=(\widetilde{C}^{\infty}_{\mathfrak{n}})^{G}$ is positive.
9.4.1. Self-Dual Case
Since it doesn’t take that much extra work in the self-dual case, we prove a statement for all $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ instead of just $G=\mathrm{SO}_{N+1}$ when $N$ is even (though note that for this eventual case of interest, $Z_{G}$ is trivial so the statement can be made simpler):
Theorem 9.4.1.
Let $\widetilde{G}_{N}$ in the self-dual case and $\widetilde{C}^{\infty}_{\mathfrak{n}}$ the test function on $\widetilde{G}_{N}^{\infty}$ defined in 4.4.5. Choose $G\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$ and let $\omega^{\prime}_{v}$ be the corresponding central characters on $Z_{G_{N},v}$ common to all transfers from $G_{v}$ (as in the discussion after Lemma 5.2.3).
Fix a central character $\omega_{\infty}$ on $Z_{G_{\infty}}\subseteq\{\pm 1\}$ and assume:
  • For all $v$, $v(\mathfrak{n})\geq c(\omega^{\prime}_{v})$,
  • If $\omega_{\infty}\neq 1$, $\mathfrak{n}\neq 1$.
Then,
$\sum_{\gamma\in Z_{G}(F)}\omega_{\infty}^{-1}(\gamma)(\widetilde{C}^{\infty}_{\mathfrak{n}})^{G}(\gamma)>0.$
Proof.
We apply Proposition 9.2.2 to the set $S$ containing all primes on which $\mathfrak{n}$ is supported and all finite places at which $\omega^{\prime}_{v}$ is nontrivial. The input for non-negativity follows from the defining property Corollary 4.4.4 of $\widetilde{C}^{\infty}_{\mathfrak{n}}$ plugged into the endoscopic character identities [Art13, Thm 2.2.1(a)] or [Mok15, Thm 3.2.1(a)].
For strict positivity, if $\omega_{\infty}\neq 1$, choose $v$ such that $v(\mathfrak{n})>0$ and set $\omega_{v}$ to be the non-trivial character on $Z_{G_{v}}$. Set all other $\omega_{v^{\prime}}=1$. Then, Lemma 9.3.2 constructs representations $\pi_{v}$ on each $G_{N,v}$ of conductor $\mathfrak{p}_{v}^{v(\mathfrak{n})}$, central character $\omega^{\prime}_{v}$, and that corresponds to a packet on $G_{v}$ with central character $\omega_{v}$.
Then, if $\pi_{S}=\prod_{v\in S}\pi_{v}$ as a representation of $G_{N,S}$, it has conductor $\mathfrak{n}$ and its corresponding $A$-packet $\Pi_{S}$ on $G_{S}$ is made of $\pi_{S}$ with central character satisfying $\omega_{\pi_{S}}|_{Z_{G}(F)}=\omega_{\pi_{\infty}}|_{Z_{G}(F)}$. Therefore, $\Pi_{S}$ is the necessary input to Proposition 9.2.2 for strict positivity (again by the endoscopic character identities). ∎
9.4.2. Conjugate Self-dual case
In the conjugate self-dual case, globalizing central characters is more complicated. For $v$ a place of $F$, let $e_{v}$ be its ramification degree in $E$. For $\mathfrak{n}$ a conductor, let $\mathcal{O}_{E}(\mathfrak{n})\subseteq\mathcal{O}_{E}^{\times}$ be the units that are $1\pmod{\mathfrak{n}}$.
We return to global indexing for conductors (where ramified exponents are half-integers).
Lemma 9.4.2.
Let $G=U^{+}_{N}\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$. Then for each finite place $v$ of $F$ there is $b_{v}\in\mathbb{Z}_{\geq-2}$ depending only on $E/F$ such that for all
  • $\omega_{\infty}$ a character on $Z_{G_{\infty}}$ that transfers to $\omega^{\prime}_{\infty}$ on $Z_{G_{N},\infty}$,
  • conductor $\mathfrak{n}$ such that
    • (i)
      $\mathfrak{n}$ is an ideal of $\mathcal{O}_{F}$ (i.e. $v(\mathfrak{n})$ is an integer for all ramified $v$),
    • (ii)
      $v(\mathfrak{n})\geq b_{v}+2/e_{v}$ for one place or $v(\mathfrak{n})\geq b_{v}+1/e_{v}$ for at least two places,
there exists a global $\omega^{\prime}$ on $Z_{G_{N}}(\mathbb{A}_{F})$ with infinite component $\omega^{\prime}_{\infty}$ and such that ${\omega^{\prime}}^{\infty}$ has conductor exactly $\mathfrak{n}$.
Proof.
Let $\phi(x)=\frac{x}{\bar{x}}$ be the map first defined in §8. Let $\Xi$ be the roots of unity in $E$ and for each finite place $v$ of $F$, consider the smallest $x\in\mathbb{Z}_{\geq 0}$ such that $\Xi\cap\phi(1+\mathfrak{p}_{v}^{x})=1$. Set $b_{v}=x$ if $x>0$ and $b_{v}=-2$ if $x=0$. Note that we are not allowing half-integral $b_{v}$ when $v$ is ramified.
The $b_{v}$ are chosen exactly so that if $\mathfrak{n}$ is a conductor satisfying conditions (i)-(ii) and $\mathfrak{n}^{\prime}$ is a proper divisor of $\mathfrak{n}$, then there is intermediate proper divisor $\mathfrak{n}^{\prime\prime}$ such that $\mathfrak{n}^{\prime}|\mathfrak{n}^{\prime\prime}$ and $\Xi\cap\phi(\mathcal{O}_{E}(\mathfrak{n}^{\prime\prime}))=1$. The integrality condition guarantees that there is at least one conjugate-orthogonal character of conductor $\mathfrak{n}$ by Lemma 9.3.4, so in addition $\phi(\mathcal{O}_{E}(\mathfrak{n}^{\prime\prime}))\neq\phi(\mathcal{O}_{E}(\mathfrak{n}))$.
Next, $Z_{G}(F)\cap Z_{G}(\widehat{\mathcal{O}}_{F})=\Xi$ since it consists of integral elements of $E^{\times}$ that have complex norm one in all their embeddings. It follows that for any compact open $K^{\infty}\subseteq Z_{G}(\widehat{\mathcal{O}}_{F})$,
$Z_{G}(F)\backslash Z_{G}(\mathbb{A}_{F})/K^{\infty}=Z_{G}(F^{\infty})/(\Xi\cap K^{\infty})\times Z_{G}(\mathbb{A}^{\infty}_{F})/Z_{G}(F)K^{\infty}.$
If $\mathfrak{n}^{\prime\prime}$ is a proper divisor of $\mathfrak{n}$ such that $\Xi\cap\phi(\mathcal{O}_{E}(\mathfrak{n}^{\prime\prime}))=1$, and $\phi(\mathcal{O}_{E}(\mathfrak{n}^{\prime\prime}))\neq\phi(\mathcal{O}_{E}(\mathfrak{n}))$, note that:
$Z_{G}(F)\backslash Z_{G}(\mathbb{A}_{F})/\phi(\mathcal{O}_{E}(\mathfrak{n}))\neq Z_{G}(F)\backslash Z_{G}(\mathbb{A}_{F})/\phi(\mathcal{O}_{E}(\mathfrak{n}^{\prime\prime}))\implies\\ Z_{G}(F_{\infty})\times Z_{G}(\mathbb{A}^{\infty}_{F})/Z_{G}(F)\phi(\mathcal{O}_{E}(\mathfrak{n}))\neq Z_{G}(F_{\infty})\times Z_{G}(\mathbb{A}^{\infty}_{F})/Z_{G}(F)\phi(\mathcal{O}_{E}(\mathfrak{n}^{\prime\prime}))\\ \implies Z_{G}(\mathbb{A}^{\infty}_{F})/Z_{G}(F)\phi(\mathcal{O}_{E}(\mathfrak{n}))\neq Z_{G}(\mathbb{A}^{\infty}_{F})/Z_{G}(F)\phi(\mathcal{O}_{E}(\mathfrak{n}^{\prime\prime})).$
Therefore, the properties (i)-(ii) of $\mathfrak{n}$ guarantees that we can globalize $\omega_{\infty}$ to a character $\omega$ of $Z_{G}(F)\backslash Z_{G}(\mathbb{A}_{F})/\phi(\mathcal{O}_{E}(\mathfrak{n}))$ that doesn’t descend to $Z_{G}(F)\backslash Z_{G}(\mathbb{A}_{F})/\phi(\mathcal{O}_{E}(\mathfrak{n}^{\prime}))$ for any proper divisor $\mathfrak{n}^{\prime}|\mathfrak{n}$.
Setting $\omega^{\prime}=\omega\circ\phi$ then suffices by Lemma 5.2.4. ∎
Note.
In many cases, $\Xi\cap(1+\mathfrak{D}_{E/F})=0$ already, so the condition on $\mathfrak{n}$ in Lemma 9.4.2 is trivial (i.e. all $b_{v}$ can be chosen to be $-2$).
Theorem 9.4.3.
Let $N\geq 4$ be even, $G=U^{+}_{N}\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$, and $\widetilde{C}^{\infty}_{\mathfrak{n}}$ the test function on $\widetilde{G}^{\infty}_{N}$ defined in 4.4.5. Let $\omega_{\infty}$ be a central character of $G_{\infty}$. Choose $b_{v}$ as in Lemma 9.4.2 and let $\mathfrak{n}$ be a conductor such that
  • For all wildly ramified $v$, $v(\mathfrak{n})\leq N/4$ or $v(\mathfrak{n})>2j_{v}-1$,
  • There is at least one $v$ such that $\max\{\lfloor v(\mathfrak{n})-(j_{v}-1/2)\rfloor,0\}\geq b_{v}+2/e_{v}$ or there are at least two $v$ such that $\max\{\lfloor v(\mathfrak{n})-(j_{v}-1/2)\rfloor,0\}\geq b_{v}+1/e_{v}$.
Then,
$\sum_{\gamma\in Z_{G}(F)}\omega_{\infty}^{-1}(\gamma)(\widetilde{C}^{\infty}_{\mathfrak{n}})^{G}(\gamma)>0.$
Proof.
Define the modified conductor
$\mathfrak{n}_{c}=\prod_{v}\mathfrak{p}_{v}^{\max\{\lfloor v(\mathfrak{n})-(j_{v}-1/2)\rfloor,0\}}.$
By Lemma 9.4.2, we can find a character $\omega^{\prime}$ on $Z_{G_{N}}(\mathbb{A}_{F})$ such that $\omega^{\prime}_{\infty}$ transfers to $\omega_{\infty}^{-1}$ and $\omega^{\prime}$ has conductor $\mathfrak{n}_{c}$. Then, our assumptions on $\mathfrak{n}$ let us proceed as in Theorem 9.4.1: multiplying together the outputs of Lemmas 9.3.5 and 9.3.6 (using the $l\neq 0$ option for inputs in Lemma 9.3.5) gives a representation $\pi^{\infty}$ of $G_{N}(\mathbb{A}^{\infty})$ with central character $(\omega^{\prime})^{\infty}$ and conductor $\mathfrak{n}$.
The argument then follows exactly as Theorem 9.4.1. ∎
We also prove a version for trivial $\omega_{\infty}$:
Theorem 9.4.4.
Let $N\geq 4$ be even, $G=U^{+}_{N}\in\widetilde{\mathcal{E}}_{\mathrm{sim}}(N)$, and $\widetilde{C}^{\infty}_{\mathfrak{n}}$ the test function on $\widetilde{G}^{\infty}_{N}$ defined in 4.4.5. Choose the conductor $\mathfrak{n}$ such that:
  • For all wildly ramified $v$, $v(\mathfrak{n})\leq N/4$ or $v(\mathfrak{n})>4j_{v}-2$.
Then,
$\sum_{\gamma\in Z_{G}(F)}(\widetilde{C}^{\infty}_{\mathfrak{n}})^{G}(\gamma)>0.$
Proof.
This is the same argument as Theorem 9.4.1 but instead of globalizing $\omega_{\infty}$ using Lemma 9.4.2, we globalize trivial $\omega_{\infty}$ to the trivial character and use the $l=0$ option for inputs in Lemma 9.3.5. ∎
Note.
A more careful analysis of representations in Lemma 9.3.5 could increase the generality of wildly ramified factors allowed in $\mathfrak{n}$. Furthermore, by inspection of the proof of Lemma 9.3.5, there are similarly results for $N=2$, but with some extra congruence conditions on the exponents of ramified primes dividing $\mathfrak{n}$.
10. Final Result
We now discuss implications for equidistribution of epsilon factors:
10.1. General Setup
We first describe sets of automorphic representations.
Definition 10.1.1.
In either the self-dual or conjugate self-dual case, let $N>0$, and consider $\mathfrak{n}$ a conductor for $F$ and $\lambda$ a (conjugate) self-dual infinitesimal character for $G_{N,\infty}$. Choose $\star$ to be either (conjugate) orthogonal or (conjugate) symplectic—denoted by “$\mathrm{orth}$” or “$\mathrm{symp}$” respectively.
Define $\mathcal{AR}_{N}^{\star}(\lambda,\mathfrak{n})$ to be the set of self-dual, type-$\star$, cuspidal automorphic representations $\pi$ of $\mathrm{GL}_{N}$ of conductor $\mathfrak{n}$ and with infinitesimal character $\lambda$ at infinity.
The most basic question we can ask about distribution of root numbers $\epsilon(1/2,\pi)$ in $\mathcal{AR}_{N,\star}(\lambda,\mathfrak{n})$, is to understand asymptotics of
$\frac{1}{|\mathcal{AR}_{N}^{\star}(\lambda,\mathfrak{n})|}\sum_{\pi\in\mathcal{AR}_{N,\star}(\lambda,\mathfrak{n})}\epsilon(1/2,\pi).$
In full generality, we also want to understand weighted distributions—if $\widetilde{f}_{S}$ is a test function at some set of places $S$ coprime to $\mathfrak{n}$, we can try to compute the asymptotics of
$\frac{\sum_{\pi\in\mathcal{AR}_{N}^{\star}(\lambda,\mathfrak{n})}\operatorname{tr}_{\widetilde{\pi}_{S}}(\tilde{f}_{S})\epsilon(1/2,\pi)}{\sum_{\pi\in\mathcal{AR}_{N}^{\star}(\lambda,\mathfrak{n})}\operatorname{tr}_{\widetilde{\pi}_{S}}(\widetilde{f}_{S})},$
where we recall $\widetilde{\pi}$ is the canonical extension of $\pi$ to $\widetilde{G}_{N}$.
We first describe the class of functions $f_{S}$ which our methods allow us to weigh the count by:
Definition 10.1.2.
In either the self-dual or conjugate self-dual case, let $S$ be a finite set of finite places of $F$ and $G\in\widetilde{\mathcal{E}}_{\mathrm{ell}}(N)$. Then a test function $\tilde{f}_{S}\in\mathscr{H}(\widetilde{G}_{N,S})$ is a full weighting for $G_{S}$ if the Fourier transform of $(\tilde{f}_{S})^{G}|_{Z_{G_{S}}}$ is strictly positive on all unramified characters.
Note that if $(\tilde{f}_{S})^{G}|_{Z_{G_{S}}}$ is supported in $K_{S}\cap Z_{G_{S}}$ (in particular, if $Z_{G_{S}}$ is compact), this simplifies to the condition that $(\tilde{f}_{S})^{G}(1)>0$.
Test functions that are not weightings produce sums over automorphic representations that cancel out except for an asymptotically trivial subset of $\mathcal{AR}_{\star}(\lambda,\mathfrak{n})$, making them much harder to study.
Some examples of full weightings for $G\in\widetilde{\mathcal{E}}_{\mathrm{ell}}(N)$:
  • Indicators of hyperspecials $\widetilde{f}_{S}=\mathbf{1}_{K_{S}\rtimes\theta}$ give unweighted counts.
  • For any test function $f_{S}\in\mathscr{H}(G_{S})$, we can find $\tilde{f}_{S}\in\mathscr{H}(\widetilde{G}_{N})$ such that $(\tilde{f}_{S})^{G}=f_{S}$ by [Mok15, Prop 3.1.1(b)] or [Art13, Prop. 2.1.1] (the argument for Mok’s extension part (b) also works in Arthur’s case).
  • If $\widetilde{f}_{S}$ satisfies that $\operatorname{tr}_{\widetilde{\pi}_{S}}(f_{S})\geq 0$ for all (conjugate) self-dual $\pi_{S}$ and there are tempered $A$-parameters $\psi_{S}$ for $G$ with general enough central character such that $\operatorname{tr}_{\widetilde{\pi}_{\psi_{S}}}(f_{S})>0$, then $f_{S}$ will be a full weighting similarly to Proposition 9.2.2.
10.2. Self-Dual Case
In the self-dual case, $\epsilon(1/2,\pi)=1$ unless $N$ is even and $\pi$ is symplectic-type/comes from $G=\mathrm{SO}_{N+1}$. We focus on this case:
Theorem 10.2.1.
Let $N>0$ be even, and consider $\mathfrak{n}$ a conductor for $F$ and $\lambda$ self-dual, symplectic-type, integral infinitesimal character for $G_{N,\infty}$ as in §5.1.2. Choose a finite set of finite places $S$ coprime to $\mathfrak{n}$ and a test function
$\widetilde{f}_{S}:=\mathbf{1}_{K_{S}\rtimes\theta}\star f_{S}\in\mathscr{H}^{\mathrm{ur}}(\widetilde{G}_{N}(F_{S}))^{\leq\kappa},$
such that $\widetilde{f}_{S}$ is a full weighting for $G=\mathrm{SO}_{N+1}$. For any self-dual automorphic representation $\pi$ of $\mathrm{GL}_{N}$, define coefficients
$w(\pi):=\operatorname{tr}_{\widetilde{\pi}_{S}}(\widetilde{f}_{S}).$
with respect to the canonical extension of $\pi$ to $\widetilde{G}_{N}$. Then, if there is place $v$ such that either
  • $v(\mathfrak{n})>N$,
  • $v(\mathfrak{n})$ is odd,
we have
$\frac{\sum_{\pi\in\mathcal{AR}_{N}^{\mathrm{symp}}(\lambda,\mathfrak{n})}w(\pi)\epsilon(1/2,\pi)}{\sum_{\pi\in\mathcal{AR}_{N}^{\mathrm{symp}}(\lambda,\mathfrak{n})}w(\pi)}=O_{\mathfrak{n}}(m(\lambda)^{-C}q_{S}^{A+B\kappa})$
for some constants $A,B,C$ with $C>1$. Otherwise,
$\frac{\sum_{\pi\in\mathcal{AR}_{N}^{\mathrm{symp}}(\lambda,\mathfrak{n})}w(\pi)\epsilon(1/2,\pi)}{\sum_{\pi\in\mathcal{AR}_{N}^{\mathrm{symp}}(\lambda,\mathfrak{n})}w(\pi)}\\ =\epsilon(1/2,\varphi_{\lambda})(-1)^{\sum_{v}v(\mathfrak{n})/2}R+O_{\mathfrak{n}}(m(\lambda)^{-C}q_{S}^{A+B\kappa})$
for some $R>0$ and where $\varphi_{\lambda}$ is the parameter with infinitesimal character $\lambda$ and trivial central character (i.e. the one coming from the discrete $L$-packet with infinitesimal character $\lambda$ on $G$).
Proof.
Let $\Sigma_{\eta}$ be the shape containing only cuspidal parameters of symplectic-type—i.e. $H(\Sigma_{\eta})=\mathrm{SO}_{2N+1}=G$. By Theorem 4.3.5 and Proposition 6.1.1,
$\sum_{\pi\in\mathcal{AR}_{N}^{\mathrm{symp}}(\lambda,\mathfrak{n})}w(\pi)\epsilon(1/2,\pi)=\epsilon(1/2,\varphi_{\lambda})\sum_{\psi\in\Sigma_{\eta,\lambda}}\operatorname{tr}_{\widetilde{\pi}_{\psi}^{\infty}}(\widetilde{E}^{\infty}_{\mathfrak{n}}\star f_{S})\\ =\epsilon(1/2,\varphi_{\lambda})S^{G}_{\Sigma_{\eta}}(\mathrm{EP}_{\lambda}(\widetilde{E}^{\infty}_{\mathfrak{n}}\star f_{S})^{G}),$
after noting that $S$ coprime to $\mathfrak{n}$ means that $\widetilde{E}_{\mathfrak{n},S}=\mathbf{1}_{K_{S}\rtimes\theta}$.
By similar arguments to [Dal22, §5.4], we have $(\widetilde{f}_{S})^{G}\in\mathscr{H}^{\mathrm{ur}}(G)^{\leq D\kappa}$ for some constant $D$ depending only on $N$ and $G$. Therefore, by the bound of Proposition 6.3.2 with our $S$:
(10.2.1) $(\dim\lambda)^{-1}S^{G}_{\Sigma_{\eta}}(\mathrm{EP}_{\lambda}(\widetilde{E}^{\infty}_{\mathfrak{n}}\star f_{S})^{G})\\ =\Lambda(G,\omega_{\lambda},(\widetilde{E}^{\infty}_{\mathfrak{n}}\star f_{S})^{G})+O_{\mathfrak{n}}(m(\lambda)^{-C}q_{S}^{A+B\kappa}(\widetilde{f}_{S})^{G}(1)),$
where $\Lambda$ is from (6.2.1) and for some $A,B,C$ with $C\geq 1$. By Corollary 7.2.2, we can compute
$\Lambda(G,\omega_{\lambda},(\widetilde{E}^{\infty}_{\mathfrak{n}}\star f_{S})^{G})=(\widetilde{E}^{\infty}_{\mathfrak{n}})^{G}(1)(\widetilde{f}_{S})^{G}(1)\\ =(\widetilde{f}_{S})^{G}(1)\begin{cases}0&\mathfrak{n}\text{ satisfies conditions}\\ M(-1)^{\sum_{v}v(\mathfrak{n})/2}&\text{else}\end{cases}$
for some constant $M>0$.
By a similar argument, now using Corollary 4.4.4,
$\sum_{\pi\in\mathcal{AR}_{N}^{\mathrm{symp}}(\lambda,\mathfrak{n})}w(\pi)=S^{G}_{\Sigma_{\eta}}(\mathrm{EP}_{\lambda}(\widetilde{C}^{\infty}_{\mathfrak{n}}\star f_{S})^{G})$
and
(10.2.2) $(\dim\lambda)^{-1}S^{G}_{\Sigma_{\eta}}(\mathrm{EP}_{\lambda}(\widetilde{C}^{\infty}_{\mathfrak{n}}\star f_{S})^{G})\\ =(\widetilde{C}^{\infty}_{\mathfrak{n}})^{G}(1)(\widetilde{f}_{S})^{G}(1)+O_{\mathfrak{n}}(m(\lambda)^{-C}q_{S}^{A+B\kappa}(\widetilde{f}_{S})^{G}(1)).$
By Theorem 9.4.1, $(\widetilde{C}^{\infty}_{\mathfrak{n}})^{G}(1)>0$. The result follows from dividing (10.2.1) by (10.2.2). ∎
In particular, when $\mathfrak{n}$ satisfies the bulleted conditions, the root numbers $\epsilon(1/2,\pi)$ for $\pi\in\mathcal{AR}_{N}^{\mathrm{symp}}(\lambda,\mathfrak{n})$ become equidistributed between ±$\pm 1$ as $m(\lambda)\to\infty$. This equidistribution is “independent” of any full weighting by Weyl-invariant polynomials on the self-dual Satake parameters at places coprime to $\mathfrak{n}$. There are also power-saving bounds on how fast equidistribution might be achieved.
If $\mathfrak{n}$ does not satisfy the bulleted conditions, then there is an “alternating” asymptotic bias between $\epsilon(1/2,\pi)=\pm 1$.
10.3. Conjugate Self-Dual Case
In the conjugate self-dual case, our techniques only apply when $N$ is even. Then, $\epsilon(1/2,\pi)=1$ unless $\pi$ is conjugate symplectic/comes from $G=U^{+}_{N}$. We focus on this case.
Isolating the conditions from Theorem 9.4.3:
Definition 10.3.1.
For a place $v$ of $F$, let $e_{v}$ be the ramification degree, and recall the numbers $b_{v}$ from Lemma 9.4.2 and $j_{v}$ from §9.3.2. Call a conductor $\mathfrak{n}$ (as in §3.1.2) valid if
  • For all wildly ramified $v$, $v(\mathfrak{n})\leq N/4$ or $v(\mathfrak{n})>2j_{v}-1$,
  • There is at least one $v$ such that $\max\{\lfloor v(\mathfrak{n})-(j_{v}-1/2)\rfloor,0\}\geq b_{v}+2/e_{v}$ or there are at least two $v$ such that $\max\{\lfloor v(\mathfrak{n})-(j_{v}-1/2)\rfloor,0\}\geq b_{v}+1/e_{v}$.
Call it 00-valid if
  • For all wildly ramified $v$, $v(\mathfrak{n})\leq N/4$ or $v(\mathfrak{n})>4j_{v}-2$,
(see note Note for potentially weaker conditions.)
We repeat again that these conditions, while not comprehensive, are quite weak: for example, only the second part of the valid condition doesn’t involve wild ramification and it is automatically satisfied if there are no roots of unity in $1+\mathfrak{D}_{E/F}$.
Theorem 10.3.2.
Let $N>2$ be even and let $\lambda$ be a conjugate self-dual infinitesimal character of $G_{N}$ as in §5.1.2. Let $\mathfrak{n}$ be a conductor for $F$. Assume that $\mathfrak{n}$ is valid or, if $\lambda$ corresponds to the trivial central character, possibly instead 00-valid.
Choose a finite set of finite places $S$ coprime to $\mathfrak{n}$ and a test function
$\widetilde{f}_{S}:=\mathbf{1}_{K_{S}\rtimes\theta}\star f_{S}\in\mathscr{H}^{\mathrm{ur}}(\widetilde{G}_{N}(F_{S}))^{\leq\kappa},$
such that $\widetilde{f}_{S}$ is a full weighting for $G=U^{+}_{N}$. For any conjugate self-dual automorphic representation $\pi$ of $G_{N}$, define coefficients
$w(\pi):=\operatorname{tr}_{\widetilde{\pi}_{S}}(\widetilde{f}_{S})$
with respect to the canonical extension of $\pi$ to $\widetilde{G}_{N}$. Then, if there is a ramified place $v$ such that either
  • $v(\mathfrak{n})$ is half-integral,
  • $v(\mathfrak{n})>N/2$,
we have
$\frac{\sum_{\pi\in\mathcal{AR}_{N}^{\mathrm{symp}}(\lambda,\mathfrak{n})}w(\pi)\epsilon(1/2,\pi)}{\sum_{\pi\in\mathcal{AR}_{N}^{\mathrm{symp}}(\lambda,\mathfrak{n})}w(\pi)}=O_{\mathfrak{n},\omega_{\lambda}}(m(\lambda)^{-C}q_{S}^{A+B\kappa})$
for some constants $A,B,C$ with $C>1$ and where $\omega_{\lambda}$ is the central character of the finite-dimensional representation of $G_{\infty}$ corresponding to $\lambda$.
Proof.
This is the same argument as Theorem 10.2.1 except we use the analogous results for the conjugate self-dual case.
Theorem 4.3.6 and Proposition 6.1.1 interpret
$\sum_{\pi\in\mathcal{AR}_{N}^{\mathrm{symp}}(\lambda,\mathfrak{n})}w(\pi)\epsilon(1/2,\pi)=\omega_{\infty}(a)^{-1}\epsilon(1/2,\varphi_{\lambda})^{-1}S^{G}_{\Sigma_{\eta}}(\mathrm{EP}_{\lambda}(\widetilde{E}^{\infty}_{\mathfrak{n}}\star f_{S})^{G})$
for some choice of $a/\bar{a}=-1$. Then, by Theorem 6.4.1:
$(\dim\lambda)^{-1}S^{G}_{\Sigma_{\eta}}(\mathrm{EP}_{\lambda}(\widetilde{E}^{\infty}_{\mathfrak{n}}\star f_{S})^{G})\\ =\Lambda(G,\omega_{\lambda},(\widetilde{E}^{\infty}_{\mathfrak{n}}\star f_{S})^{G})+O_{\mathfrak{n}}(m(\lambda)^{-C}q_{S}^{A+B\kappa}(\widetilde{f}_{S})^{G}(1)),$
where our bulleted conditions on $\mathfrak{n}$ give that
$\Lambda(G,\omega_{\lambda},(\widetilde{E}^{\infty}_{\mathfrak{n}}\star f_{S})^{G})=0$
by Corollary 8.4.1.
On the other hand, Corollary 4.4.6 gives that
$\sum_{\pi\in\mathcal{AR}_{N}^{\mathrm{symp}}(\lambda,\mathfrak{n})}w(\pi)=S^{G}_{\Sigma_{\eta}}(\mathrm{EP}_{\lambda}(\widetilde{C}^{\infty}_{\mathfrak{n}}\star f_{S})^{G})$
and Theorem 6.4.1 again gives
$(\dim\lambda)^{-1}S^{G}_{\Sigma_{\eta}}(\mathrm{EP}_{\lambda}(\widetilde{C}^{\infty}_{\mathfrak{n}}\star f_{S})^{G})\\ =\Lambda(G,\omega_{\lambda},(\widetilde{C}^{\infty}_{\mathfrak{n}}\star f_{S})^{G})+O_{\mathfrak{n}}(m(\lambda)^{-C}q_{S}^{A+B\kappa}(\widetilde{f}_{S})^{G}(1)).$
Let the support of $\mathfrak{n}$ be $S_{0}$. Then by (9.2.2),
$\Lambda(G,\omega_{\lambda},(\widetilde{C}^{\infty}_{\mathfrak{n}}\star f_{S})^{G})=\sum_{\omega_{S_{0}}\omega_{S}\in\Omega_{\lambda}(K^{S,S_{0},\infty})}E^{\mathrm{pl}}((\widetilde{C}_{S_{0},\mathfrak{n}})^{G}f_{S}^{G}|\omega_{S_{0}}\omega_{S})$
for the set of characters $\Omega_{\lambda}(K^{S,S_{0},\infty})$ of $Z_{G,S\cup S_{0}}$ defined there. Each of the summands factors into:
$E^{\mathrm{pl}}((\widetilde{C}_{S_{0},\mathfrak{n}})^{G}|\omega_{S_{0}})E^{\mathrm{pl}}(f_{S}^{G}|\omega_{S}).$
Since $f_{S}^{G}$ is unramified, the second factor is zero unless $\omega_{S}|_{K_{S}}$ is trivial. Therefore, we can rewrite $\Lambda(G,\omega_{\lambda},(\widetilde{C}^{\infty}_{\mathfrak{n}}\star f_{S})^{G})$ as
$\sum_{\omega_{S_{0}}\in\Omega_{\lambda}(K^{S_{0},\infty})}E^{\mathrm{pl}}((\widetilde{C}^{\infty}_{\mathfrak{n}})^{G}|\omega_{S_{0}})\left(\sum_{\omega_{S_{0}}\omega_{S}\in\Omega_{\lambda}(K^{S,S_{0},\infty})}E^{\mathrm{pl}}(f_{S}^{G}|\omega_{S})\right)$
Since $f_{S}$ is a full weighting, the values of the inner sum are always strictly positive. By Theorem 9.4.3 or 9.4.4, trace-positivity of $\widetilde{C}^{\infty}_{\mathfrak{n}}$, and validity of $\mathfrak{n}$, all the terms $E^{\mathrm{pl}}((\widetilde{C}^{\infty}_{\mathfrak{n}})^{G}|\omega_{S_{0}})$ are non-negative and at least one of them is strictly positive. In total, $\Lambda(G,\omega_{\lambda},(\widetilde{C}^{\infty}_{\mathfrak{n}}\star f_{S})^{G})>0$.
Therefore, we get our result by dividing these two computations. ∎
Because we cannot prove Conjecture 8.4.3, we do not have a non-equidistribution result analoguous to the second part of Theorem 10.2.1. However, we expect that it does hold for valid (or 00-valid) $\mathfrak{n}$ that do not satisfy the bulleted conditions.
As a replacement, we could use the $\tilde{f}_{\mathfrak{n}}$ from Corollary 8.4.1 in our trace formula computation instead of $\widetilde{E}^{\infty}_{\mathfrak{n}}$. This would prove a non-equidistribution result for an average over $\pi$ with $c(\pi)|\mathfrak{n}$ that is weighted by a product depending on $\mathfrak{n}/c(\pi)$ of the (signed!) coefficients from Proposition 4.2.5.